A Bibliography of Publications in Notices of the American Mathematical Society (2010–2019)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

#1 [Dal13b].

α [GT15]. β [GT15]. D [Ber19b]. G [BS13]. G_{2} [Kar11]. K [Zak19, Mor18]. m [Duj16]. R [AdS19]. p [MS19, Sil13]. π [Mar16b]. π^{2} [BBMW13].

-adic [MS19, Sil13]. -functions [BS13]. -manifold [Kar11]. -Schur [Mor18]. -theory [Zak19]. -tuple [Duj16]. -wave [Ber19b].

//www.topsecretrosies.com [Gre12b].

000-year-old [Car13].

1 [Ano16a]. 1-dimensional [BG11b]. 100-year [May19b]. 120-cell [SS15a]. 1900 [Pen11]. 1904 [Zit11]. 1st [RTKB12]. 1st- [RTKB12].

2 [Ano16b]. 2.0 [Bea11]. 2000 [Ben10]. 2001 [Ano11a]. 2010 [Ano10e, Ano10b, Ano10c, Ano10d, Ano11v, HWWA11, Fol11]. 2011 [Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11m, Ano11i, Ano11h, Ano11g, Ano12g]. 2012 [Ano12k, Keh12a, Keh12h, Keh12c, Keh12b, Keh12d, Keh12e, Keh12f, Keh13a, Smi14]. 2014 [dMPQZ15, Fol16]. 2018 [Ano18b, Ano18a, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18i, Ano18h, Ano18j, Keh18]. 21st [Edw12]. 21st-century [Edw12].

248-dimensional [BG11b]. 2nd [RTKB12]. 2nd-line [RTKB12].

3-headed [Mur12]. 3-manifold [Ada18]. 3N [Zie11].

91st [Jac14].

Abel [Ano11m, Ano12k, Ano18-39, Ano19a, DS18, DS19, RS11, RS16a, RS17]. Abhyankar [Ano14e]. ability [Tho14]. Abrosimov [BKK^{+}12]. abstract [Ros14]. academic [Sen13]. Accumulation [TW19]. Achievement [Ano18-27]. Action [Pet10, Kob18]. actually [GTTZ10, Ser14]. acylindrical [Kob18]. Ad [BAD^{+}18, CMJ^{+}18, CST^{+}17]. Adams [DL16a]. addition [Lar18]. adic [MS19, Sil13]. Advanced [BDJ^{+}17]. adventures [Coo10]. aesthetics [Fen12]. affiliation [PM12]. affine [Fox12]. Africa [BGL^{+}18]. African [GK18, Hou18b, Hou19b]. after [Bae19, Coo12]. against [Hol13]. Age [BCLO11, Chu10b, Pér17]. Agnesi [Dum19d]. ago [Ano17a, Gra16, Qui12]. AIDs [RTKB12]. airliner [CGM^{+}15]. airlines [CGM^{+}15]. Alan [Lip11, Coo12, Don14, Mas19, Avi14, Smi14]. Albert [Ano18a, Dra15a]. Aleksander [Ano17b]. Alexander [BKK^{+}12, Car15, GM19]. Alexandre [Ano16a, Ano16b]. Algebra [DPV14, KN08, Abr16, Ano18f, BG11b, Flø15, Keh12b, MS19, Raz13, Sob12]. Algebraic [De 19, ACP18, Ano16e, Gar10, Str17, Zak19]. algebras [Set10, Shl18, Vid19]. algorithm [BKK14]. algorithms [HO19, Shu12]. Alice [Ano18k, BSG^{+}10]. Alignment [Don14]. Alliance [GK18]. almanac [Axl17]. Almost [Duk14]. Am [Ger16]. Am./Farrar [Ger16]. Amer [Ewi11a]. America [Por10]. American [Ben10, Bat11, GK18, Hou18b, Hou19b, Nat14, Zit11]. Amie [Ano18l]. amplituhedron [BT18]. AMS [Keh12a, Ano18-27, Fre18, Jac16]. Anabelian [Wic16]. analysis [Ano16c, Ano16d, GW19, KW18, Nee19, Oko19, Per19a, Ros14, FF12, Ros19]. analytic [Bar13, DKNS10]. Anatole [Has19]. Ancient [Dam10]. André [Aud11, Wei18]. Andrea [Ano18m]. Andreevich [KKK^{+}14]. Andrew [RS17, Ano13a, Dar17, SRM^{+}17]. Andrews [Jac11]. animal [HH10]. Animating [Gof15]. Animus [And16]. anniversary [CR15, Spe10b, Zie18]. annotated [Lip11]. answering [Duk14]. Antibiotics [TY17]. aperiodic [BDG16]. apologies [McC18]. Applications [Per00, PN12, Ano16g, Ber19b, CMT17, E19, GW19, MS19, Mór14, NPS14, SS17, Sto19, Tor19]. applied [Ano18g, MOT10, Pap17]. apprenticeship [Dam10]. approach [Saa13]. approaches [Ran14]. approximate [Gre12a]. approximately [Fel14]. approximation [OSS19]. arch [Oss10]. Archimedes [GDG^{+}15, PR12]. areas [KvP16]. Arf [Sna13]. Arithmetic [Fre10b, Han12, Wic18]. Arnold [Ano12f, Ano12m]. Arroyo [Lan18c]. art [Ati10, YHH19, RTKB12]. artful [Wal15]. Arthur [Ano15c, Ano18-27]. article [Pér11]. arts [Phi12b]. Arveson [Ano15i]. Aschbacher [Keh13a]. Assessment [HO19]. Assn [Ewi11a]. Assoc [Por10]. Association [And16]. Assumptions [KM10]. astronomy [Lin12]. Astrophysics [KN08]. asymptotic [KW18]. Atiyah [Ano19b, CK19]. atom [Bla11b]. Attitudes [Zak19]. Audin [Jac17]. Austin [AK12]. Autobiography [Wu17]. automatic [Row15]. Ávila [Ver13]. avuncular [RR13]. Award [Keh12g, Ano12g, Ano18b, Keh13b, Ano11v]. awarded [Ano10b, Ano10f, Ano10c, Ano10d, Ano11j, Ano11n, Ano12b, Ano12e, Ano14a, Ano14c, Ano15a, Ano18-39, Ano19a, Keh12h, Keh13a, Keh16, Keh18]. Awareness [Rob10, Bar17]. awesome [MOT10]. axiom [APW13].

B [Ano12c, Ano12d, Ano15i, Ano17c, BF12b]. Bôcher [Ano11b]. back [Cor13, Hal14, Vid19]. backtest [BBdPZ14, BBLZ14]. Baernstein [Dra15a]. bake [Mar16b]. Baker [Mas19]. Balzan [Ano11n]. Banach [Giv19, Ros11]. Barry [Ano16c, Ano16d]. Baseball [D’A10, Buk18, Por10]. based [OH13]. Basic [Daw18, Fel14, Van15]. Basil [Ano13e]. basis [Pet19]. BBVA [Keh13b]. be [Blu16, Ewi11a, Mil16, Por10]. beautiful [Fen12]. beauty [Ham19]. became [Sac11]. bedside [Don18]. been [Mer19a]. Beethoven [FGLG13]. beginning [Far12a]. behavior [HH10]. Behind [Fre10b, Was18, Dal12b]. Beilinson [Keh18]. beliefs [Fol12b]. Bell [GB13]. Bellevue [Sch13b]. Belonging [Pet18]. Benford [BH17]. Benoît [BF12a, BF12b, Ano12c, Ano12d]. Bento [AK12]. Berger [LBG^{+}17]. Berlin [Oor13]. Bertrand [Ano18c]. best [Fol11]. better [DEG^{+}17]. between [GB13, Mil15, Pap19]. Beyond [DSW16, Mar12, MOT10]. bias [Grc10]. bicycles [Pap17]. big [NPS14, Zor12]. biholomorphic [Bed12]. biography [Sti14]. Biology [CCBB^{+}16, Ada11, Fri10, Mag15, Mil15, Mur12]. Birds [Fol16]. Birkhoff [Ano18g, Keh12a]. Birman [Ano18-29, Mar19]. birth [Ano15f, Ano15g, PM12, Row19, Str15]. Black [Ald11, DSW16, Goi18, Wil18a, GHPW18, Pei17]. blackboard [Don18]. Blackwell [Ano11t, Hou19a]. blender [BCDW16]. Bloomsbury [Bau15]. blowing [Abb19]. Bodacious [KM10]. Bolyai [Dén11]. Book [Bla14a, Hod11, O’S11, AK12, Ald11, Arc13, Ari13, Ari15, Aud10, Aud11, Aus18, Avi11, Avi14, Avi16, Axl17, Bae19, Bat14, Bau15, Bay12, Bea11, Ben10, Bla11a, Bla11b, Bla14b, Blu16, Bom16, Bra10, Bre16, Bur14, Car13, Col16, Cur12, Dal12a, Dal12b, Dal13a, Dal14a, Dal14b, Dal17, Dam10, Dav11, Dav13, Daw18, DeT17, Des11, Dra15b, Ede13, Eis10, Elr19, Ewi12, Far12a, FF12, Fel14, Fol11, Fol12a, Fol12b, Fol16, Fre10a, Fuj10, Ger16, Glu14, Gof15, Gol19, Gra13a, Gra17, GT15, Häg11, Häg12, Häg13, Han12, Hay13, Hay14, Hay17, Hel13a, Hel13b, Hit11, Hub19, Hur14, Joh14, Jon12, Jon16, Jon19, KH19, Kna14, Kno15, Kob16, Lar10a, Lar18, Law19, LS16, Lin12, Lip11, Lyn13, Mag13, Mag14, Man15, Man10, Mar10, Mar16b, McC17]. book [McC18, Mil18, Mil19, Mon13, Mum10, Mum18, O’S18, Oor13, Pen11, Phi12b, Poo17, Por10, Que19, Roc11, Roi10, Roi17, Ros19, Row19, Sch13b, Sen13, SS15c, Sil10, Smi14, Spe10a, Ste11, Ste14, Sti14, Str15, Van12, Van15, Wal15, Wig14, Zor12]. Books [AK12, Daw18, Fel14, Spe10a, Van15, Smi14, Roi17]. Boole [Cha19]. Boolean [BS19]. Boothe [Dum19a]. bordered [DHV11]. Borel [For18]. Bott [Ano13f]. bottle [Fra12]. bounded [DG14]. Bourgain [Ano10e, Keh12h]. box [DSW16]. braid [Sch18a]. brain [Pap19, Smi14]. Brave [KM10]. breaking [Lar18]. breakout [Dum19c]. Breakthrough [Coh17, Dum19c, O’S11]. Brennie [Ano18e]. Bridges [Fen12]. bridging [Bou13]. brief [Pér11]. brilliant [Sau16]. bringing [KSZ17]. broken [DMVW10]. bronze [FF10]. Brouwer [Joh14]. Browder [Ano18u]. Bryant [Jac16]. build [TY17]. building [Dum19a]. bunk [Häg11]. Burden [Ede13]. Busemann [Pap18]. Business [Ste11].

C [Ano18-28, Sch10]. Caffarelli [Keh13a]. Calabi [Tos18]. calculate [Bou13]. calculus [Bra10, Dal14a, Don19, Häg12, Ran14].

Caltech [GB13]. Cambridge [Car13, Phi12b]. Can [Sax18, Sch11, DPV14, Por10]. Cancer [Dur13]. Can’t [GS10]. Capo [Car13]. Cárdenas [AGLP^{+}13]. cardiac [Cai11]. care [BC13]. Cari [Reh10]. Carina [Ano18n]. Carmichael [Mer19a]. Carolyn [Ano18o]. Cartan [BRH10, CDD^{+}10, DKNS10, EMS11, SA10]. case [Cha19, FH11, Sch18b]. Casebook [Van15]. Catalan [BBMW13]. catastrophe [Rot17]. categories [GJO19]. Cathedral [Bla14a, Bla14b]. Cathleen [Ano18-35, Jac14]. Cats [Sil14]. Cauchy [Bar13, DT10]. celebrates [Ric14]. Celebrating [FF10, AABT^{+}19]. cell [SS15a]. centenarian [Hou19b]. centenary [Ano15f, Ano15g, Avi16]. centennial [Hou19a]. central [Flø15]. century [Bae19, Car13, Duk14, Edw12, Has19, Qui12, O’S11]. century-long [Car13]. chains [D’A10]. challenges [Cai11]. Chan [Ano18-37]. chance [Hub19]. Chang [Ano18k]. change [CHKM11, Wal10]. changed [Kuz15]. changing [NPS14]. chaos [Zel19]. charlatanism [BBdPZ14, BBLZ14]. Charles [CST^{+}17]. charts [TW19]. Chasing [Lin12]. chat [RR13]. Chelsea [Ano18p, DL18a]. Chen [KW17]. Chern [Ano10b, Ano11s]. Chevalley [Ano18d]. Chicago [Häg11, Häg11]. Chinese [Ji16]. Christodoulou [Ano11j]. chronology [Dia13]. Chudnovsky [Ano18-33]. Chung [Ano18t]. Circle [Daw18]. circulant [KS12b]. citation [Yon14]. Civita [Row19]. Claire [BAD^{+}18]. Clare [Dum19a]. Clarence [Hou18a]. Clarke [Ano17a]. class [JS14]. Classic [Dal13a]. classical [Ano16g, Ros14]. classification [Mon19, Ros11, Sol18]. Clerk [Sil14]. Clifford [KK18]. climate [Wal10]. climb [Sha19]. Closed [BC13]. closer [KSZ17]. Co [Han12, Shl18]. Coble [Gla13]. Codá [DL16c]. codebreaker [Ano11o]. Cohen [Sar10]. cohomology [Tu11]. Coimbra [Fen12]. coin [Orn13]. coincidences [Dal14b]. Cole [Ano11c, Ano18f, Keh12b]. Coleman [Bal19]. Colin [DL16a]. College [Pei17]. colored [Zie11]. Combinatorial [Pak19, Nel11, Yon14, Zak19]. combinatorics [Mor18, Str17]. commemoration [CR15]. commercial [War13]. communication [Set10]. communities [AJ19]. community [Dum19c, GK18, HS19, VJCC19]. commutative [MS19]. comparisons [Saa13]. competitions [Gar19]. completeness [BG11a]. complex [DKNS10, EMS11, Gar17, Moo10, WS11, Gra17]. complexity [DMVW10, Vid19, Roc11]. component [Spe10b]. comprehensive [Ros19]. compromise [CCF^{+}12]. Computability [DH19, Lip11]. computation [BB11, BBMW13, LS16]. computational [CXM^{+}14]. Computer [DPV14, Ber19b, Car13, Kal16, War13]. computers [Gre12b]. Computing [Mor18, MOT10, Wig14]. Conant [Ano11a, Ano18i, Keh12c]. concept [Men17, SS15b]. concepts [BBH^{+}14]. Conceptual [Coh17]. concerning [Gan18]. conclusions [Wil18b]. confidence [Dum19a]. conflict [SZ13]. Conformal [Sob12]. Congress [Sax18, Zit11]. congruence [Dup12]. conjecture [BR18, Chu10a, DKM17, GGW14, Ios19, Mer19a, Mon19]. conjugate [BL12]. Constance [Ale11]. constant [BBMW13]. continuity [KS12a]. contribution [Bat11]. contributions [Hou19a, Hou19b]. conversation [Ewi11a, KW17]. convex [DHV11, Gho18, Zie18]. conviction [Hel13a]. Conway [Sch13a, Mil18]. cooperation [Dum19b]. cop [Bon12]. correct [Fel14]. corrections [Grc13]. cost [Gab10]. count [Wic18]. counterexamples [BBZ16]. country [Car15]. course [Ros19]. court [BD17, Sax18]. Cox [Loh19]. CR [HR17]. Cradle [Sil14]. Crafoord [Keh12h]. Crashing [MOT10]. Creating [Dum19a, Wal15]. creativity [Glu14]. crest [Mon13]. Critique [Yon14]. Cronin [Abo19]. Crown [Ste11]. cryptographers [PQ13]. cryptographic [BS19]. Cryptography [Cas10, Fre10b, KM10]. crystal [KŠŠ12]. crystallization [Rad17]. crystals [BBF11]. CT [Dal12a]. cubic [Wic18]. cultural [Bou13]. culture [Sen13]. cumulant [NŚ11]. curious [Mil18]. curse [Kol15]. Curto [Ano18n]. curvature [CM16]. curve [DLR17]. curves [Goi19]. curvitas [SS15b]. cyclic [RSW14]. cyclotomy [GHL15].

D.C. [Por10]. damping [OH13]. Dance [Sch18a]. dangerous [Ger16]. Daniel [Ano12a, BHMM15, Boy13, GK11]. Data [Mag12, Ewi11b, Hay14, Mag15, NPS14, Nee19]. databases [BT13]. Daubechies [Ano11i, Ano12b, Ano18y, Keh13b]. David [Ano18g, Ano11t, GM19, Hou19a, Jac13a, Zit15, Oor13]. dawn [Sil10]. Day [Rou19, Dal14b]. days [Zor12, McC17]. DC [Ewi11a]. decide [GS10]. Decoding [Car13]. decomposition [Con12, DEG^{+}17]. decoupling [dlP19]. Deep [Hel13b]. default [BBS13]. Deformations [DHV11]. degenerate [CP10, Tum16]. degrees [Mon19]. Dehn [Pei17]. Deligne [RS14]. delusions [Bre16]. democracy [Hod11]. Democritus [Wig14]. derivatives [Ser14]. derived [Saa13, Vez11]. Descartes [Ran14]. Descriptive [Ros11]. deserts [Smi14]. design [Loh19]. destroyed [Nat14]. destruction [Poo17]. detect [BG11b]. development [Arc13]. dialogue [HH10]. Dickson [DSG17]. Did [BG11b]. Dido [Ban17]. Diestel [Ano18-44]. Differential [Mil11, CW17, DMP10, Riv16, Shu12]. Digital [Bla14a, BCLO11, Bla14b]. digits [BBMW13]. dilemmas [Avi11]. dimension [Rei12]. dimensional [BG11b, Don18, Way11]. Dimensions [YN11, Mar16a]. Diophantine [Duj16]. Dirac [Bla11b]. Directional [JP11]. discontinuous [Shu12]. discover [Car13, Jon16]. discovery [FH11]. Discrete [BV11, CW17]. disease [Nee19]. distinguished [Ano18b, Keh12g]. distortion [Pap19]. diversity [Tho14]. divide [Bou13]. divine [Dam10]. Division [Set10, BKK14]. do [Dal12b, Man15, Sau16]. Does [Tho14, CCF^{+}12, GTTZ10, Ser14]. Doing [Hay14]. domain [Ran12]. Donald [GCCUD18]. Donoho [Ano14a]. don’t [HS10]. Doob [Ano11d]. door [Hal14, KR12]. Dorman [Spe10a]. Dorothea [Pei17]. double [Ran14]. doubter [Axl17]. Douglas [YHH19]. dramatist [Reh10]. Drinfeld [Keh18]. driven [Ewi11b]. drops [CMT17]. Duel [Sil10]. Duistermaat [Ano11q]. Dürer [Gho18]. Dusa [Ano18q]. DVD [Cor13]. Dyer [BCCTS19]. Dynamical [Str17]. dynamics [CXM^{+}14, Has19, Hyn19, vdBL15].

E. [GB13]. Earle [KK18]. Early [RT18a, Jac16, Lin12]. earth [Jon16]. Earthquakes [Wal10]. East [SZ13]. Ebola [CCBB^{+}16]. eclipse [Lin12]. economics [Gui19]. Edgerton [KR12]. edition [Kno15]. Edizioni [Bom16]. education [BV11, Ewi12, KSZ17, PM12]. Edward [And16, Oog15]. Edwards [DL17b]. Eells [CR15]. Effectiveness [Nic12]. effects [BBdPZ14, BBLZ14, MOT10]. efficiency [BD17]. Efficient [Shu12, BKK14]. Ehrenpreis [Ano11r]. eigenfunctions [HS10]. eigenvector [Saa13]. eigenvectors [Stu16]. Einstein [Fre18, Row19]. Eisenbud [Ano11e]. Eisenhart [Gla13]. Eisenstein [BBF11]. elections [TW19]. electronic [Smi14]. electrophysiology [Cai11]. elementary [San11]. elimination [Grc11]. Ellen [Ano15e]. Ellipse [Adl12]. ellipsoids [KL14]. elliptic [DLR17, Goi19]. Eloquent [Adl12]. else [Str18]. Elsevier [Avi14]. emerging [Pen11]. Emily [Ano18r]. enabled [Ber19b]. encyclopedia [Slo18]. endomorphism [Dek11]. Endre [RS13]. engineer [Dav13]. Enlightening [Ari15]. enough [NPS14]. Enriques [BG11a, Mum11]. entertaining [Hay17]. Entertainment [And16, And16]. entire [NS10]. entry [CGM^{+}15]. Environmental [BFH10, Rot17]. envy [BKK14]. envy-free [BKK14]. epic [Daw18, Roi10]. Episodes [NB18]. equations [CP10, DMP10, Gui19, Riv16, Rob15, Shu12]. equilibrium [SW16]. equity [Gar19]. equivariant [Ger19, Tu11]. Erdos [Ano15f, Ano15g]. ergodic [Sil16]. Eric [Ano11p, Jac13b]. Ernest [NB18]. Ernst [Buc13]. Errors [Grc13]. Escher [Sch10]. essays [Kob14]. essential [Rei12]. establishment [FH11]. ethical [Rob10]. Euclidean [OH13]. Eugene [DSG17]. Euler [BBH^{+}14, Mar10]. Euphemia [KSZ17]. Europe [BRH10]. European [Mon13]. Éva [Ano18s]. Evanston [Aud11]. Évariste [Cur12, Neu11]. Evelyn [DL17a]. events [AABT^{+}19, Dal14b]. Every [Lin10, Dal14b, Was18]. Everything [Dun19a, Dun19b]. evidence [Ewi12]. evolution [Ari13]. examination [KR10]. example [Tho14, Zie18]. Excerpt [Rib19]. Exit [And16]. expanded [Kno15]. expansion [Zit11]. experimentation [BB11]. experiments [War13]. expert [Dun19a, Dun19b]. Explanations [Far12a]. Exploratory [BB11]. Exploring [Riv16]. explosions [MOT10]. expression [Rot17]. extensions [BBZ16]. Extensive [Dum19b]. exterior [Flø15]. extraordinary [Cha19].

F [Ano18x, Hou18a, RS16a]. Fa [KW17]. face [Dén11]. faces [Kob14]. fair [BKK14]. Falconer [Ios19]. Faltings [Ano15a]. family [Cha19, SP18]. Fan [Ano18t, Lin10].

Farey [Phi12a]. Farrar [Ger16, Str15]. Fasquelle [Hur14]. father [Wei18]. Fe [Hay13]. Federer [Ano12h]. Federico [AGLP^{+}13]. Federigo [BG11a]. Fefferman [CST^{+}17]. Felix [Ano18u]. fellowships [Dum19c]. female [Gre12b]. Fenchel [BL12]. Fermat [BBH^{+}14, Per00]. Fermi [Yon14]. Fern [Ano18v]. Fernando [DL16c]. Ferrand [Pan18]. few [Was18]. Fibonacci [Han12]. field [JS14]. Fielding [Ano13a]. Fields [Ano10f, Gra16, HWWA11, Kol15, dMPQZ15]. fiftieth [CR15, Zie18]. fighting [Des11]. figure [Dum19e, Sac11]. figures [Coo10, Dum19e, Hay13]. Film [Gre12b, And16, Ewi11a, Sau14, Sau16]. Films [And16, Sau16]. finally [Fri15]. financial [BBdPZ14, BBLZ14]. Fingerprint [BT13]. finite [Sol18]. Firecracker [And16]. first [Bon13, GM19]. five [KŠŠ12]. five-fold [KŠŠ12]. fixed [Tu15]. flag [Raz13]. Flashbacks [Bel16]. flattening [Abb19]. fleeing [Aud10]. flies [Mil15]. flight [CGM^{+}15]. flood [May19b]. flows [LSW19]. fluid [CXM^{+}14, Way11]. flux [Olv11]. fold [KŠŠ12]. Fonseca [Ano18z]. Forbes [Ano16f]. Forbidden [Far12b]. form [Fol10]. forms [BC13]. Formula [Adl12, BD17]. forty [Mil11]. forty-six [Mil11]. found [Fri15]. Foundation [Grä13b, Keh13b]. foundations [Daw18, Bae19]. four [Smi14]. Fourier [Ros14, Sto19]. foxes [Rob19]. fractal [AH14]. frame [Hei13]. frameworks [SS17]. Francesco [GB12]. Francisco [AGLP^{+}13]. Frank [Ano18f, Ano18e]. Free [Ser14, BKK14, GTTZ10, NŚ11, Sax18]. freedom [BDJ^{+}17]. Fresh [GB12]. Friedlander [Ano11p, Jac13b]. Friedrich [Ano14b]. friend [CK19]. friendship [Bra10]. frogs [Fol16]. frontier [Coo10]. Fruit [Mil15]. Fulkerson [Ano10a]. function [BS19, Duk14, NS10]. Functions [BCLO11, Bea11, BS13, Beu14, Gun10, KvP16, Mor18, WS11, Dra15b]. Fundamental [KN08].

G [Ano18-27, Hou19b, Nai13, YHH19]. Gabor [Oko19]. Gaetana [Dum19d]. Galileo [Phi12b]. Galois [Bon13, Edw12, Neu11, Cur12]. Gambler [Dal12b]. games [Row14]. Gamow [CMT17]. gang [BBH^{+}14]. gap [BD17, Fri15]. Garabedian [KLN^{+}14]. Garden [Dys13, Ano16c, Ano16d]. Gardner [Ano11l]. gas [CLWZ11]. gateway [Oss10]. Gauge [Tum16]. Gaunce [May19a]. Gauss [Ano10c, GHL15, PR12]. Gaussian [Grc11, NS10]. Gelfand [Ano13c, Ano13b, Bei16]. gem [Mar10]. gender [Gar19]. General [Per00, Row19]. generalist [Grc10]. generated [Gui19]. genesis [Tu15]. genius [NB18, O’S11, Ono10, Sau14, Mil18, Avi16]. gentle [BS13]. Gentzen [Avi16]. geodesic [OH13]. geography [Pap19]. Geometric [Tor19, CM19, KW18, Sob12]. geometries [DT10]. Geometry [Mag15, YN11, Ano16e, Ard18, BR18, CW17, DL19, DEG^{+}17, DKNS10, DMVW10, Don18, Gui19, Hol16, Kat17, Lan18a, Moo10]. George [Ano18g, Dum19b, Jac11]. Geraldo [Ver13]. Gerhard [Ano11k]. Germain [Mer19b]. German [Sen13]. German-speaking [Sen13]. Germany [Aud10]. Gerrymandering [Sob17, BD17]. Gerrymanders [Sax18]. Gertrude [Loh19]. gets [Mer19b]. getting [EMS11]. Gettysburgians [Gla13]. ghost [Man10]. giant [Spe10b]. Gifted [GHPW18]. Gigliola [Ano18w]. Gioia [Reh10]. girls [AK12]. Giroux [Ger16, Str15]. given [HN17]. glimpse [CW17]. Glimpses [Ano12c, BF12a]. global [Rot17]. goal [KSZ17]. goes [BD17]. Going [Bor12, Ano12j, Ano13i]. golden [Pér17, Spe10b]. Goldwyn [Sau16]. good [Was18]. Gordon [Ano13e, Ano18o]. Goss [GM19]. Gösta [Hel13a]. got [Dal12b, Man15]. Gowers [DL16d]. Grace [Aue19]. graduate [Dum19c]. Graham [Ano18t]. grandmaster [Abb19]. Graph [Chu10b, HiK13]. graphics [Roi17]. graphon [Gla15]. graphs [SP18]. Grasset [Hur14]. Grauert [Ano14f]. Gravitational [BGY17, HN17]. Gravity [Pet10]. great [BZ11, Hub19, Mil16]. green [HN17, Hay13, Kha16]. Greenhouse [CLWZ11]. Grenander [Ano18j]. Gromov [RS10]. Grothendieck [Rib19, Ano16a, Ano16b, Ano16e, Car15, Hon16, IBB^{+}10]. group [CF11, Gar10, Gre12a, JS14, Kob18, KLLR19, And16]. groups [Sch18a, Sol18]. growth [DRAS16]. Grünbaum [Zie18]. guesses [Ewi12]. guide [Dum19b, Dun19a, Dun19b, KH19]. guided [Lip11, Roc11]. Guzman [Ano18-40].

H [CR15, Hou19a]. Haagerup [CJMR16]. Haas [DL18b]. Hackensack [Dal14a]. Hadwiger [Chu10a]. half [Has19]. half-century [Has19]. Hall [Don14]. Halmos [Ewi11a]. halting [CC10]. Hamilton} [Mer19b, Ano11j]. Handbook [BCLO11]. Handbooks [Bea11]. Hans [Ano14f, Ano18x, Bru15]. happen [Dal14b]. happened [Qui12]. Happy [Jac14]. hard [TY17]. harmonic [CR15]. Harmonious [KN08]. Harper [Bla11a]. Harris [Fla18]. Harvard [Phi12b]. Haven [Dal12a]. Haynes [KSZ17]. headed [Mur12]. health [VJCC19]. hear [Sch11]. Hearing [GM13]. heart [Cai11, Kna14, Qua18]. heavens [Car13]. Helen [DL16b]. help [Sax18]. Helson [Ano11u]. Henri [SA10, Sti14]. Henry [Ano11u]. Henryk [Ano11i]. her [Pan18]. Herbert [Ano12h, Ano15b, Pap18]. here [HS10]. hereditarily [Giv19]. heroes [Sil10]. Hess [Lan18b]. Hidden [YN11, Ari15, Bla11b, Dum19e, GHL15, Kna14, Dum19e]. hiding [Rib19]. high [Don18]. high-dimensional [Don18]. higher [GJO19]. highly [Ald11]. Hilbert [Bae19]. Hilton [Ano11o]. him [Kha16]. Hirsch [Yon14]. Hirzebruch [Ano14b].

Hispanics [HLS19]. historic [Lip11, Pér11]. historical [Dia13]. histories [Lar18]. history [Ari15, BBE^{+}13, Bar13, BDJ^{+}17, BBH^{+}14, Bat14, Dal12b, Lin12, RT18b, Sna13, Wil18a, Dam10]. Hitchin [Keh16]. hitting [D’A10]. HIV [RTKB12]. HIV/AIDs [RTKB12]. Hochschild [Ano11k]. Hocking [Nai13]. hocus [Mag14]. hocus-pocus [Mag14]. Hodge [AHK17]. Hole [Tu15]. Holm [Ano18-42]. holonomy [Gro18]. Holtz [Ano18-38]. home [Aud11]. homogeneity [KS12a]. homology [Shl18, Wei11]. homotopy [APW13, Ber19a, Ger19]. honor [Wil18a]. honorem [BAD^{+}18, CMJ^{+}18, CST^{+}17]. Hopper [Aue19]. Hörmander [Ano15h]. Horne [Dum19e]. Horton [Mil18, Sch13a]. House [Ald11, DeT17]. housekeeper [Fuj10]. Houston [DL17b]. Houston-Edwards [DL17b]. http [Gre12b]. human [BRH10, Dal14a, KH19, Rob15]. Humanizing [Aus18]. hundred [AABT^{+}19, McC17]. hunt [BZ11, Ano18v]. Hunting [Rob19]. hurricane [May19b]. Huxley [Ano13a]. hydrodynamic [Li18]. hyperbolic [Ada18, CP10]. Hypergeometric [Beu14]. hypoellipticity [Ano18-46, Str18]. hypothesis [O’S18].

I. [Bei16]. ICCM [Ji16]. ice [BBF11]. idea [AABT^{+}19, Ran14, Ari13]. idealism [Hol13]. ideas [Ano16g, Bar10, Dal14a, Hub19]. ideology [Hol13]. II [Ano13c, Ano13i, Ano15g, Ano16d, Bar17, Dra15a, Dun19b, Gre12b, Oor13]. Illusion [Dal12b]. illustration [WS11]. Ilya [Ano10g]. imagination [Fre18]. Imaging [Bor19]. Immigration [BDJ^{+}17]. impact [Ald11, Ban17, LSW19, Avi14]. implementing [SZ13]. improbability [Dal14b]. improbable [Ald11]. improve [May19b]. Improvement [RTKB12]. IMU [CDW14]. inaccessible [BBMW13]. inclusive [Hos19]. incompleteness [KR10].

Incomputability [Coo12].

indecomposable [Giv19]. index [Yon14]. India [RTKB12, Mum10]. Indianapolis [Lip11]. individualism [Dum19b]. indivisible [BKK14]. industrial [Boh14]. industry [Don18, MOT10]. inequalities [Mór14, Pak19]. Infinitesimal [Ger16]. infinity [Far12a, Glu14, And16]. Influence [BF12b, Ano12d, DSG17]. Influential [PM12]. information [Chu10b, Nie18]. infra [Dek11]. infra-nilmanifold [Dek11]. Infrastructure [JS14]. Ingrid [Ano11i, Ano18y]. inner [Eis10, Hit11]. inspiration [CK19]. inspire [CMT17]. inspired [Hou18b]. instant [TW19]. instant-runoff [TW19]. Institute [CCBB^{+}16, BDJ^{+}17]. insulation [BBN17]. intangibles [Saa13]. integer [Slo18]. integral [Bal19]. interactions [Mil15]. Intercepts [Don14]. interdisciplinary [HH10]. intermediate [Bar13]. International [Rou19]. Interview [DL16c, DL16d, DS18, DS19, Oog15, RS10, RS11, RS12, RS13, RS14, RS16a, RS16b, RS17, Sch13a, Ano11p, DL16a, DL16b, DL16e, DL17a, DL17b, DL18a, DL18b, DL18c, Fla17, Fla18, Jac11, Jac13b, Jac13a, Lan17, Lan18b, Lan18c, Reh10]. intimidation [Ewi11b]. introduction [BS13, Boh14, Wil18a, Jon12]. Intuition [Mum11]. intuitionist [Joh14]. invariance [Tum16]. invariant [Sna13]. Invariants [Ger19, GJO19]. Inverse [Tre16]. Invisible [Lyn13]. invitation [DT10, GW19, Oko19]. iPad [Grä13b]. Iraq [Dam10]. Irene [Ano18z]. Isoperimetric [CFL^{+}12]. Israel [Ano13b, Ano13c]. Italian [Row19]. items [BKK14]. IV [Grä11]. Iwaniec [Ano11i, Ano15a]. Iwasawa [Sha19].

J [ABDS18, Ano11q, Ano12l, Boy13, Joh14, KK18, RS17]. J. [NB18]. Jack [MZ12]. Jackson [Ano18-43]. Jacob [ADF^{+}15]. Jacobian [Gui19]. Jacqueline [Pan18].

James [Sil14, Ano13d, Ano18-27, CR15]. Jane [Abo19]. János [Dén11]. Japanese [Don14]. Jerry [Ano12i]. Jewish [Sen13]. Jill [Ano18-28]. Joan [Ano17a, Ano18-29, Mar19]. Joe [Ano18-44, Fla18]. Johannes [Ano11q]. John [Ano11i, Mil18, RS11, Ano16f, Bat11, Nai13, RS12, RS16a, Sch13a]. Johnny [Dys13]. Johnson [Hou19b]. Joining [HS19]. joint [JP11]. Joram [Ano15d]. Joseph [Ano17c, CR15]. Journals [Olv11, CHKM11, Grc10]. journey [Bla11a, Goi18, San11, WS11]. Jr [ABDS18, Ano16f, Bat11, May19a, RS16a]. Jr. [NB18]. judging [Rib19]. July [Ano14e]. just [Smi14].

Kadison [GJRR19]. Kalman [Jac10]. Kalton [Ano12l]. Kan [BHMM15]. Kardar [Cor16]. Karen [Ano18-30, Ano19a, Don19, Fla17]. Katherine [Hou19b]. Kathryn [Lan18b]. Katok [Has19]. Keller [Ano17c]. Kelsey [DL17b]. Kervaire [Sna13]. kinetic [CP10]. Klaus [HSPM15]. Klein [Fra12]. knew [And16, Kha16, Rib19]. knight [CK19]. knot [Nel11, SW17]. Knots [Sil14]. knowledge [Fre18]. known [Car15]. Kobayashi [Ano14d]. Kontsevich [Ano12e]. Kunze [GW15]. Kurt [BBC15]. Ky [Lin10]. Kyoto [Ano10h].

L [Ano18i, Ano18-43]. L. [May19a, Joh14]. Lamb [DL17a]. lament [Sch13b]. Langlands [Ano18-39, DS19, Fri18]. Large [Nee19]. Lars [Ano15h]. last [Ono10]. Late [Cor13]. later [Mil11]. Lathisms [HLS19]. Latinxs [HLS19]. lattice [KŠŠ12, Zon13]. Laubenbacher [Lan17]. Laureate [DS18, DS19, RS11, RS17, RS16a]. Lauren [Ano18-31]. Laurie [Roc12]. Law [BH17]. Lawrence [SLL12]. laws [KS12a, Orn13]. learning [E19, dlP19]. least [Pap19]. Leavitt [Abr16]. Lecce [Bom16]. lecture [Fre18]. lectures [FF12]. led [Gla13]. Lee [Dum19e]. Leech [Zon13]. Leffler [Hel13a]. legacy [BBC15, KKK^{+}14]. legendary [Coo10]. Leibniz [BBH^{+}14, KS12a, Kno15]. Lelong [SDS14]. Leon [Ano18a, Ano11r]. Leonard [DSG17]. Leroy [Ano18-27, Ano18h]. Level [CM16, Row14]. Levi [Ano18i, Row19]. Levi-Civita [Row19]. Lewis [May19a]. LGBT [BBKL19]. Library [Sch13b]. Lie [Ano18d, BG11b]. Life [Coo10, Bar10, Bel16, Bla11b, Cor13, Dra15b, Hou19a, Hou19b, Jac16, KK18, McD17, NB18, Ada11, Joh14]. Lifetime [Ano18-27]. Light [Pet10, HN17]. Lillian [Ano18-32]. limit [BBH^{+}14]. limits [LS16, Tos18]. Lindenstrauss [Ano15d]. line [Hyn19, RTKB12, Slo18]. Linear [Li18, De 19, DG14, Gar10]. lineman [Mil16]. lines [Wic18]. Lion [Hay13]. liquid [CMT17]. literature [Dun19c, Grc13]. Liveright [Mil19]. lives [Zor12]. living [AdS19, RTKB12]. Localized [HS10]. Lofton [KSZ17]. Logic [Dra15b, Avi16, Gra13a]. logician [Dav13]. Logicomix [Roi10]. Loh [DL16e]. London [Fuj10, Zor12]. long [Car13]. Look [GB12]. Looking [Dun19c, Cor13]. lost [Avi16, Kob14]. Louis [Zit11, Don11, RS16b, Riv16]. Lovász [Ano10h]. Love [Kna14, Col16]. Ltd [Zor12]. Luce [Dum19a]. luck [Dal12b]. Lunch [Ser14]. Lusztig [Ano14c]. Lyme [Nee19].

M [BHMM15, Bei16, Loh19, GT15]. M. [Sch10]. MA [Car13, Phi12b]. Machine [E19, dlP19, Lip11]. machines [TY17]. Mackey [Dum19b]. magic [Per19b]. Magical [Van12]. magnet [BG11b]. Mahler [BBC15]. Mahowald [Rav16]. Maker [Mil19]. Making [DEG^{+}17, Lar18]. Malaysia [CGM^{+}15]. Malliavin [SYK^{+}11]. man [And16, Bla11b, Han12, Hel13a, Rob15]. Mandelbrot [Ano12c, Ano12d, BF12a, BF12b, Sil13]. manifold [Ada18, Kar11, Lár10b]. manifolds [Tos18]. Manin [Cor13]. mapping [Bed12]. mappings [Sob12]. Maps [Pap19, CR15]. Marc [LP14]. Marcel [LBG^{+}17]. Maria [Ano18-33, Dum19d]. Marina [Dan19]. Mark [May19a, Rav16]. Markov [D’A10, Gun10, Pet19]. Marques [DL16c]. Marsden [Ano12i]. Marshall [Don14]. Martin [Ano11l, Mon19]. martyrs [Sil10]. marvelous [Dar17]. Mary [Ano15e]. Maryam [Ano18-34]. master [YHH19]. Matchett [Ano18-36]. material [Rad13]. math [Bre16, Cai11, Dum19a, Kna14, Man15, Poo17, Sax18, Ver13, GK18, AK12, Ede13, Roi17]. Math. [Ewi11a, Por10]. Mathematical [Cur12, Dia13, Ewi11b, Fen12, Neu11, Rot17, Ano16c, Ano16d, Ari15, BBE^{+}13, Bar10, Bay12, Bea11, Bel16, BFH10, Cai11, CLWZ11, Coo10, Dal14a, Dal17, E19, Fre10a, Fri10, Ger16, Glu14, GK11, Gof15, HWWA11, HLS19, HS19, Jon19, Lar10a, Lar18, O’S11, RTKB12, Reh10, Sac11, San11, Sch10, SRM^{+}17, Van15, CCBB^{+}16, Mur12, Jab19, RR13, Ric14]. Mathematically [Buk18, GHPW18]. mathematician [Ano11o, CK19, DeT17, Ewi11a, Goi18, Hou19a, Hou19b, Jac17, Mil16, Nat14, Sch13b, Zit15]. Mathematicians [DPV14, Grc11, BBKL19, Hou18b, Jon16, PM12, RT18a, Row19, Sen13, Aud10, Eis10, Kob14]. Mathematics [BF12b, BJ12, CHKM11, Dum19c, Nic12, Oss10, Pap17, Zel19, Ada11, AABT^{+}19, Ano12d, Ano18e, Ano18g, Ano18-35, Ati10, Aue19, Aus18, Avi11, Avi18, Bae19, BBdPZ14, BBLZ14, Ban17, BGL^{+}18, Bar17, Bat11, Ben10, Ber19b, BGY17, Boh14, BME^{+}18, Bur14, Col16, Dal12b, Dud10, Dum19b, Dun19c, FF10, Flø15, Fol11, Gla13, Gof15, GB13, Grc10, Grc13, Ham19, Hay17, Hel13b, HH10, Hod11, Hol13, Hos19, Jac16, Ji16, Joh14, KK18, Keh18, KvP16, LS16, Lin12, Lyn13, Mar19, MOT10, Mil15, NPS14, Phi12b, Qua18, Qui12, Rad17, Sau16, Tho14, VJCC19, Wal15, Wal10, Zit11, Zor12, Rou19, Bau15, Dam10, Jon16, McC18, Mum10, Pen11, Arc13, Cor13, Dav11, Des11, Hel13b, Man10, Mon13, Mum18, Que19, Sil10, Ste14, Van12, Zor12]. MathSciNet [Dun19a, Dun19b]. matrices [DEG^{+}17, KS12b]. matrix [AdS19, Lan18a]. matroidal [SP18]. matroids [AHK17, Ard18]. matter [Fol12b]. matters [Qui12]. Max [Pei17]. maxima [Mór14]. Maxwell [Sil14]. May [AABT^{+}19, CHKM11]. McClelland [Zor12]. McDuff [Ano18q]. me [Hou18b]. mean [CM16]. Meaning [Hel13b]. measure [Fur11, Tor19]. measurement [Saa13]. mechanization [Avi18]. Medal [Jac10, Keh11, Keh12i, Bar15, Ano10b, Gra16, Kol15]. medalists [HWWA11, dMPQZ15]. Medals [Ano10f]. media [Bor19, Hay14]. medieval [Sac11, SS15b]. Meeting [Hon16]. Meijer [BS13]. Melanie [Ano18-36]. Melody [Ano18-37]. memoir [Bon13]. Memorial [May19a, Dan19]. memoriam [Abo19, Ano10g, Ano15d, CJMR16, GJRR19]. Memories [Abb19, Ano11l, Ano12f, Ano15e, Ano19b]. memory [ADF^{+}15, Ano15c, Ano15h, AGLP^{+}13, GW15, SDS14]. men [Was18]. menagerie [GHL15]. Mental [VJCC19]. Merrill [KR12]. metagenomics [AJ19]. method [CM16, GDG^{+}15, Ran14]. methods [Byr10, Dia13, MS19, Sto19]. Metric [Tos18]. metrics [Tum16]. metronome [FGLG13]. Meyer [CMJ^{+}18, DS18]. MH370 [CGM^{+}15]. Michael [Ano19b, CK19, CK19]. Michel [Abb19, Ill19]. Michèle [Jac17]. Michell [Gan18]. microbial [AJ19]. middle [Bar13, SZ13]. Mikhail [RS10]. Milella [Bom16]. million [Roc12]. Mills [Mar12].

Milnor [Ano11i, Ano11m, RS12]. mind [Mil18, Sau16]. minimal [Pér17]. mining [Mag12]. miracles [Dal14b]. mirror [Per19b]. Mirzakhani [Ano18-34]. Misfortunes [DPV14]. misguided [Ewi12]. misuse [Tho14]. Mittag [Hel13a]. Mittag-Leffler [Hel13a]. mock [Duk14, Fol10]. Model [Moo10, RTKB12]. Modeling [Ber19a, CCBB^{+}16, Gar19, Ano18j, BFH10, Buk18, Dur13, Mur12]. Modelling [San11]. models [DRAS16]. modern [Ban17, Boh14, Ger16, KvP16, Sil10, Zit15]. modernist [Man10]. modular [Fol10]. Moduli [Kov18, Mil15]. Moiseevich [Ano13b, Ano13c]. molecules [CLWZ11]. Moment [Mór14, Lin10]. momentum [Rao14]. monsters [Mur12]. month [RT18b]. Moore [DL16b]. Morawetz [Ano18-35, Jac14]. Morel [Ano18-41]. Morgan [Ano11f, Ano18e, Keh12d]. Moscow [Nat14]. motion [CM16, Way11]. Mountain [Pei17]. movie [And16, Sau16]. MR1970295 [FF12]. MR1976398 [FF12]. MR1996199 [Dav11]. MR2129625 [FF12]. MR2265113 [Lar10a]. MR2363580 [Avi16]. MR2426819 [Ben10]. MR2435289 [Avi11]. MR2440945 [Mar10]. MR2440977 [Dam10]. MR2452344 [Man10]. MR2468443 [Mum10]. MR2473212 [Pen11]. MR2522825 [Aud10]. MR2526973 [Glu14]. MR2536166 [Bra10]. MR2542778 [Eis10]. MR2583232 [Des11]. MR2590059 [Häg12]. MR2590934 [Bay12]. MR2598223 [O’S11]. MR2604887 [Hod11]. MR2641048 [Roc11]. MR2654369 [Sil10]. MR2674693 [Fol12b]. MR2722198 [Hit11]. MR2722967 [Mon13]. MR2723248 [Bea11]. MR2743189 [Hel13a]. MR2759991 [Ada11]. MR2777703 [Fol11]. MR2807493 [Arc13]. MR2807927 [Bla11b]. MR2827930 [FF12].

MR2848884 [Jon12]. MR2858033 [Van12]. MR2866515 [LS16]. MR2882171 [Cur12]. MR2884886 [Roi10]. MR2895623 [Hel13b]. MR2896048 [Fol12a]. MR2905809 [Joh14]. MR2906064 [Mag13]. MR2908506 [Fol12a]. MR2919681 [Smi14]. MR2919869 [Lin12]. MR2954056 [Dav13]. MR2963548 [Smi14]. MR2963841 [Häg13]. MR2976816 [Dal13a]. MR2986502 [Sti14]. MR2986979 [RS13]. MR2994999 [Ari13]. MR3013691 [Gra13a]. MR3024839 [Lyn13]. MR3025050 [Ede13]. MR3058839 [Wig14]. MR3087339 [Sen13]. MR3100544 [Bat14]. MR3114790 [RS14]. MR3136845 [Dal17]. MR3137670 [Ste14]. MR3155039 [Gof15]. MR3155773 [Kna14]. MR3183832 [GT15]. MR3183982 [Ari15]. MR3185259 [Smi14]. MR3222396 [Bur14]. MR3234572 [Dal14b]. MR3236985 [Blu16]. MR3243751 [SS15c]. MR3287223 [Fol16]. MR3289987 [McC18]. MR3329687 [Mil18]. MR3330133 [Kno15]. MR3362188 [Wal15]. MR3363668 [Mar16b]. MR3364090 [Ros19]. MR3364494 [Ros19]. MR3380969 [Gra17]. MR3408723 [Avi16]. MR3408971 [Ros19]. MR3409067 [RS16b]. MR3409221 [RS16a]. MR3410783 [Ros19]. MR3443339 [Ros19]. MR3444773 [Hay17]. MR3486163 [Jon16]. MR3495254 [Kob16]. MR3558830 [Roi17]. MR3561047 [Roi17]. MR3561130 [Poo17]. MR3585691 [Lar18]. MR3616260 [O’S18]. MR3618873 [KH19]. MR3674587 [Axl17]. MR3702017 [Hub19]. MR3702027 [Jon19]. MR3726777 [DS18]. MR3729276 [Aus18]. MR3729321 [Mum18]. MR3753635 [Elr19]. MR3775369 [Que19]. MR3791446 [Gol19]. MR3793646 [Bae19]. MR3821554 [Row19]. MRI [Don18]. MTBI [CCBB^{+}16]. Mu [KW17]. Mu-Fa [KW17]. much [Fri15]. multilayer [Por18]. multiple [FH11, MFV16]. multiplication [Lan18a]. Mumford [Oor13, Keh11, Keh13b]. Murray [Aue19]. muse [Phi12b]. Music [DMVW10, KK18]. My [McD17, Wei18, Dra15b, Kob16]. Mysteries [LZ12, Van15]. mystery [SS15b, Fre10a]. mystic [Bla11b]. mysticism [Glu14]. myth [Bar15, Bre16].

Nahmod [Ano18m]. name [Car15]. Naming [Glu14]. Nash [Ano16f, RS16a, DL19, SW16]. National [Jac10, Keh11, Keh12i]. natural [Cho11]. naturally [Mon19]. nature [Fol12b, Hel13b, Sau14, Bla11a]. Naval [Don14]. Nazi [Aud10]. near [PN12]. need [Mer19b]. needed [Fri15]. needs [CDW14]. Nefarious [AF11]. Nelson [Ano18f]. Nemmers [Ano12b]. network [Por18]. Neumann [Dys13]. Nevanlinna [Ano10d]. Newton [Orn13]. NFL [Mil16]. Nicole [SS15b]. Nigel [Ano12l]. nilmanifold [Dek11]. Nirenberg [Ano16g, Don11, RS16b, Riv16]. NIST [Hal14]. NJ [Dal12b, Dal14a, Ewi12, Fre10a, Mag14]. NM [Hay13]. no [Mer19a, Mur12, Mer19a, Ser14]. non [Mon13]. non-European [Mon13]. noncommutative [Sch18b]. nonlinear [Byr10]. nonlocally [CMT17]. Nonsense [Häg11]. nonstandard [GW19]. normalization [dlP19]. Northwestern [Aud11]. notation [Ari15]. nothing [Ste11]. notions [Flø15]. November [Ano14e]. NSA [Hal14]. null [BR18]. number [Ano11c, Bon12, GM19]. numbered [Zor12]. Numbers [Shp10, AF11, Fri15, Han12, O’S18, Hod11, Jon12, Law19, Mag13, Spe10a]. numerate [Dam10]. Numeric [War13]. numerical [OSS19]. numerics [vdBL15]. nursery [Sil14]. NY [Han12, Str15].

Oberwolfach [KvP16]. objects [HO19]. occasion [Zie18]. occurring [Mon19]. octopus [Con12]. oeuvre [Pan18]. offensive [Mil16]. Oka [Lár10b]. old [Car13, CMT17, Ran14]. Olek [Ano17b]. Olga [Ano18-38]. on-line [Slo18]. One [Mer19b, Roc12, AABT^{+}19, Bae19, Goi18, McC17, Mil16, Sch11, McC17]. Only [KH19, Car15]. Onsager [DL19]. ook [GT15]. Open [Gan18, BBS13, KW18, Rad17]. opened [KR12]. OpenFOAM [CXM^{+}14]. operator [Gar17, Vid19]. operators [DG14]. Opinion [BBS13]. opportunities [Dum19a, Mol10]. Optics [Per00, Gui19]. optimization [BBN17, De 19]. Optimizing [Ser14]. order [BDG16]. orders [Zor12]. O’Reilly [Hay14]. Oresme [SS15b]. Organic [BJ12]. organization [BBKL19]. organize [Jon16]. Origami [DMP10]. Origins [Bla14a, Bar17, Bla14b, BBKL19, Dal17]. orthogonal [MMB17, MFV16, Mór14]. oscillations [Byr10]. Ostrowski [Keh10]. other [Avi11, Bre16]. our [Jon16, Zor12, Zor12]. out-of-sample [BBdPZ14, BBLZ14]. outer [Eis10]. outstanding [Ano18e]. overfitting [BBdPZ14, BBLZ14]. Oxford [Smi14].

P [Ano15j, Ano16h, Ano18-27, Ano18h, DS19]. Packing [Coh17, LZ12, Chu10a]. pages [Kuz15]. pair [Tef10]. paired [Saa13]. Palis [Ano11n]. Pamela [Spe10a]. paper [Lip11]. papers [Fol16, Oor13]. parabolic [CP10]. paradox [KR10]. Paradoxes [Häg13]. parallel [OH13]. parallel-transport-based [OH13]. paraproduct [BMN10]. Paris [Hur14]. Parisi [Cor16]. Parker [Bat11]. Part [Ano13c, Ano12j, Ano13i, Ano15g, Ano16a, Ano16b, Dun19b]. partial [DMP10, Mór14, Riv16, Shu12]. particle [Hyn19]. partisan [BD17]. partitionability [DKM17]. Path [KN08, Abr16]. patterns [Bla11a, Mil19, Wal15]. Paul [Bla11b, Ewi11a, ABDS18, Ano15f, Ano15g, KLN^{+}14, Sar10, SYK^{+}11]. Paying [CR15]. Payne [SLL12]. peacock [Mon13]. Pearl [Ano12g]. peg [Mat14]. Pelczynski [Ano17b]. Penguia [Far12a]. Penguin [DeT17, Man15]. Penrose [BR18]. pentagonal [CFL^{+}12]. people [RTKB12]. perfection [KSZ17]. perfectoid [Bha14]. performance [BBdPZ14, BBLZ14]. Perimeter [Adl12, Ano17d]. period [MS14]. persistent [Wei11]. person [BKK14]. Personal [Wil18b, Dur13, Lar10a]. Perspective [Nic12, CLWZ11, Dur13, Lar10a]. perspectives [Bae19]. perverse [dCM10]. Peter [Ano11o, BCCTS19]. Peters [HSPM15]. Phase [WS11]. PhD [Sch18a]. philosopher [Joh14]. philosophies [Lar18]. philosophy [Bay12, Bur14, Dav11, Hos19, Lar18, Aus18, Avi11]. physics [Ano16c, Ano16d, Bae19]. Pi [BB14, Bor12]. Piatetski [Ano10g]. Piatetski-Shapiro [Ano10g]. Picadoz [Fuj10]. Pierce [Ano18-32]. Pierre [RS14, SDS14]. pioneering [CR15, Hou19b]. pioneers [Hou18b]. Pipher [Ano18-28]. pitfalls [Rob10]. place [Dun19a, Dun19b]. plan [Mil16]. planar [Shl18]. plane [Zie11]. planet [Jon16, Jon16]. Plato [Hod11, Man10]. Play [Shp10, Mil18]. plots [WS11]. Po-Shen [DL16e]. pocus [Mag14]. Poincaré [Sti14]. Point [Law19, Ran14, Tu15]. points [Zie11]. polar [DEG^{+}17]. policies [Ewi12]. pollution [CCF^{+}12]. polyhedra [Gho18]. polynomial [MMB17, MFV16]. polytopes [DHV11, Zie18]. Popa [Keh10]. popular [Gof15]. Population [Rao14]. possessing [KŠŠ12]. potential [KL14, SZ13]. power [D’A10, Law19]. powers [Ari15]. practice [Lar18, Bay12, Lar18]. practitioners [Dum19b]. precedent [BB14]. Predicting [May19b]. prediction [HO19]. Prefect [O’S11]. presence [Bei16]. present [Hod11]. President [Jac16]. Presidential [Ano11p, Jac11, Jac13b, Jac13a]. Press [Aud11, Bre16, Car13, Dal12a, Dal12b, Ewi12, Fre10a, Häg11, Hay13, Mag14, Phi12b, Sch13b, Smi14, Dra15b]. Pressman [And16]. prevent [CCF^{+}12]. previously [BBMW13]. Prime [Fri15, Spe10a, O’S18]. primer [Gui19]. Princeton [Dal12b, Ewi12, Fre10a, Mag14, Dal12b, Ewi12, Fre10a, Mag14, RT18a, FF12]. principal [Saa13]. Principle [Per00, Dal14b, FH11]. Principles [SZ13]. Prize [Ano18-27, Ano10e, Ano10c, Ano10h, Ano10d, Ano11b, Ano11j, Ano11a, Ano11c, Ano11d, Ano11e, Ano11f, Ano11m, Ano11n, Ano11g, Ano12b, Ano12e, Ano12k, Ano14a, Ano14c, Ano15a, Ano18a, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18i, Ano18-39, Ano18j, Ano19a, GM19, Keh10, Keh12a, Keh12h, Keh12c, Keh12b, Keh12d, Keh12f, Keh13a, Keh16, Keh18]. Prizes [Ano11i, Dum19c, Ano10a, Ano11h, Ano18h, Keh12e]. probability [CC10, Häg13, Mór14, Dal13a]. Probably [Fel14]. problem [Ban17, CMT17, Gho18, Mar16a, Mat14, Sna13, Yon14]. problems [BFH10, BBN17, Dal13a, Gan18, KW18, Lar10a, San11, Sto19, YHH19]. process [Gun10]. Productions [And16]. professional [Rob10]. professor [Ver13, AGLP^{+}13, Van15, Fuj10]. Profiling [AJ19]. program [Fri18]. progress [Sol18]. projection [Nie18]. promenade [Jon19]. promotion [Str17]. Proof [BR18, Cho11, Dar17, Ede13, Avi11]. proofs [Bou13]. propositions [GDG^{+}15]. protect [Jon16]. protection [Rob10]. prove [BG11a, GTTZ10]. proven [Mer19a]. Pseudo [BBdPZ14, BBLZ14]. Pseudo-mathematics [BBdPZ14, BBLZ14]. pseudoconvex [Ran12]. psychology [Dal12b]. Publ [Dal14a]. public [Ano18b, Fre18, Keh12g].

Publishing [Law19, Zor12]. puro [Bom16]. pursuit [Fol12a, LS16]. Putting [Ada11]. puzzle [Kal16]. Puzzling [SS15a]. Pythagora [Fre10a].

quadratic [JS14]. quandle [Nel16]. quantification [Dam10]. Quants [Ste11]. quantum [Ber19b, Kal16, War13, Zel19, Wig14]. quantum-enabled [Ber19b]. Quercus [Zor12]. quest [BG11a, Daw18, Mum11]. question [Duk14]. questions [Goi18, Hel13b, Zor12]. Quillen [Ano12a].

R [And16]. Race [Law19]. Raggi [AGLP^{+}13]. Rallis [Ano13g]. Ramanujan [Ano12j, Ano13i, Bor12, Sau14, Kob16, Sau14]. Random [DRAS16, DeT17, Bor19, Ald11]. randomness [DH19]. ranking [Dal13b, HO19]. Raoul [Ano13f]. rare [Dal14b]. rating [Dal13b]. Ratner [Dan19, LSW19]. Rauzy [AH14]. Ray [Per00, GW15, Rob15]. Raynaud [Abb19, Ill19]. readers [Edw12]. Real [Dén11, Dav11, EMS11, Gra17, Hyn19, JS14]. reality [Kna14]. really [Qui12]. Recalling [Ano13d]. Receive [Keh13b, Keh12i]. receives [Ano10e, Ano10h, Ano11m, Ano11v, Ano12g, Ano12k, Jac10, Keh10, Keh11]. reckoning [Lin12]. rectifiable [Ano18-45]. reduction [For18]. reflect [Jon16]. reflection [Hou19a, NPS14]. Reflections [Ano15f, Ano15g, Fol12b]. Regular [LZ12]. regularity [JP11]. Reich [SS15c]. Reid [Ale11]. Reinhard [Lan17]. relative [Saa13]. Relativity [Per00, Row19]. religious [Glu14]. remembered [LBG^{+}17]. Remembering [Ale11, Ano11q, Ano11r, Ano12h, Ano12i, Ano13e, Ano13f, Ano13g, Ano13h, Ano14d, GCCUD18, HSPM15, Rav16, Sar10, SYK^{+}11]. remembrances [BSG^{+}10, Roc12]. Reminiscences [IBB^{+}10]. Renaissance [Phi12b]. report [And16, Sol18]. representations [Sch18a]. reprint [DS18, RS13, RS14, RS16a, RS16b]. requirement [RTKB12, Rob10]. research [Ano18e, CMT17, HS19, San11]. resilience [May19b]. resonance [Str17]. resplendent [Kos11]. restriction [Sto19]. resurfaces [CMT17]. revenge [Fre10a]. Reverse [Mum18]. Review [Bla14a, AK12, Ada11, Ald11, And16, Arc13, Ari13, Ari15, Aud10, Aud11, Aus18, Avi11, Avi14, Axl17, Bae19, Bat14, Bau15, Bay12, Bea11, Ben10, Bla11a, Bla11b, Bla14b, Blu16, Bom16, Bra10, Bre16, Bur14, Car13, Col16, Cor13, Cur12, Dal12a, Dal12b, Dal13a, Dal14a, Dal14b, Dal17, Dam10, Dav11, Dav13, Daw18, DeT17, Des11, Dra15b, Ede13, Eis10, Elr19, Ewi11a, Ewi12, Far12a, Fel14, Fol11, Fol12a, Fol12b, Fol16, Fre10a, Fuj10, Ger16, Glu14, Gof15, Gol19, Gra13a, Gra17, Gre12b, GT15, Häg11, Häg12, Häg13, Han12, Hay13, Hay14, Hay17, Hel13a, Hel13b, Hit11, Hod11, Hub19, Hur14, Joh14, Jon12, Jon16, Jon19, KH19, Kna14, Kno15, Kob16, Lar10a, Lar18, Law19, LS16, Lin12, Lip11, Lyn13, Mag13, Mag14, Man15, Man10, Mar10]. review [Mar16b, McC17, McC18, Mil18, Mil19, Mon13, Mum10, Mum18, O’S11, O’S18, Oor13, Pen11, Phi12b, Poo17, Por10, Que19, Roc11, Roi10, Roi17, Ros19, Row19, Sau14, Sau16, Sch13b, Sen13, SS15c, Sil10, Smi14, Spe10a, Ste11, Ste14, Sti14, Str15, Van12, Van15, Wal15, Wig14, Zor12]. reviewing [RR13]. reviews [Avi16, FF12, Smi14, RR13, Ric14]. Reviews(R) [Jab19]. revolution [Nel11, Qui12, Han12]. revolutionary [Rad13]. Ricci [Row19]. Richard [BME^{+}18, GJRR19]. Riehl [Ano18r]. Riemann [O’S18, DT10]. Riemannian [DT10, Gro18, Tum16]. right [Dun19a, Dun19b]. rights [BRH10]. rigidity [SS17]. rigor [Bar13, Mum11, O’S11]. Rigorous [vdBL15]. rings [Ger19]. rise [Sil10]. Rittenhouse [Zit15]. Robert [DS19, Jac16, Ano18-39]. Rockburne [Pei17]. Roesel [KLN^{+}14]. role [Lyn13, NPS14]. Ronald [YHH19]. rooted [Joh14]. roots [GK18, Mon13]. Rosemary [Ano18-40]. Rosies [Gre12b]. Rota [GGW14]. rowmotion [Str17]. Rubik [PQ13]. Rudin [Ano13h, Ano15e]. Rudolph [Boy13]. Rudy [Dum19e]. rugged [Dum19b]. rule [Hod11]. runoff [TW19]. Russell [Ano18c]. Ruth [DL18b].

S [Ano15b, Ano18-42]. s.r.l [Bom16]. Saharan [BGL^{+}18]. salesman [LS16]. Sally [ABDS18]. sample [BBdPZ14, BBLZ14]. sampling [FH11]. Sampson [CR15]. Samuel [Sau16]. sandpile [LP10]. sandwiches [Sob17]. Sandy [Kha16]. Santa [Hay13]. Sarason [GCCUD18]. Satter [Ano11g]. say [Ser14]. scaling [Gun10]. Scanlon [Abo19]. Schafer [BSG^{+}10]. scheduling [HiK13]. scheme [Wic16]. Schneider [Bru15]. Schoen [BME^{+}18]. school [IBB^{+}10]. Schur [Mor18]. Schuster [Col16]. Schwartz [ADF^{+}15]. Sci [Dal14a, Avi14]. Sci. [Ger16]. science [Dal13b, Dum19a, Häg11, Jac10, Keh11, Keh12i, Daw18, Fol12b, Hay14]. sciences [HLS19, Tho14]. scientific [E19, FH11, MOT10, Sti14]. scissors [Dup12]. seagull [Chu10a]. search [Car13, CM19, HN17, Kob16, Roi10]. searching [Dun19a, Dun19b]. second [KR10]. secret [Gre12b]. secrets [Gof15, Car13]. Seduced [Gra13a]. see [HS10]. Seized [Mol10]. selected [Fol16, Oor13]. selection [Sto19]. self [dlP19]. self-normalization [dlP19]. selfishness [Häg12]. seminar [Bei16]. sense [Lar18]. separable [Ros11]. sequence [Row15]. sequences [Slo18]. series [BBF11, D’A10, Mór14, Per19a]. Serrin [Ano13d]. service [Ano18b, Keh12g]. Set [BBS13, Ano18-45, CM16, Sil13]. sets [PN12, Row14]. Severi [GB12]. Severo [Ver13]. shadow [BBH^{+}14]. shadows [Lin12]. shape [GM13, Hit11]. shaped [Ger16]. Shapiro [Ano10g]. Sharing [Ham19]. Sharon [Lan18c]. Shaw [Ano10e, Ano11j, Ano12e, Ano14a, Ano14c, Ano15a, Keh16]. sheaf [dCM10]. sheaves [Rob19]. Shen [Ano11s, DL16e]. Shiing [Ano11s]. Shiing-Shen [Ano11s]. Shimura [Mil12]. shiva [DeT17]. short [Ari15, Jon12]. Shoshichi [Ano14d]. shotgun [AJ19]. shows [Mil16]. Shreeram [Ano14e]. SIAM [Keh12a]. side [Sch10]. Siegel [Mar16a, Pér11]. sieving [RSW14]. sights [Per19b]. Simon [Ano16c, Ano16d, Col16]. Simone [Aud11]. simple [Sol18]. simplified [Ano16e]. Simpsons [Gof15]. since [Wig14]. singular [DEG^{+}17, Jon19]. singularities [Kov18]. Sir [RS17, Ano19b, CK19, DL16d]. six [Mil11]. Smiley [Abo19]. Smith [Ano18-30, Fla17, GM19]. smoke [MOT10]. smooth [Wic18]. Snapper [Buc13]. Snapshots [BGL^{+}18, KvP16]. Snell [Roc12]. Sobolev [MMB17]. social [Dam10, Tho14]. solitude [Spe10a]. solution [CP10, SZ13]. solutions [Shu12]. Solving [GGW14, Shu12, Lar10a]. Some [KW18]. something [FGLG13]. Song [Sil14]. Sophie [Ano18-41]. sound [Sch11]. sourcebook [Pen11]. Sources [Arc13]. Souza [Ver13]. Soviet [Hol13, Nat14]. space [Bha14, Giv19, Hit11]. spaces [Ros11]. spark [Mer19b]. Spazio [Bom16]. speaking [Sen13]. Special [BCLO11, Beu14, KvP16]. spectra [BBKL19]. spectrahedron [Vin14]. spectral [DG14]. Sphere [Coh17, Fox12, OH13]. spheres [Sto19]. Spielman [GK11]. spirito [Bom16]. sports [HiK13]. spring [OH13]. Springer [Oor13]. square [Mat14]. Srinivasa [Sau14, Ano12j, Ano13i, Bor12]. St [Zit11]. stability [Li18, Rao14]. stable [CM19]. stack [Vez11]. Staffilani [Ano18w]. stationary [Fur11]. Statistical [OSS19, Loh19]. statistics [ACP18, Por10]. Steele [Ano11i, Ano11h, Ano18-27, Ano18h, Keh12e]. Stein [DL18c, PR12]. Steinberger [May19a]. Steklov [KKK^{+}14]. STEM [Bre16]. Stephens [Hou18a]. Steve [Ano13g]. Stewart [Zor12, Van15]. Sticky [Hyn19]. stilts [Häg11]. stochastic [Ano18j]. stone [FF10]. storm [Lyn13, May19b]. story [Glu14]. strangest [Bla11b]. Strategy [And16]. Straus [Ger16, Str15]. Street [Des11]. Street-fighting [Des11]. striking [FH11]. String [YN11]. Strong [Ano15c, Ano12j, Ano13i, Bor12]. structure [JS14]. structures [CM19]. struggle [Hol13]. Stuck [Bar13]. stucture [Kos11]. student [Ano18e]. study [Dal14a, Dia13, Pap17, Sch18b, BDJ^{+}17]. Sturm [Cal19a]. style [Cor13]. sub [DT10, BGL^{+}18]. sub-Riemannian [DT10]. Sub-Saharan [BGL^{+}18]. subfactors [Shl18]. subjects [Hay17]. submanifold [HR17]. successful [Was18]. suggests [Ran14]. sum [Bat14]. sums [Mór14]. supercharacters [GHL15]. support [Dum19a]. Supreme [Sax18]. surface [Wic18]. surfaces [DHV11, Pap19, Pér17]. surge [May19b]. surprise [KR10]. survey [Mat14]. survival [Bar10, RTKB12]. Swan [Ald11]. Swinnerton [BCCTS19]. Swinnerton-Dyer [BCCTS19]. symbols [Ari15]. symmetric [Gar17]. symmetries [Far12b]. symmetry [DMVW10, KŠŠ12, Per19b, Wal15, Bla11a]. symplectic [Hol16]. Synge [Ano18-35, Jac14]. Systems [DPV14]. Szemerédi [Ano12k, RS13].

T [GB13]. tail [Phi12a]. Taking [Cai11, Qua18]. tale [KL14]. Tanaka [EMS11]. tangents [Ran14]. Tao [Keh12h]. Tapas [ACP18]. Tapia [Keh12i]. Tara [Ano18-42]. Tardos [Ano18s]. Tate [RS11]. teach [Por10]. teacher [CDD^{+}10]. teaching [Ben10]. tell [Häg11]. telling [Rib19]. Ten [Hou18b, Hub19]. tensegrity [Con13]. Tensors [Stu16]. tessellations [Elr19]. Tetrahedra [LZ12]. th [GT15]. their [BGY17, CR15, Coo10, Gof15, Kob14, Stu16]. Them [DPV14, HS10]. theme [Fra12]. Theorem [KN08, BG11a, Bar13, BBZ16, GTTZ10, KR10, Sch11, Tu15, Ser14, Str15]. Théorème [Hur14]. theorems [BT13]. Theoretical [CCBB^{+}16]. theories [Ber19a]. Theory [YN11, AHK17, Ano11c, Ano18j, APW13, Chu10b, Dal12a, DG14, E19, Ger16, GM19, HiK13, KL14, Kov18, Kuz15, Moo10, Mór14, Nel11, Ros11, Sha19, Shl18, SS17, Tor19, Vid19, Zak19, Ano18d, Häg13]. there [Bur14, DG14, HS10, Kol15, KŠŠ12, Mur12, Was18]. thermal [BBN17]. theta [Duk14]. thin [KLLR19]. Third [SS15c]. Thompson [CF11]. thought [Dal14a, Hay13]. Thoughts [Sau16]. Three [Goi18, Mar16a]. Thurston [Ano15j, Ano16h]. Tilings [Gun10, CFL^{+}12, Elr19]. time [TY17]. Times [Ste14, Per19a]. Timothy [DL16d]. today [CMT17, Qui12]. tools [De 19, Ben10]. Top [Gre12b]. Topical [Grc10]. Topological [Byr10, GJO19, Per19a, De 19, Zak19]. topologist [Joh14]. topology [DEG^{+}17, Mil11, Sch18b, Sob17]. Toronto [Zor12]. tossing [Orn13]. totient [Mer19a]. tour [Lip11, Roc11]. tower [Sha19]. town [Fen12]. Trachette [Ano18-43]. trademarking [BB14]. tradition [Sen13]. train [Cal19a]. transcendental [Bau15]. transcending [YHH19, Sen13]. transform [Far12a]. transformation [Man10, Sil16]. transformations [BV11]. transforming [Don18]. transport [OH13]. traveling [LS16]. treatment [Mer19b]. Trial [Ede13]. triangle [GM13]. triangles [KvP16]. Tribute [Ano12m, Ano18-44, SLL12, Ano11t, Ano11u, Ano12l, Ano14f, Ano15i, CR15, CK19, Dan19, May19a, SA10]. trinomial [BZ11]. Trio [DPV14]. trip [Ros14]. tropical [Kat17]. troubadour [Sac11]. troubling [BB14]. true [BBH^{+}14, Glu14]. trump [Tho14]. trusses [Gan18]. Trust [DPV14]. truth [Roi10]. tubes [Sto19]. tuple [Duj16]. turbulence [DL19]. turbulent [MOT10]. Turing [Ano11v, Ano12g, Avi14, KH19, Lip11, Smi14, Coo12, Don14, Mon19, Bla14a, Bla14b]. Turner [BSG^{+}10]. Tverberg [BBZ16]. twenty [McC17]. twenty-one [McC17]. Twists [Elr19]. Two [BKK14, BBN17, Kob14, Gla13, Roi17, Way11]. two-dimensional [Way11]. Two-person [BKK14]. TX [AK12]. type [APW13]. typical [Row14].

ubiquity [Goi19]. Uffe [CJMR16]. Uhlenbeck [Ano19a, Don19]. Ulf [Ano18j]. uncertainty [BFH10]. uncover [Ewi12]. Undecide [GS10]. undergraduate [Ano18e, BV11]. understanding [Lyn13]. Undiluted [Mag14]. Uneducated [Ewi12]. unexpected [Dum19a]. unfolding [Gho18]. Union [Nat14]. unipotent [LSW19]. Univ [Aud11, Dal12a, Dal12b, Ewi12, Mag14, Phi12b]. univalence [APW13]. universality [Cor16]. Universe [YN11, Fol12a, Bla14a, Bla14b]. University [Fre10a, Smi14, Häg11]. unknown [Riv16, Fol12a]. Unreasonable [Nic12]. update [Per19b]. used [Por10]. useful [Fri10]. Using [DPV14, Ewi12].

V [GJRR19, Grä13b]. Valiant [Ano11v]. value [Bar13, DEG^{+}17, Ste11]. valued [CHKM11]. values [KvP16]. Varadhan [Keh12i]. variations [Don19, Fra12]. variety [Mil12]. varifold [Men17]. various [Hay17]. versions [Bea11]. very [Jon12]. vexing [Hod11]. via [Ger19]. victors [BBE^{+}13]. Vienna [Daw18]. view [Eis10]. views [Ano11p, Jac11, Jac13b, Jac13a, Jac16]. Viking [Far12a]. Vintage [Bla14b]. virtual [SW17]. Visible [Cas10]. visual [MOT10]. vivant [Hur14]. Vivo [Bom16]. Vladimir [Ano12f, Ano12m, KKK^{+}14, Wei19]. Voevodsky [APW13, Wei19]. Vogan [Jac13a]. Voisin [BAD^{+}18]. volume [Oor13]. Vortices [Way11]. vs [JP11].

W [Ano11i]. Waerden [SS15c]. waking [Lin10]. Walk [Dys13, Ano16c, Ano16d]. Walker [Han12]. wallpaper [Wal15]. Walter [Ano13h]. Walton [Ano18p, DL18a]. want [Ewi11a]. War [Bar17, Gre12b]. Warga [MZ12]. warping [Pap19]. wars [GB13]. Was [FGLG13, HN17]. Washington [Ewi11a, Por10]. water [CGM^{+}15, Kuz15]. wave [Ber19b, HN17]. wavelet [BV11]. Wavelets [LWW13]. Waves [Sto19, BGY17, Bor19, Kuz15, MOT10]. wealth [Mag13]. Weapons [Poo17]. weather [Lyn13]. weatherquakes [Wal10]. website [GHPW18, Gre12b]. Weil [Wei18, Aud11]. Weinberger [Ano18x]. welcome [Lar18]. Wentsun [Wu17]. westward [Zit11]. Whitehead [Cal19b]. Whiteman [Ano18a, Keh12f]. Who [Dal13b, Gra16, And16, Gla13, Hou18b, Kob14]. wide [DSG17]. Wightman [Ano15c]. Wigner [Nic12]. Wild [Cal19b]. Wiles [Dar17, RS17, SRM^{+}17]. Wiley [Lip11]. Wilf [Ano15b, Tef10]. Wilkins [NB18]. Wilkinson [Ano18l]. will [GTTZ10]. William [Ano15i, Ano15j, Ano16h, DL18c]. Williams [Ano18-31]. Winifred [KR12]. wins [GM19]. wireless [Set10]. without [McC18, Ser14]. Witten [Oog15]. Wolf [Keh13a, Keh18]. woman [Was18]. Women [RT18b, AABT^{+}19, Dum19a, RT18a, Que19]. won [Gra16]. Wood [Ano18-36]. Woods [Tu15]. word [San11]. words [Ono10, Dal17, Fol12a]. Work [Shp10, Avi14, CR15, Don11, GK11, HWWA11, LSW19, Riv16, dMPQZ15]. works [SRM^{+}17]. World [Gre12b, KM10, EMS11, Bar17, Dal14a, Eis10, Far12a, Ger16]. would [Dal12a, Gra16]. writer [Jac17]. Writing [Jab19, Fol11]. Writings [Cur12, Neu11]. written [BBE^{+}13]. wrong [FGLG13, Blu16]. Wu [Wu17]. WW2 [Don14]. www.topsecretrosies.com [Gre12b]. WYSIWYG [Grä11].

X [Ada11]. Xeitgeist [And16].

Yale [Dal12a]. Yang [Mar12]. Yau [Tos18]. year [Car13, Hou19a, May19b]. years [Ano17a, Gra16, Mil11, Ric14]. Yogiisms [Tre16]. Yor [LP14]. York [Han12, Ste14, Str15, Ald11, Bau15, Bla11a, Bla14b, Bre16, Col16, Daw18, DeT17, Far12a, Fel14, Ger16, Law19, Man15, Mil19, Sch13b, Spe10a, Ste11, Van15]. young [Sau16]. Yuri [Cor13]. Yves [DS18, CMJ^{+}18].

Zeilberger [Tef10]. zero [Fol12a]. zeta [KvP16]. Zhang [Cor16].

References

Agarwal:2019:MCW

[AABT^{+}19] Nikita Agarwal, Carolina Araujo, Petra Bonfert-Taylor, Mojgan Mahmoudi, Marie Françoise Ouedraogo, Olga Paris-Romaskevich, Marie-Françoise Roy, Elisabetta Strickland, and Andrea Vera Gajardo. May 12: celebrating women in mathematics. From one idea to one hundred events. Notices of the American Mathematical Society, 66(11):1879– 1886, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Abbes:2019:MGM

[Abb19] Ahmed Abbes. Memories of a grandmaster: Michel Raynaud and the flattening by blowing up. Notices of the American Mathematical Society, 66(1):42– 44, January 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Adler:2018:PJS

[ABDS18] Jeffrey Adler, John Boller, Stephen DeBacker, and Loren Spice. Paul J. Sally Jr. (1933– 2013). Notices of the American Mathematical Society, 65 (5):555–565, May 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With remembrances.

Aboufadel:2019:MJS

[Abo19] Edward F. Aboufadel. In memoriam: Jane Smiley Cronin Scanlon. Notices of the American Mathematical Society, 66(9): 1448–1452, October 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Abrams:2016:WLL

[Abr16] Gene Abrams. What is … a Leavitt path algebra? Notices of the American Mathematical Society, 63(8):910–911, September 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Amendola:2018:TAS

[ACP18] Carlos Améndola, Marta Casanellas, and Luis David García Puente. Tapas of algebraic statistics. Notices of the American Mathematical Society, 65(8): 936–938, September 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Adam:2011:PXB

[Ada11] John Adam. Putting the X in biology: a review of The mathematics of life [MR2759991]. Notices of the American Mathematical Society, 58(11):1572–1578, December 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Adams:2018:WLH

[Ada18] Colin Adams. What is … a hyperbolic 3-manifold? Notices of the American Mathematical Society, 65(5):544–546, May 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anastasio:2015:MJS

[ADF^{+}15] Sal Anastasio, Ronald G.

Douglas, Ciprian Foias, Wai-Mee Ching, Martin Davis, Micha Sharir, Michael Wigler, Joseph A. Fisher, Peter D. Lax, Louis Nirenberg, Paul M. Willig, and Horacio Porta. In memory of Jacob Schwartz. Notices of the American Mathematical Society, 62(5):473–490, May 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Sal Anastasio, coordinating editor.

Adlaj:2012:EFP

[Adl12] Semjon Adlaj. An eloquent formula for the perimeter of an ellipse. Notices of the American Mathematical Society, 59(8):1094–1099, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

                                        www.ams.org/notices/201208/
                                        rtx120801094p.pdf.

Araiza:2019:WLM

[AdS19] Roy Araiza and Rolando de Santiago. R we living in the matrix? Notices of the American Mathematical Society, 66 (8):1216–1224, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arnold:2011:NN

[AF11] Douglas N. Arnold and Kristine K. Fowler. Nefarious numbers. Notices of the American Mathematical Society, 58 (3):434–437, March 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arroyo:2013:MPF

[AGLP^{+}13] María José Arroyo, Rogelio Fernández-Alonso González, Sergio R. López-Permouth, José Ríos Montes, and Carlos Signoret. In memory of Professor Francisco Federico Raggi Cárdenas. Notices of the American Mathematical Society, 60 (10):1342–1343, November 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arnoux:2014:WLR

[AH14] Pierre Arnoux and Edmund Harriss. What is … a Rauzy fractal? Notices of the American Mathematical Society, 61(7):768–770, August 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Adiprasito:2017:HTM

[AHK17] Karim Adiprasito, June Huh, and Eric Katz. Hodge theory of matroids. Notices of the American Mathematical Society, 64 (1):26–30, January 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Allen:2019:PMC

[AJ19] Jonathan E. Allen and Crystal Jaing. Profiling microbial communities with shotgun metagenomics. Notices of the American Mathematical Society, 66(9): 1498–1499, October 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Abe:2012:MGB

[AK12] Mari Abe and Mei Kobayashi.

Math girls [book review, Bento Books, Austin, TX, 2011]. Notices of the American Mathematical Society, 59(7):956–959, August 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aldous:2011:BSI

[Ald11] David Aldous. The Black Swan: the impact of the highly improbable [book review; Random House, New York, 2007]. Notices of the American Mathematical Society, 58(3):427–431, March 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Alexanderson:2011:RCR

[Ale11] Gerald L. Alexanderson. Remembering Constance Reid (1918–2010). Notices of the American Mathematical Society, 58(10):1458–1459, November 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Andrews:2016:MWK

[And16] George E. Andrews. The man who knew infinity: a report on the movie [film review, Animus Films, Edward R. Pressman Film, Exit Strategy Productions, Firecracker Entertainment, in association with Xeitgeist Entertainment Group, [2015?]]. Notices of the American Mathematical Society, 63(2):178–190, February 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:FP

[Ano10a] Anonymous. 2009 Fulkerson Prizes. Notices of the American Mathematical Society, 57 (11):1475–1476, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:CMA

[Ano10b] Anonymous. 2010 Chern Medal awarded. Notices of the American Mathematical Society, 57 (11):1472–1474, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:GPA

[Ano10c] Anonymous. 2010 Gauss Prize awarded. Notices of the American Mathematical Society, 57 (11):1466–1469, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:NPA

[Ano10d] Anonymous. 2010 Nevanlinna Prize awarded. Notices of the American Mathematical Society, 57(11):1469–1471, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:BRS

[Ano10e] Anonymous. Bourgain receives 2010 Shaw Prize. Notices of the American Mathematical Society, 57(8):987, September 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:FMA

[Ano10f]  Anonymous.  Fields Medals

awarded. Notices of the American Mathematical Society, 57 (11):1459–1465, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2010:IPS

[Ano10g] Anonymous. Ilya Piatetski-Shapiro, in memoriam. Notices of the American Mathematical Society, 57(10):1260–1275, November 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Jim Cogdell, Steve Gelbart, and Peter Sarnak, coordinating editors.

Anonymous:2010:LRK

[Ano10h] Anonymous. Lovász receives Kyoto Prize. Notices of the American Mathematical Society, 57 (9):1136, October 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:CP

[Ano11a] Anonymous. 2001 Conant Prize. Notices of the American Mathematical Society, 58(4):597–598, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:BP

[Ano11b] Anonymous. 2011 Bôcher Prize. Notices of the American Mathematical Society, 58(4):603–605, April 2011. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:CPN

[Ano11c] Anonymous. 2011 Cole Prize in number theory. Notices of the American Mathematical Society, 58(4):610–611, April 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:DP

[Ano11d] Anonymous. 2011 Doob Prize.

Notices of the American Mathematical Society, 58(4):606–607, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:EP

[Ano11e] Anonymous. 2011 Eisenbud Prize. Notices of the American Mathematical Society, 58 (4):608–609, April 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:MP

[Ano11f] Anonymous. 2011 Morgan Prize.

Notices of the American Mathematical Society, 58(4):599–600, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:SPb

[Ano11g] Anonymous. 2011 Satter Prize.

Notices of the American Mathematical Society, 58(4):601–602, April 2011. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:SPa

[Ano11h] Anonymous. 2011 Steele Prizes. Notices of the American Mathematical Society, 58(4):593–596, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:SPH

[Ano11i]  Anonymous.  2011 Steele

Prizes [Henryk Iwaniec, Ingrid Daubechies, John W. Milnor]. Notices of the American Mathematical Society, 58(4):593–596, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201104/
rtx110400593p.pdf.

Anonymous:2011:CHA

[Ano11j] Anonymous. Christodoulou and Hamilton awarded Shaw Prize. Notices of the American Mathematical Society, 58(9):1293, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:GH

[Ano11k] Anonymous. Gerhard Hochschild (1915–2010). Notices of the American Mathematical Society, 58(8):1078–1099, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Walter Ferrer Santos and Martin Moskowitz, coordinating editors.

Anonymous:2011:MMG

[Ano11l] Anonymous. Memories of Martin Gardner. Notices of the American Mathematical Society, 58 (3):418–422, March 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Steven G. Krantz, coordinating editor.

Anonymous:2011:MRA

[Ano11m] Anonymous. Milnor receives 2011 Abel Prize. Notices of the American Mathematical Society, 58(6):831–832, June/July 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:PAB

[Ano11n] Anonymous. Palis awarded Balzan Prize. Notices of the American Mathematical Society, 58(1):66–67, January 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:PHC

[Ano11o] Anonymous. Peter Hilton:

codebreaker and mathematician (1923–2010). Notices of the American Mathematical Society, 58(11):1538–1551, December 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Jean Pedersen, coordinating editor.

Anonymous:2011:PVI

[Ano11p] Anonymous. Presidential views:

interview with Eric Friedlander.

Notices of the American Mathematical Society, 58(3):473–475, March 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2011:RJJ

[Ano11q] Anonymous. Remembering Johannes J. Duistermaat (1942– 2010). Notices of the American Mathematical Society, 58(6): 794–802, June/July 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Victor Guillemin, Álvaro Pelayo, San Vũ Ngo.c and Alan Weinstein, coordinating editors.

Anonymous:2011:RLE

[Ano11r] Anonymous. Remembering Leon Ehrenpreis (1930–2010). Notices of the American Mathematical Society, 58(5):674–681, May 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With obituaries by Daniele C. Struppa, Hershel Farkas, Takahiro Kawai, Peter Kuchment, Eric Todd Quinto, Shlomo Sternberg and Alan Taylor.

Anonymous:2011:SSC

[Ano11s] Anonymous. Shiing-Shen Chern (1911–2004). Notices of the American Mathematical Society, 58(9):1226–1249, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Shing-Tung Yau, coordinating editor.

Anonymous:2011:TDB

[Ano11t] Anonymous. A tribute to David Blackwell. Notices of the American Mathematical Society, 58 (7):912–928, August 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). George G. Roussas, coordinating editor.

Anonymous:2011:THH

[Ano11u] Anonymous. A tribute to Henry Helson. Notices of the American Mathematical Society, 58(2): 274–288, February 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Donald Sarason, coordinating editor.

Anonymous:2011:VRT

[Ano11v] Anonymous. Valiant receives 2010 Turing award. Notices of the American Mathematical Society, 58(6):833–834, June/ July 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:DQ

[Ano12a] Anonymous. Daniel Quillen. Notices of the American Mathematical Society, 59(10):1392–1406, November 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Eric Friedlander and Daniel Grayson, coordinating editors.

Anonymous:2012:DAN

[Ano12b] Anonymous. Daubechies awarded Nemmers Prize. Notices of the American Mathematical Society, 59(7):962, August 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:GBB

[Ano12c] Anonymous. Glimpses of Benoît B. Mandelbrot (1924–2010). Notices of the American Mathematical Society, 59(8):1056–1063, September 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:IBB

[Ano12d] Anonymous. The influence of Benoît B. Mandelbrot on mathematics. Notices of the American Mathematical Society, 59(9): 1208–1221, October 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:KAS

[Ano12e] Anonymous. Kontsevich awarded Shaw Prize. Notices of the American Mathematical Society, 59 (9):1261, October 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:MVA

[Ano12f]  Anonymous.  Memories of

Vladimir Arnold. Notices of the American Mathematical Society, 59(4):482–502, April 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Boris Khesin and Serge Tabachnikov, coordinating editors.

Anonymous:2012:PRT

[Ano12g] Anonymous. Pearl receives 2011 Turing Award. Notices of the American Mathematical Society, 59(6):837–838, June/July 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:RHF

[Ano12h] Anonymous. Remembering Herbert Federer (1920–2010). Notices of the American Mathematical Society, 59(5):622–631, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Harold Parks, coordinating editor.

Anonymous:2012:RJM

[Ano12i] Anonymous. Remembering Jerry Marsden (1942–2010). Notices of the American Mathematical Society, 59(6):758–775, June/ July 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Tudor Ratiu and Alan Weinstein, coordinating editors.

Anonymous:2012:SRG

[Ano12j] Anonymous. Srinivasa Ramanujan: going strong at 125, Part I. Notices of the American Mathematical Society, 59(11):1522– 1537, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Krishnaswami Alladi, editor.

Anonymous:2012:SRA

[Ano12k] Anonymous. Szemerédi receives 2012 Abel Prize. Notices of the American Mathematical Society, 59(6):836–837, June/July 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2012:TNJ

[Ano12l] Anonymous. A tribute to Nigel J. Kalton (1946–2010). Notices of the American Mathematical Society, 59(7):942–951, August 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Peter G. Casazza, coordinating editor.

Anonymous:2012:TVA

[Ano12m] Anonymous. Tribute to Vladimir Arnold. Notices of the American Mathematical Society, 59 (3):378–399, March 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Boris Khesin and Serge Tabachnikov, coordinating editors.

Anonymous:2013:AFH

[Ano13a] Anonymous. Andrew Fielding Huxley (1917–2012). Notices of the American Mathematical Society, 60(5):576–584, May 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Michael G. Mackey and Moisés Santillán, coordinating editors.

Anonymous:2013:IMGa

[Ano13b] Anonymous. Israel Moiseevich Gelfand. Notices of the American Mathematical Society, 60 (1):24–49, January 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Vladimir Retakh, coordinating editor.

Anonymous:2013:IMGb

[Ano13c] Anonymous. Israel Moiseevich Gelfand, Part II. Notices of the American Mathematical Society, 60(2):162–171, February 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Vladimir Retakh, coordinating editor.

Anonymous:2013:RJS

[Ano13d] Anonymous. Recalling James Serrin. Notices of the American Mathematical Society, 60(6): 700–712, June/July 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Howard Levine and Hans Weinberger, coordinating editors.

Anonymous:2013:RBG

[Ano13e] Anonymous. Remembering Basil Gordon, 1931–2012. Notices of the American Mathematical Society, 60(7):856–865, August 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Krishnaswami Alladi, coordinating editor.

Anonymous:2013:RRB

[Ano13f]  Anonymous.  Remembering

Raoul Bott (1923–2005). Notices of the American Mathematical Society, 60(4):398–416, April 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Loring W. Tu, coordinating editor.

Anonymous:2013:RSR

[Ano13g] Anonymous. Remembering Steve Rallis. Notices of the American Mathematical Society, 60(4):466–469, April 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Jim Cogdell and Dihua Jiang, coordinating editors.

Anonymous:2013:RWR

[Ano13h] Anonymous. Remembering Walter Rudin (1921–2010). Notices of the American Mathematical Society, 60(3):295–301, March 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Alexander Nagel and Edgar Lee Stout, coordinating editors.

Anonymous:2013:SRG

[Ano13i] Anonymous. Srinivasa Ramanujan: going strong at 125, Part II. Notices of the American Mathematical Society, 60(1):10– 23, January 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Krishnaswami Alladi, Editor.

Anonymous:2014:DAS

[Ano14a] Anonymous. Donoho awarded 2013 Shaw Prize. Notices of the American Mathematical Society, 61(3):284–285, March 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2014:FH

[Ano14b] Anonymous. Friedrich Hirzebruch (1927–2012). Notices of the American Mathematical Society, 61(7):706–727, August 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Michael Atiyah and Don Zagier, coordinating editors.

Anonymous:2014:LAS

[Ano14c] Anonymous. Lusztig awarded Shaw Prize. Notices of the American Mathematical Society, 61 (7):774–775, August 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2014:RSK

[Ano14d] Anonymous. Remembering Shoshichi Kobayashi. Notices of the American Mathematical Society, 61(11):1322–1332, December 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Gary R. Jensen, coordinating editor.

Anonymous:2014:SAJ

[Ano14e] Anonymous. Shreeram Abhyankar (July 22, 1930–November 2, 2012). Notices of the American Mathematical Society, 61 (10):1196–1216, November 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Shashikant Mulay and Avinash Sathaye, coordinating editors.

Anonymous:2014:THG

[Ano14f] Anonymous. A tribute to Hans Grauert. Notices of the American Mathematical Society, 61 (5):472–483, May 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Alan Huckleberry and Thomas Peternell, coordinating editors.

Anonymous:2015:FIA

[Ano15a] Anonymous. Faltings and Iwaniec awarded Shaw Prize. Notices of the American Mathematical Society, 62(9):1074– 1075, October 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2015:HW

[Ano15b] Anonymous. Herbert S. Wilf (1931–2012). Notices of the American Mathematical Society, 62(4):346–360, April 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Fan Chung, Curtis Greene and Joan Hutchinson, coordinating editors.

Anonymous:2015:MAS

[Ano15c] Anonymous. In memory of Arthur Strong Wightman. Notices of the American Mathematical Society, 62(3):249–257, March 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Barry Simon, coordinating editor.

Anonymous:2015:JLM

[Ano15d] Anonymous. Joram Lindenstrauss, in memoriam. Notices of the American Mathematical Society, 62(1):26–37, January 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Complied by William B. Johnson and Gideon Schechtman.

Anonymous:2015:MME

[Ano15e] Anonymous. Memories of Mary Ellen Rudin. Notices of the American Mathematical Society, 62(6):617–629, June/July 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2015:RPEa

[Ano15f]  Anonymous.  Reflections on

Paul Erdős on his birth centenary. Notices of the American Mathematical Society, 62(2): 121–143, February 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Krishnaswami Alladi and Steven Krantz, coordinating editors.

Anonymous:2015:RPEb

[Ano15g] Anonymous. Reflections on Paul Erdős on his birth centenary, Part II. Notices of the American Mathematical Society, 62 (3):226–247, March 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Krishnaswami Alladi and Steven G. Krantz, coordinating editors.

Anonymous:2015:MLH

[Ano15h] Anonymous. To the memory of Lars Hörmander (1931–2012). Notices of the American Mathematical Society, 62(8):890–907, September 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Jan Boman and Ragnar Sigurdsson, coordinating editors.

Anonymous:2015:WBA

[Ano15i] Anonymous. William B. Arveson: a tribute. Notices of the American Mathematical Society, 62(7):761–769, August 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Daniel Markiewicz and Palle E. T. Jorgensen, coordinating editors.

Anonymous:2015:WPT

[Ano15j]  Anonymous.  William P.

Thurston, 1946–2012. Notices of the American Mathematical Society, 62(11):1318–1332, December 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). David Gabai and Steve Kerchoff, coordinating editors.

Anonymous:2016:AGPa

[Ano16a] Anonymous. Alexandre Grothendieck 1928–2014, Part 1. Notices of the American Mathematical Society, 63(3):242–255, March 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Michael Artin, Allyn Jackson, David Mumford and John Tate, coordinating editors.

Anonymous:2016:AGPb

[Ano16b] Anonymous. Alexandre Grothendieck 1928–2014, Part 2. Notices of the American Mathematical Society, 63(4):401–413, April 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2016:MPAa

[Ano16c] Anonymous. From mathematical physics to analysis: a walk in Barry Simon’s mathematical garden. Notices of the American Mathematical Society, 63 (7):740–752, August 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Fritz Gesztesy, coordinating editor.

Anonymous:2016:MPAb

[Ano16d] Anonymous. From mathematical physics to analysis: a walk in Barry Simon’s mathematical garden, II. Notices of the American Mathematical Society, 63(8): 878–889, September 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Fritz Gesztesy, coordinating editor.

Anonymous:2016:HGS

[Ano16e] Anonymous. How Grothendieck simplified algebraic geometry. Notices of the American Mathematical Society, 63(3):256–265, March 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2016:JFN

[Ano16f] Anonymous. John Forbes Nash Jr. (1928–2015). Notices of the American Mathematical Society, 63(5):492–504, May 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Camillo De Lellis, coordinating editor.

Anonymous:2016:RAN

[Ano16g] Anonymous. Recent applications of Nirenberg’s classical ideas. Notices of the American Mathematical Society, 63(2):126–134, February 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Communicated by Christina Sormani.

Anonymous:2016:WPT

[Ano16h] Anonymous. William P. Thurston, 1946–2012. Notices of the American Mathematical Society, 63(1):31–41, January 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). David Gabai and Steve Kerckhoff, coordinating editors.

Anonymous:2017:YAJ

[Ano17a] Anonymous. 100 years ago: Joan Clarke. Notices of the American Mathematical Society, 64 (3):252–255, March 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2017:AOP

[Ano17b] Anonymous. Aleksander (Olek)

Pelczynski 1932–2012. Notices of the American Mathematical Society, 64(1):54–58, January 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions.

Anonymous:2017:JBK

[Ano17c] Anonymous. Joseph B. Keller (1923–2016). Notices of the American Mathematical Society, 64(6):606–615, June/July 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2017:WLP

[Ano17d] Anonymous. What is … perimeter? Notices of the American Mathematical Society, 64 (9):1009, October 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:ALW

[Ano18a] Anonymous. 2018 Albert Leon Whiteman Prize. Notices of the American Mathematical Society, 65(4):472–474, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:ADP

[Ano18b] Anonymous. 2018 Award for distinguished public service. Notices of the American Mathematical Society, 65(4):475–476, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:BRP

[Ano18c] Anonymous. 2018 Bertrand Russell Prize. Notices of the American Mathematical Society, 65 (4):470–471, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:CPL

[Ano18d] Anonymous. 2018 Chevalley Prize in Lie Theory. Notices of the American Mathematical Society, 65(4):459–460, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:FBM

[Ano18e] Anonymous. 2018 Frank and Brennie Morgan Prize for outstanding research in mathematics by an undergraduate student. Notices of the American Mathematical Society, 65(4):465–467, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:FNC

[Ano18f] Anonymous. 2018 Frank Nelson Cole Prize in algebra. Notices of the American Mathematical Society, 65(4):461–462, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:GDB

[Ano18g] Anonymous. 2018 George David Birkhoff Prize in applied mathematics. Notices of the American Mathematical Society, 65 (4):477–478, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:LPS

[Ano18h] Anonymous. 2018 Leroy P.

Steele Prizes. Notices of the American Mathematical Society, 65(4):455–458, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:LCP

[Ano18i] Anonymous. 2018 Levi L. Conant Prize. Notices of the American Mathematical Society, 65 (4):463–464, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:UGP

[Ano18j] Anonymous. 2018 Ulf Grenander Prize in stochastic theory and modeling. Notices of the American Mathematical Society, 65(4):468–469, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:AC

[Ano18k] Anonymous. Alice Chang. Notices of the American Mathematical Society, 65(3):254–255, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:AW

[Ano18l] Anonymous. Amie Wilkinson. Notices of the American Mathematical Society, 65(3):298–299, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:AN

[Ano18m] Anonymous. Andrea Nahmod. Notices of the American Mathematical Society, 65(3):282–283, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:CC

[Ano18n] Anonymous. Carina Curto. Notices of the American Mathematical Society, 65(3):258–259, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:CG

[Ano18o] Anonymous. Carolyn Gordon. Notices of the American Mathematical Society, 65(3):264–265, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:CW

[Ano18p] Anonymous. Chelsea Walton.

Notices of the American Mathematical Society, 65(3):296–297, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:DM

[Ano18q] Anonymous. Dusa McDuff. Notices of the American Mathematical Society, 65(3):278–279, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:ER

[Ano18r] Anonymous. Emily Riehl. Notices of the American Mathematical Society, 65(3):288–289, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:ET

[Ano18s] Anonymous. Éva Tardos. Notices of the American Mathematical Society, 65(3):294–295, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:FCG

[Ano18t] Anonymous. Fan Chung Graham. Notices of the American Mathematical Society, 65 (3):266–267, March 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:FB

[Ano18u] Anonymous. Felix Browder (1927–2016). Notices of the American Mathematical Society, 65(11):1398–1411, December 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions.

Anonymous:2018:FH

[Ano18v] Anonymous. Fern Y. Hunt. Notices of the American Mathematical Society, 65(3):274–275, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:GS

[Ano18w] Anonymous. Gigliola Staffilani. Notices of the American Mathematical Society, 65(3):292–293, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:HFW

[Ano18x] Anonymous. Hans F. Weinberger (1928–2017). Notices of the American Mathematical Society, 65(7):850–855, August 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Don Aronson, Peter Olver and Fadil Santosa, coordinating editors, With remembrances.

Anonymous:2018:ID

[Ano18y] Anonymous. Ingrid Daubechies. Notices of the American Mathematical Society, 65(3):260– 261, March 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic).  URL http://www.
ams.org/journals/notices/
201803/201803FullIssue.pdf.

Special issue in honor of Women’s History Month, with profiles of 27 women mathematicians.

Anonymous:2018:IF

[Ano18z] Anonymous. Irene Fonseca. Notices of the American Mathematical Society, 65(3):262–263, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:JGA

[Ano18-27] Anonymous. James G. Arthur:

AMS 2017 Leroy P. Steele Prize for Lifetime Achievement. Notices of the American Mathematical Society, 65(6):637–645, June/July 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). James Cogdell and Freydoon Shahidi, Guest Editors.

Anonymous:2018:JCP

[Ano18-28] Anonymous. Jill C. Pipher. Notices of the American Mathematical Society, 65(3):286–287, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:JB

[Ano18-29] Anonymous. Joan Birman. Notices of the American Mathematical Society, 65(3):250–251, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:KS

[Ano18-30] Anonymous. Karen E. Smith. Notices of the American Mathematical Society, 65(3):290–291, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:LW

[Ano18-31] Anonymous. Lauren Williams. Notices of the American Mathematical Society, 65(3):300–301, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:LP

[Ano18-32] Anonymous. Lillian Pierce. Notices of the American Mathematical Society, 65(3):284–285, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:MCb

[Ano18-33] Anonymous. Maria Chudnovsky. Notices of the American Mathematical Society, 65(3):256–257, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:MM

[Ano18-34] Anonymous. Maryam Mirzakhani: 1977–2017. Notices of the American Mathematical Society, 65(10):1221–1247, November 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Hélène Barcelo and Stephen Kennedy, coordinating editors, With remembrances.

Anonymous:2018:MCS

[Ano18-35] Anonymous. The mathematics of Cathleen Synge Morawetz. Notices of the American Mathematical Society, 65(7):764–778, August 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With remembrances.

Anonymous:2018:MMW

[Ano18-36] Anonymous. Melanie Matchett Wood. Notices of the American Mathematical Society, 65 (3):302–303, March 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:MCa

[Ano18-37] Anonymous. Melody Chan. Notices of the American Mathematical Society, 65(3):252–253, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:OH

[Ano18-38] Anonymous. Olga Holtz. Notices of the American Mathematical Society, 65(3):272–273, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:RLA

[Ano18-39] Anonymous. Robert Langlands awarded Abel Prize. Notices of the American Mathematical Society, 65(6):670–672, June/ July 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With a biography written by Alexander Bellos.

Anonymous:2018:RG

[Ano18-40] Anonymous. Rosemary Guzman. Notices of the American Mathematical Society, 65(3):268–269, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:SM

[Ano18-41] Anonymous. Sophie Morel. Notices of the American Mathematical Society, 65(3):280–281, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:TH

[Ano18-42] Anonymous. Tara S. Holm. Notices of the American Mathematical Society, 65(3):270–271, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:TJ

[Ano18-43] Anonymous. Trachette L. Jackson. Notices of the American Mathematical Society, 65(3): 276–277, March 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:TJD

[Ano18-44] Anonymous. Tribute to Joe Diestel (1943–2017). Notices of the American Mathematical Society, 65(7):845–847, August 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Richard Aron, coordinating editor, With remembrances.

Anonymous:2018:WIR

[Ano18-45] Anonymous. What is … a rectifiable set? Notices of the American Mathematical Society, 65(9): 1086–1087, October 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2018:WLH

[Ano18-46] Anonymous. What is … hypoellipticity? Notices of the American Mathematical Society, 65(4):418–419, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2019:KUA

[Ano19a] Anonymous. Karen Uhlenbeck awarded Abel Prize. Notices of the American Mathematical Society, 66(6):939–941, June/ July 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2019:MSM

[Ano19b] Anonymous. Memories of Sir Michael Atiyah. Notices of the American Mathematical Society, 66(11):1834–1852, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With remembrances.

Awodey:2013:VUA

[APW13] Steve Awodey, Álvaro Pelayo, and Michael A. Warren. Voevodsky’s univalence axiom in homotopy type theory. Notices of the American Mathematical Society, 60(9):1164–1167, October 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Archibald:2013:SDM

[Arc13] Tom Archibald. Sources in the development of mathematics [book review of MR2807493]. Notices of the American Mathematical Society, 60(10):1331– 1333, November 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ardila:2018:GM

[Ard18] Federico Ardila. The geometry of matroids. Notices of the American Mathematical Society, 65(8): 902–908, September 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arianrhod:2013:EIB

[Ari13] Robyn Arianrhod. The evolution of an idea [book review of MR2994999]. Notices of the American Mathematical Society, 60(7):916–919, August 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arianrhod:2015:ESS

[Ari15] Robyn Arianrhod. Enlightening symbols: a short history of mathematical notation and its hidden powers [book review of MR3183982]. Notices of the American Mathematical Society, 62(2):148–151, February 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Atiyah:2010:AM

[Ati10] Michael Atiyah. The art of mathematics. Notices of the American Mathematical Society, 57(1): 8, January 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Audin:2010:MFN

[Aud10] Michèle Audin. Mathematicians fleeing from Nazi Germany [book review of MR2522825]. Notices of the American Mathematical Society, 57(10):1300–1302, November 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Audin:2011:HAS

[Aud11] Michèle Audin. At home with André and Simone Weil [book review, Northwestern Univ. Press, Evanston, IL, 2010]. Notices of the American Mathematical Society, 58(5):697–698, May 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Auel:2019:MGM

[Aue19] Asher Auel. The mathematics of Grace Murray Hopper. Notices of the American Mathematical Society, 66(3):330–340, March 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Auslander:2018:HMP

[Aus18] Joseph Auslander. Humanizing mathematics and its philosophy [book review of MR3729276]. Notices of the American Mathematical Society, 65(11):1424– 1429, December 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Avigad:2011:POD

[Avi11]  Jeremy Avigad.  Proof and

other dilemmas: mathematics and philosophy [book review of MR2435289]. Notices of the American Mathematical Society, 58(11):1580–1584, December 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Avigad:2014:ATH

[Avi14] Jeremy Avigad. Alan Turing: his work and impact [book review, Elsevier Sci., 2013]. Notices of the American Mathematical Society, 61(8):886–890, September 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201408/
rnoti-p886.pdf.

Avigad:2016:LLG

[Avi16] Jeremy Avigad. Logic’s lost genius and Gentzen’s centenary [book reviews of MR2363580 and MR3408723]. Notices of the American Mathematical Society, 63(11):1288–1292, December 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Avigad:2018:MM

[Avi18] Jeremy Avigad. The mechanization of mathematics. Notices of the American Mathematical Society, 65(6):681–690, June/July 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Axler:2017:DAB

[Axl17] Sheldon Axler. A doubter’s almanac [book review of MR3674587]. Notices of the American Mathematical Society, 64(3):265–267, March 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beauville:2018:AHC

[BAD^{+}18] Arnaud Beauville, Ekaterina Amerik, Olivier Debarre, Phillip Griffiths, Daniel Huybrechts, and François Charles. Ad honorem Claire Voisin. Notices of the American Mathematical Society, 65(4):390–399, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2019:FMP

[Bae19] John C. Baez. Foundations of mathematics and physics one century after Hilbert: new perspectives [book review of MR3793646]. Notices of the American Mathematical Society, 66(10):1690–1692, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Balakrishnan:2019:WLC

[Bal19] Jennifer S. Balakrishnan. What is … a Coleman integral? Notices of the American Mathematical Society, 66(5):738–740, May 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bandle:2017:DPI

[Ban17] Catherine Bandle. Dido’s problem and its impact on modern mathematics. Notices of the American Mathematical Society, 64(9):980–984, October 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barnsley:2010:LSM

[Bar10] Michael F. Barnsley. The life and survival of mathematical ideas. Notices of the American Mathematical Society, 57(1):10– 22, January 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barany:2013:SMC

[Bar13] Michael J. Barany. Stuck in the middle: Cauchy’s intermediate value theorem and the history of analytic rigor. Notices of the American Mathematical Society, 60(10):1334–1338, November 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barany:2015:MM

[Bar15] Michael J. Barany. The myth and the medal. Notices of the American Mathematical Society, 62(1):15–20, January 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barany:2017:WWI

[Bar17] Michael J. Barany. The World War II origins of mathematics awareness. Notices of the American Mathematical Society, 64 (4):363–367, April 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Batterson:2011:CJP

[Bat11] Steve Batterson. The contribution of John Parker Jr. to American mathematics. Notices of the American Mathematical Society, 58(2):262–273, February 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Batterson:2014:HSB

[Bat14]  Steve Batterson.  A his-

tory in sum [book review of MR3100544]. Notices of the American Mathematical Society, 61(6):603–605, June/July 2014.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bauer:2015:MTB

[Bau15] Andrej Bauer. Mathematics of the transcendental [book review, Bloomsbury, New York, 2014]. Notices of the American Mathematical Society, 62(9): 1070–1071, October 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bays:2012:PMP

[Bay12] Timothy Bays. The philosophy of mathematical practice [book review of MR2590934]. Notices of the American Mathematical Society, 59(3):424–428, March 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bailey:2011:EEC

[BB11] David H. Bailey and Jonathan M. Borwein. Exploratory experimentation and computation. Notices of the American Mathematical Society, 58 (10):1410–1419, November 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.
edu.au/1396/;   https://www.
ams.org/notices/201110/rtx111001410p.

pdf. Chinese translation in Mathematical Advances in Translation (Chinese Academy of Science), June 2012.

Bailey:2014:RTP

[BB14] D. H. Bailey and J. M. Borwein.

The recent trademarking of pi: a troubling precedent. Notices of the American Mathematical Society, 61(10):121–124, November 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http:/

/docserver.carma.newcastle.
edu.au/id/eprint/1501;   http:
//www.ams.org/notices/201410/
rnoti-p1224.pdf.

Borwein:2015:LKM

[BBC15] Jonathan M. Borwein, Yann Bugeaud, and Michael Coons. The legacy of Kurt Mahler. Notices of the American Mathematical Society, 62(5):526–531, May 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201505/
rnoti-p526.pdf.

Bailey:2014:PMF

[BBdPZ14] David H. Bailey, Jonathan M.

Borwein, Marcos López de Prado, and Qiji Jim Zhu. Pseudomathematics and financial charlatanism: the effects of backtest overfitting on out-of-sample performance. Notices of the American Mathematical Society, 61 (5):458–471, May 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bair:2013:MHW

[BBE^{+}13] Jacques Bair, Piotr Błaszczyk, Robert Ely, Valérie Henry, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Semen S. Kutateladze, Thomas McGaffey, David M. Schaps, David Sherry, and Steven Shnider. Is mathematical history written by the victors? Notices of the American Mathematical Society, 60 (7):886–904, August 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brubaker:2011:ESC

[BBF11] Benjamin Brubaker, Daniel Bump, and Solomon Friedberg. Eisenstein series, crystals, and ice. Notices of the American Mathematical Society, 58(11): 1563–1571, December 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bascelli:2014:FLE

[BBH^{+}14] Tiziana Bascelli, Emanuele Bottazzi, Frederik Herzberg, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Tahl Nowik, David Sherry, and Steven Shnider. Fermat, Leibniz, Euler, and the gang: the true history of the concepts of limit and shadow. Notices of the American Mathematical Society, 61(8):848–864, September 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bryant:2019:OSO

[BBKL19] Robert Bryant, Ron Buckmire, Lily Khadjavi, and Douglas Lind. The origins of spectra, an organization for LGBT mathematicians. Notices of the American Mathematical Society, 66(6): 875–882, June/July 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bailey:2014:PMFb

[BBLZ14] David H. Bailey, Jonathan M.

Borwein, Marcos López de Prado, and Qiji Jim Zhu. Pseudo-mathematics and financial charlatanism: the effects of backtest overfitting on outof-sample performance. Notices of the American Mathematical Society, 61(5):458–471, May 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http:/

/docserver.carma.newcastle.
edu.au/1506/;   http://www.
ams.org/notices/201405/rnoti-
p458.pdf.

Bailey:2013:CPI

[BBMW13] David H. Bailey, Jonathan M.

Borwein, Andrew Mattingly, and Glenn Wightwick. The computation of previously inaccessible digits of π^{2} and Catalan’s constant. Notices of the American Mathematical Society, 60(7):844–854, August 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.
edu.au/1436/;   http://www.
ams.org/notices/201307/rnoti-
p844.pdf.

Bucur:2017:TOP

[BBN17] Dorin Bucur, Giuseppe Buttazzo, and Carlo Nitsch. Two optimization problems in thermal insulation. Notices of the American Mathematical Society, 64(8):830–835, September 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bailey:2013:OSD

[BBS13] D. H. Bailey, J. M. Borwein, and Victoria Stodden. Opinion: Set the default to “open”. Notices of the American Mathematical Society, 60(6):679–680, June/ July 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

                                        www.ams.org/notices/201306/
                                        rnoti-p679.pdf.

Barany:2016:TTE

[BBZ16] Imre Bárány, Pavle V. M. Blagojević, and Günter M. Ziegler. Tverberg’s theorem at 50: extensions and counterexamples. Notices of the American Mathematical Society, 63(7):732–739, August 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borwein:2013:CFW

[BC13] Jonathan M. Borwein and Richard E. Crandall. Closed forms: What they are and why we care. Notices of the American Mathematical Society, 60(1):50–65, January 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.
edu.au/767/;   http://www.
ams.org/notices/201301/rnoti-
p50.pdf.

Birch:2019:PSD

[BCCTS19] Bryan Birch, John Coates, Jean-Louis Colliot-Thélène, and Alexei Skorobogatov. Peter Swinnerton-Dyer (1927–2018). Notices of the American Mathematical Society, 66(7):1058– 1067, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonatti:2016:WIB

[BCDW16] Ch. Bonatti, S. Crovisier, L. J.

Díaz, and A. Wilkinson. What is … a blender? Notices of the American Mathematical Society, 63(10):1175–1178, November 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Boisvert:2011:SFH

[BCLO11] Ronald F. Boisvert, Charles W.

Clark, Daniel W. Lozier, and Frank W. Olver. A special functions handbook for the digital age. Notices of the American Mathematical Society, 58(7): 905–911, August 1, 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.
ams.org/journals/notices/
201107/201107-full-issue.
pdf;   https://www.nist.gov/

                                        publications/special-functions-
                                        handbook-digital-age.

Bernstein:2017:FGC

[BD17] Mira Bernstein and Moon Duchin. A formula goes to court: partisan gerrymandering and the efficiency gap. Notices of the American Mathematical Society, 64(9):1020–1024, October 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baake:2016:WLA

[BDG16] Michael Baake, David Damanik, and Uwe Grimm. What is … aperiodic order? Notices of the American Mathematical Society, 63(6):647–650, June/July 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bardawil:2017:IFH

[BDJ^{+}17] Fadi Bardawil, Thomas Dodman, Ian Jauslin, Pascal Marichalar, Klaus Oschema, and Peter Redfield. Immigration, freedom, and the history of the Institute for Advanced Study. Notices of the American Mathematical Society, 64(10):1160–1168, November 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beals:2011:HMF

[Bea11] Richard Beals. Handbooks of mathematical functions, versions 1.0 and 2.0 [book review of MR2723248]. Notices of the American Mathematical Society, 58(8):1115–1118, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bedford:2012:WLB

[Bed12] Eric Bedford. What is … a biholomorphic mapping? Notices of the American Mathematical Society, 59(6):812–814, June/July 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beilinson:2016:MGH

[Bei16] A. Beilinson. I. M. Gelfand and his seminar — a presence. Notices of the American Mathematical Society, 63(3):295–298, March 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bellow:2016:FML

[Bel16] Alexandra Bellow. Flashbacks of a mathematical life. Notices of the American Mathematical Society, 63(8):931–934, September 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bennett:2010:TAM

[Ben10] Andrew G. Bennett. Tools of American mathematics teaching, 1800–2000 [book review of MR2426819]. Notices of the American Mathematical Society, 57(1):53–55, January 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bergner:2019:MHT

[Ber19a] Julia E. Bergner. Modeling homotopy theories. Notices of the American Mathematical Society, 66(9):1423–1432, October 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berwald:2019:MQE

[Ber19b] Jesse J. Berwald. The mathematics of quantum-enabled applications on the D-wave quantum computer. Notices of the American Mathematical Society, 66(6):832–841, June/July 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beukers:2014:HFH

[Beu14] Frits Beukers. Hypergeometric functions, how special are they? Notices of the American Mathematical Society, 61(1):48– 56, January 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barnsley:2012:GBM

[BF12a] Michael F. Barnsley and Michael Frame. Glimpses of Benoît Mandelbrot (1924–2010). Notices of the American Mathematical Society, 59(8):1056–1063, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

                                        www.ams.org/notices/201208/
                                        rtx120801056p.pdf.

Barnsley:2012:IBB

[BF12b] Michael F. Barnsley and Michael Frame. The influence of Benoît B. Mandelbrot on mathematics. Notices of the American Mathematical Society, 59 (9):1208–1221, October 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201209/
rtx120901208p.pdf. Other con-

tributors are Roger Howe, Ian Stewart, David Mumford, Hillel Furstenberg, Kenneth Falconer, Bruce J. West, Marc-Olivier Coppens, Nathan Cohen, Stéphane Jaffard, and Sir Michael Berry.

Boland:2010:EPU

[BFH10] John W. Boland, Jerzy A. Filar, and Phil G. Howlett. Environmental problems, uncertainty, and mathematical modeling. Notices of the American Mathematical Society, 57(10):1286–1294, November 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Babbitt:2011:FEQ

[BG11a] Donald Babbitt and Judith Goodstein. Federigo Enriques’s quest to prove the “completeness theorem”. Notices of the American Mathematical Society, 58(2): 240–249, February 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borthwick:2011:DDM

[BG11b] David Borthwick and Skip Garibaldi. Did a 1-dimensional magnet detect a 248-dimensional Lie algebra? Notices of the American Mathematical Society, 58(8):1055–1066, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Banyaga:2018:SMS

[BGL^{+}18] Augustin Banyaga, Wilfrid Gangbo, Florian Luca, Nancy Ann Neudauer, Carol Shubin, Cédric Villani, and Michel Waldschmidt. Snapshots of mathematics in Sub-Saharan Africa. Notices of the American Mathematical Society, 65(7):798–808, August 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bieri:2017:GWT

[BGY17] Lydia Bieri, David Garfinkle, and Nicolás Yunes. Gravitational waves and their mathematics. Notices of the American Mathematical Society, 64(7):693–707, August 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berger:2017:WLB

[BH17] Arno Berger and Theodore P. Hill. What is … Benford’s Law? Notices of the American Mathematical Society, 64(2): 132–134, February 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/publications/journals/
notices/201702/rnoti-p132.
pdf.

Bhatt:2014:WLP

[Bha14] Bhargav Bhatt. What is … a perfectoid space? Notices of the American Mathematical Society, 61(9):1082–1084, October 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barwick:2015:DMK

[BHMM15] Clark Barwick, Michael Hopkins, Haynes Miller, and Ieke Moerdijk. Daniel M. Kan (1927– 2013). Notices of the American Mathematical Society, 62(9): 1042–1045, October 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borwein:2012:OMT

[BJ12] Jonathan Borwein and Veselin Jungić. Organic mathematics: then and now. Notices of the American Mathematical Society, 59(3):416–419, March 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http:/

/docserver.carma.newcastle.
edu.au/1397/;   http://www.
ams.org/notices/201203/rtx120300416p.
pdf.

Beloshapka:2012:AA

[BKK^{+}12] Valery Beloshapka, Aleksey Kalinin, Michael Kuznetsov, Kirill Mayorov, Grigory Polotovskiy, Ivan Remizov, Yakov Sennikovskiy, George Shabat, Mikhail Tokman, Alexander Tumanov, and Grigoriy Zhislin. Alexander Abrosimov. Notices of the American Mathematical Society, 59(11):1569–1570, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brams:2014:TPF

[BKK14] Steven J. Brams, D. Marc Kilgour, and Christian Klamler. Two-person fair division of indivisible items: an efficient, envyfree algorithm. Notices of the American Mathematical Society, 61(2):130–141, February 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bauschke:2012:WLF

[BL12] Heinz H. Bauschke and Yves Lucet. What is … a Fenchel conjugate? Notices of the American Mathematical Society, 59(1):44– 46, January 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blank:2011:SJP

[Bla11a] Brian E. Blank. Symmetry: a journey into the patterns of nature [book review, Harper, New York, 2008]. Notices of the American Mathematical Society, 58(2):304–310, February 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blank:2011:SMH

[Bla11b] Brian E. Blank. The strangest man: the hidden life of Paul Dirac, mystic of the atom [book review of MR2807927]. Notices of the American Mathematical Society, 58(9):1278–1284, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blank:2014:BRT

[Bla14a] Brian E. Blank. Book review:

Turing’s Cathedral: The Origins of the Digital Universe. Notices of the American Mathematical Society, 61(7):759–767, August 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201407/
rnoti-p759.pdf.

Blank:2014:TCO

[Bla14b] Brian E. Blank. Turing’s cathedral: the origins of the digital universe [book review, Vintage, New York, 2012]. Notices of the American Mathematical Society, 61(7):759–767, August 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blumberg:2016:HWB

[Blu16] Andrew J. Blumberg. How not to be wrong [book review of MR3236985]. Notices of the American Mathematical Society, 63(3):301–303, March 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bray:2018:MRS

[BME^{+}18] Hubert L. Bray, William P. Minicozzi, II, Michael Eichmair, Lan-Hsuan Huang, Shing-Tung Yau, Karen Uhlenbeck, Rob Kusner, Fernando Codá Marques, Chikako Mese, and Ailana Fraser. The mathematics of Richard Schoen. Notices of the American Mathematical Society, 65(11):1349–1376, December 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Benyi:2010:WLP

[BMN10] Árpád Bényi, Diego Maldonado, and Virginia Naibo. What is … a paraproduct? Notices of the American Mathematical Society, 57(7):858–860, August 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bohun:2014:IMI

[Boh14] C. Sean Bohun. An introduction to modern industrial mathematics. Notices of the American Mathematical Society, 61 (4):364–375, April 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bombieri:2016:SPB

[Bom16] Enrico Bombieri. Uno spirito puro [book review, Edizioni Milella di Lecce Spazio Vivo s.r.l., Lecce, 2015]. Notices of the American Mathematical Society, 63(6):640–642, June/July 2016.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonato:2012:WLC

[Bon12] Anthony Bonato. What is … cop number? Notices of the American Mathematical Society, 59(8):1100–1101, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bondoni:2013:GFM

[Bon13] Davide Bondoni. Galois’s first memoir. Notices of the American Mathematical Society, 60 (5):598–599, May 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borwein:2012:RPS

[Bor12] Jonathan M. Borwein. Ramanujan and Pi: Srinivasa Ramanujan: Going strong at 125. Notices of the American Mathematical Society, 59 (11):1534–1537, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.
edu.au/1379/;   http://www.
ams.org/journals/notices/
201211/rtx121101522p.pdf.

Borcea:2019:IWR

[Bor19] Liliana Borcea. Imaging with waves in random media. Notices of the American Mathematical Society, 66(11):1800–1812, December 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Boute:2013:HCP

[Bou13] Raymond T. Boute. How to calculate proofs: bridging the cultural divide. Notices of the American Mathematical Society, 60(2):173–191, February 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Boyle:2013:DJR

[Boy13] Mike Boyle. Daniel J. Rudolph (1949–2010). Notices of the American Mathematical Society, 60(9):1177, October 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bray:2018:PNG

[BR18] Hubert L. Bray and Henri P. Roesch. Proof of a null geometry Penrose conjecture. Notices of the American Mathematical Society, 65(2):156–162, February 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Braden:2010:CFB

[Bra10] Lawrence S. Braden. The calculus of friendship [book review of MR2536166]. Notices of the American Mathematical Society, 57(6):743–745, June/July 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bressoud:2016:MMO

[Bre16] David M. Bressoud. The math myth and other STEM delusions [book review, The New Press, New York, 2016]. Notices of the American Mathematical Society, 63(10):1181–1183, November 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bourguignon:2010:CEH

[BRH10] Jean-Pierre Bourguignon, Reinhold Remmert, and Friedrich Hirzebruch. Cartan, Europe, and human rights. Notices of the American Mathematical Society, 57(8):972–975, September 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted from Newslet. Eur. Math. Soc., December 2008.

Brualdi:2015:HS

[Bru15]  Richard A. Brualdi.  Hans

Schneider, 1927–2014. Notices of the American Mathematical Society, 62(11):1380, December 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beals:2013:MFG

[BS13] Richard Beals and Jacek Szmigielski. Meijer G-functions: a gentle introduction. Notices of the American Mathematical Society, 60(7):866–872, August 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Budaghyan:2019:WLC

[BS19] Lilya Budaghyan and Pantelimon Stănică. What is … a cryptographic Boolean function? Notices of the American Mathematical Society, 66(1):60– 63, January 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Benkart:2010:ATS

[BSG^{+}10] Georgia Benkart, Bhama Srinivasan, Mary Gray, Ellen Maycock, and Linda Rothschild. Alice Turner Schafer (1915–2009): remembrances. Notices of the American Mathematical Society, 57(9):1116–1119, October 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Billey:2013:FDT

[BT13] Sara C. Billey and Bridget E. Tenner. Fingerprint databases for theorems. Notices of the American Mathematical Society, 60(8):1034–1039, September 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bourjaily:2018:WLA

[BT18] Jacob Bourjaily and Hugh Thomas. What is … the amplituhedron? Notices of the American Mathematical Society, 65(2): 167–169, February 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Buckley:2013:ES

[Buc13] Joseph Buckley. Ernst Snapper (1913–2011). Notices of the American Mathematical Society, 60(1):86, January 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bukiet:2018:MMB

[Buk18] Bruce G. Bukiet. Mathematically modeling baseball. Notices of the American Mathematical Society, 65(4):428–430, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Burgess:2014:WTP

[Bur14]  John P. Burgess.  Why is

there philosophy of mathematics at all? [book review of MR3222396]. Notices of the American Mathematical Society, 61(11):1345–1348, December 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beneteau:2011:DWT

[BV11] Catherine Bénéteau and Patrick J.

Van Fleet. Discrete wavelet transformations and undergraduate education. Notices of the American Mathematical Society, 58(5):656–666, May 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Byrnes:2010:TMN

[Byr10] Christopher I. Byrnes. Topological methods for nonlinear oscillations. Notices of the American Mathematical Society, 57(9): 1080–1091, October 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brent:2011:GTH

[BZ11] Richard P. Brent and Paul Zimmermann. The great trinomial hunt. Notices of the American Mathematical Society, 58(2): 233–239, February 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cain:2011:TMH

[Cai11] John W. Cain. Taking math to heart: mathematical challenges in cardiac electrophysiology. Notices of the American Mathematical Society, 58(4):542–549, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2019:ST

[Cal19a] Danny Calegari. Sturm und train. Notices of the American Mathematical Society, 66(9): 1487–1489, October 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2019:WWW

[Cal19b] Danny Calegari. Wild wild Whitehead. Notices of the American Mathematical Society, 66 (4):581–583, April 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Carman:2013:DHY

[Car13] Christián Carman. Decoding the heavens: a 2,000-year-old computer — and the centurylong search to discover its secrets [book review, Da Capo Press, Cambridge, MA, 2009]. Notices of the American Mathematical Society, 60(6):741–744, June/July 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cartier:2015:AGC

[Car15]  Pierre Cartier.  Alexander

Grothendieck: a country known only by name. Notices of the American Mathematical Society, 62(4):373–382, April 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted from Inference Int. Rev. Sci. 1 (2014), no. 1 [revised version of MR3287701].

Casselman:2010:VC

[Cas10] Bill Casselman. Visible cryptography. Notices of the American Mathematical Society, 57 (3):378–379, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/notices/201003/.

Calude:2010:WLH

[CC10] Cristian S. Calude and G. J.

Chaitin. What is … a halting probability? Notices of the American Mathematical Society, 57(2):236–237, February 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Castillo-Chavez:2016:MEM

[CCBB^{+}16] C. Castillo-Chavez, K. Barley, D. Bichara, D. Chowell, E. Diaz Herrera, B. Espinoza, V. Moreno, S. Towers, and K. E. Yong. Modeling Ebola at the Mathematical and Theoretical Biology Institute (MTBI). Notices of the American Mathematical Society, 63(4):366–371, April 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clemons:2012:WDC

[CCF^{+}12] C. Clemons, J. Cossey, M. Ferrara, S. Forcey, T. Norfolk, G. Obeng, D. Ricciardi, and G. Young. When does compromise prevent more pollution? Notices of the American Mathematical Society, 59(9): 1223–1234, October 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cartier:2010:CT

[CDD^{+}10] Pierre Cartier, Jacques Dixmier, Adrien Douady, Christian Houzel, Jean-Pierre Kahane, and Max Karoubi. Cartan as a teacher. Notices of the American Mathematical Society, 57(8):961–971, September 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clemens:2014:NIN

[CDW14] Herbert Clemens, Ingrid Daubechies, and Carol Wood. The new IMU needs you! Notices of the American Mathematical Society, 61(6): 632–534, June/July 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cannon:2011:WTG

[CF11] J. W. Cannon and W. J. Floyd.

What is … Thompson’s group? Notices of the American Mathematical Society, 58(8):1112– 1113, September 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chung:2012:IPT

[CFL^{+}12] Ping Ngai Chung, Miguel A. Fernandez, Yifei Li, Michael Mara, Frank Morgan, Isamar Rosa Plata, Niralee Shah, Luis Sordo Vieira, and Elena Wikner. Isoperimetric pentagonal tilings. Notices of the American Mathematical Society, 59(5):632–640, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2015:MAF

[CGM^{+}15] Goong Chen, Cong Gu, Philip J.

Morris, Eric G. Paterson, Alexey Sergeev, Yi-Ching Wang, and Tomasz Wierzbicki. Malaysia airlines flight MH370: water entry of an airliner. Notices of the American Mathematical Society, 62(4):330–344, April 2015.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chas:2019:ECB

[Cha19] Moira Chas. The extraordinary case of the Boole family. Notices of the American Mathematical Society, 66(11):1853–1866, December 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crowley:2011:MJW

[CHKM11] James Crowley, Susan Hezlet, Robion Kirby, and Don Mc-Clure. Mathematics journals: what is valued and what may change. Notices of the American Mathematical Society, 58 (8):1127–1130, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Report on the workshop held at MSRI, Berkeley, CA, February 14–16, 2011.

Chow:2011:WLN

[Cho11] Timothy Y. Chow. What is … a natural proof? Notices of the American Mathematical Society, 58(11):1586–1587, December 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chudnovsky:2010:HCS

[Chu10a] Maria Chudnovsky. Hadwiger’s conjecture and seagull packing. Notices of the American Mathematical Society, 57(6):733–736, June/July 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chung:2010:GTI

[Chu10b] Fan Chung. Graph theory in the information age. Notices of the American Mathematical Society, 57(6):726–732, June/July 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Connes:2016:UHM

[CJMR16] Alain Connes, Vaughan Jones, Magdalena Musat, and Mikael Rørdam. Uffe Haagerup — in memoriam. Notices of the American Mathematical Society, 63 (1):48–49, January 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Connes:2019:SMA

[CK19] Alain Connes and Joseph Kouneiher. Sir Michael Atiyah, a knight mathematician: a tribute to Michael Atiyah, an inspiration and a friend. Notices of the American Mathematical Society, 66(10):1660–1671, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2011:GGM

[CLWZ11] Goong Chen, Jaan Laane, Steven E. Wheeler, and Zhigang Zhang. Greenhouse gas molecules: a mathematical perspective. Notices of the American Mathematical Society, 58 (10):1421–1434, November 2011.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Colding:2016:LSM

[CM16] Tobias Holck Colding and William P. Minicozzi, II. Level set method for motion by mean curvature. Notices of the American Mathematical Society, 63 (10):1148–1153, November 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Colding:2019:SSG

[CM19] Tobias Holck Colding and William P. Minicozzi, II. In search of stable geometric structures. Notices of the American Mathematical Society, 66 (11):1785–1791, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coifman:2018:AHY

[CMJ^{+}18] Ronald Coifman, Stéphane Mallat, Stéphane Jaffard, Alexander Olevskii, and Albert Cohen. Ad honorem Yves Meyer. Notices of the American Mathematical Society, 65(11):1378–1396, December 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Choksi:2017:OPR

[CMT17] Rustum Choksi, Cyrill B. Muratov, and Ihsan Topaloglu. An old problem resurfaces nonlocally: Gamow’s liquid drops inspire today’s research and applications. Notices of the American Mathematical Society, 64 (11):1275–1283, December 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cohn:2017:CBS

[Coh17] Henry Cohn. A conceptual breakthrough in sphere packing. Notices of the American Mathematical Society, 64(2): 102–115, February 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/publications/journals/
notices/201702/rnoti-p102.
pdf.

Colyvan:2016:MLB

[Col16] Mark Colyvan. The mathematics of love [book review, Simon & Schuster, New York, 2015]. Notices of the American Mathematical Society, 63(7):821–822, August 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Conlon:2012:WLO

[Con12] Larry Conlon. What is … an octopus decomposition? Notices of the American Mathematical Society, 59(7):952–954, August 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Connelly:2013:WLT

[Con13] Robert Connelly. What is … a tensegrity? Notices of the American Mathematical Society, 60(1):78–80, January 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cooke:2010:LMF

[Coo10] Roger Cooke. Life on the mathematical frontier: legendary figures and their adventures. Notices of the American Mathematical Society, 57(4):464–475, April 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cooper:2012:IAA

[Coo12] S. Barry Cooper. Incomputability after Alan Turing. Notices of the American Mathematical Society, 59(6):776–784, June/ July 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

                                        www.ams.org/notices/201206/
                                        rtx120600776p.pdf.

Cornelissen:2013:LSY

[Cor13] Gunther Cornelissen. Late style — Yuri I. Manin looking back on a life in mathematics [DVD review of 2976698]. Notices of the American Mathematical Society, 60(1):84–85, January 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Corwin:2016:KPZ

[Cor16] I. Corwin. Kardar–Parisi–Zhang universality. Notices of the American Mathematical Society, 63(3):230–239, March 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2010:WLK

[CP10] Gui-Qiang G. Chen and Benoît Perthame. What is … a kinetic solution for degenerate parabolic–hyperbolic equations? Notices of the American Mathematical Society, 57(6):737–739, June/July 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chiang:2015:PTJ

[CR15] Yuan-Jen Chiang and Andrea Ratto. Paying tribute to James Eells and Joseph H, Sampson: in commemoration of the fiftieth anniversary of their pioneering work on harmonic maps. Notices of the American Mathematical Society, 62(4):388–393, April 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cordoba:2017:AHC

[CST^{+}17] Antonio Córdoba, Elias Stein, Terence Tao, Louis Nirenberg, Joseph J. Kohn, Sun-Yung Alice Chang, C. Robin Graham, Diego Córdoba, Bo’az Klartag, Jürg Fröhlich, Luis Seco, and Michael Weinstein. Ad honorem Charles Fefferman. Notices of the American Mathematical Society, 64(11):1254–1273, December 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Curtis:2012:MWE

[Cur12] Charles W. Curtis. The Mathematical Writings of Évariste Galois [book review MR2882171]. Notices of the American Mathematical Society, 59(11):1565– 1568, December 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [Neu11].

Crane:2017:GDD

[CW17] Keenan Crane and Max Wardetzky. A glimpse into discrete differential geometry. Notices of the American Mathematical Society, 64(10):1153–1159, November 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2014:OCF

[CXM^{+}14] Goong Chen, Qingang Xiong, Philip J. Morris, Eric G. Paterson, Alexey Sergeev, and Yi-Ching Wang. OpenFOAM for computational fluid dynamics. Notices of the American Mathematical Society, 61(4):354–363, April 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DAngelo:2010:BMC

[D’A10] John P. D’Angelo. Baseball and Markov chains: power hitting and power series. Notices of the American Mathematical Society, 57(4):490–495, April 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2012:TWB

[Dal12a] Andrew Dale. The theory that would not die [book review, Yale Univ. Press, New Haven, CT, 2011]. Notices of the American Mathematical Society, 59 (5):657–660, May 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2012:WLG

[Dal12b] Andrew I. Dale. What’s luck got to do with it?: The history, mathematics, and psychology behind the Gambler’s Illusion [book review; Princeton Univ. Press, Princeton, NJ, 2010]. Notices of the American Mathematical Society, 59(2):304–307, February 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2013:CPP

[Dal13a] Andrew I. Dale. Classic problems of probability [book review of MR2976816]. Notices of the American Mathematical Society, 60(10):1328–1330, November 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2013:WSR

[Dal13b] Andrew I. Dale. Who’s #1?

The science of rating and ranking. Notices of the American Mathematical Society, 60(1):81– 83, January 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2014:CIM

[Dal14a] Andrew I. Dale. A calculus of ideas: a mathematical study of human thought [book review, World Sci. Publ., Hackensack, NJ, 2012]. Notices of the American Mathematical Society, 61 (1):58–60, January 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2014:IPW

[Dal14b] Andrew I. Dale. The improbability principle: why coincidences, miracles, and rare events happen every day [book review of MR3234572]. Notices of the American Mathematical Society, 61(11):1351–1354, December 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dale:2017:OMW

[Dal17] Andrew I. Dale. Origins of mathematical words [book review of MR3136845]. Notices of the American Mathematical Society, 64(1):47–48, January 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Damerow:2010:NAD

[Dam10] Peter Damerow. From numerate apprenticeship to divine quantification [book review of Mathematics in Ancient Iraq: a social history, MR2440977]. Notices of the American Mathematical Society, 57(3):380–384, March 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dani:2019:MRM

[Dan19]  S. G. Dani.  Marina Rat-

ner [Memorial tribute]. Notices of the American Mathematical Society, 66(3):368–372, March 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Darmon:2017:AWM

[Dar17] Henri Darmon. Andrew Wiles’s marvelous proof. Notices of the American Mathematical Society, 64(3):209–216, March 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davies:2011:TPR

[Dav11] E. Brian Davies. Towards a philosophy of real mathematics [book review of MR1996199]. Notices of the American Mathematical Society, 58(10):1454– 1457, November 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:2013:LEB

[Dav13] Martin Davis. The logician and the engineer [book review of MR2954056]. Notices of the American Mathematical Society, 60(9):1170–1172, October 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dawson:2018:VCE

[Daw18] John W. Dawson, Jr. The Vienna Circle and the epic quest for the foundations of science [book review, Basic Books, New York, 2017]. Notices of the American Mathematical Society, 65(8):1002–1005, September 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

deCataldo:2010:WLP

[dCM10] Mark Andrea de Cataldo and Luca Migliorini. What is … a perverse sheaf? Notices of the American Mathematical Society, 57(5):632–634, May 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeLoera:2019:ATT

[De 19] Jesús A. De Loera. Algebraic and topological tools in linear optimization. Notices of the American Mathematical Society, 66(7):1023–1033, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeTurck:2017:MMB

[DEG^{+}17] Dennis DeTurck, Amora Elsaify, Herman Gluck, Benjamin Grossmann, Joseph Ansel Hoisington, Anusha M. Krishnan, and Jianru Zhang. Making matrices better: geometry and topology of singular value and polar decomposition. Notices of the American Mathematical Society, 64(8):

876–886, September 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dekimpe:2011:WLI

[Dek11] Karel Dekimpe. What is … an infra-nilmanifold endomorphism? Notices of the American Mathematical Society, 58 (5):688–689, May 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Denes:2011:RFJ

[Dén11] Tamás Dénes. Real face of János Bolyai. Notices of the American Mathematical Society, 58(1):41– 51, January 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Deschamp:2011:SFM

[Des11] Brent Deschamp. Street-fighting mathematics [book review of MR2583232]. Notices of the American Mathematical Society, 58(7):960–962, August 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeTurck:2017:MSB

[DeT17] Dennis DeTurck. The mathematician’s shiva [book review, Penguin Random House, New York, 2014]. Notices of the American Mathematical Society, 64(9):1043–1045, October 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ding:2014:TST

[DG14] Longyun Ding and Su Gao. Is there a spectral theory for all bounded linear operators? Notices of the American Mathematical Society, 61(7):730–735, August 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Downey:2019:CR

[DH19] Rod Downey and Denis R. Hirschfeldt. Computability and randomness. Notices of the American Mathematical Society, 66(7):1001–1012, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devadoss:2011:DBS

[DHV11] Satyan L. Devadoss, Timothy Heath, and Wasin Vipismakul. Deformations of bordered surfaces and convex polytopes. Notices of the American Mathematical Society, 58(4):530–541, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diacu:2013:MMS

[Dia13] Florin Diacu. Mathematical methods in the study of historical chronology. Notices of the American Mathematical Society, 60(4):441–449, April 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Duval:2017:PC

[DKM17] Art M. Duval, Caroline J. Klivans, and Jeremy L. Martin. The partitionability conjecture. Notices of the American Mathematical Society, 64(2):117–122, February 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Demailly:2010:CCA

[DKNS10] Jean-Pierre Demailly, Shoshichi Kobayashi, Raghavan Narasimhan, and Yum-Tong Siu. Cartan and complex analytic geometry. Notices of the American Mathematical Society, 57(8):952–960, September 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2016:CAI

[DL16a] Alexander Diaz-Lopez. Colin Adams interview. Notices of the American Mathematical Society, 63(10):1172–1174, November 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2016:HMI

[DL16b]  Alexander Diaz-Lopez.  He-

len Moore interview. Notices of the American Mathematical Society, 63(7):768–770, August 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2016:IFC

[DL16c] Alexander Diaz-Lopez. Interview with Fernando Codá Marques. Notices of the American Mathematical Society, 63(2): 142–143, February 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2016:IST

[DL16d] Alexander Diaz-Lopez. Interview with Sir Timothy Gowers. Notices of the American Mathematical Society, 63(9): 1026–1028, October 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2016:PSL

[DL16e] Alexander Diaz-Lopez. Po-Shen Loh interview. Notices of the American Mathematical Society, 63(8):905–908, September 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2017:ELI

[DL17a] Alexander Diaz-Lopez. Evelyn Lamb interview. Notices of the American Mathematical Society, 64(9):1005–1008, October 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2017:KHE

[DL17b] Alexander Diaz-Lopez. Kelsey Houston-Edwards interview. Notices of the American Mathematical Society, 64(8):870–872, September 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2018:CWI

[DL18a] Alexander Diaz-Lopez. Chelsea Walton interview. Notices of the American Mathematical Society, 65(2):164–166, February 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2018:RHI

[DL18b] Alexander Diaz-Lopez. Ruth Haas interview. Notices of the American Mathematical Society, 65(9):1084–1085, October 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaz-Lopez:2018:WSI

[DL18c] Alexander Diaz-Lopez. William Stein interview. Notices of the American Mathematical Society, 65(5):540–543, May 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeLellis:2019:TGN

[DL19] Camillo De Lellis and László Székelyhidi, Jr. On turbulence and geometry: from Nash to Onsager. Notices of the American Mathematical Society, 66 (5):677–685, May 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

delaPena:2019:DSN

[dlP19] Victor H. de la Pena. From decoupling and self-normalization to machine learning. Notices of the American Mathematical Society, 66(10):1641–1646, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Daniels:2017:WLE

[DLR17] Harris B. Daniels and Álvaro Lozano-Robledo. What is … an elliptic curve? Notices of the American Mathematical Society, 64(3):241–243, March 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dacorogna:2010:OPD

[DMP10] Bernard Dacorogna, Paolo Marcellini, and Emanuele Paolini. Origami and partial differential equations. Notices of the American Mathematical Society, 57 (5):598–606, May 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

deMelo:2015:WFM

[dMPQZ15] Welington de Melo, Bjorn Poonen, Jeremy Quastel, and Anton Zorich. The work of the 2014 Fields medalists. Notices of the American Mathematical Society, 62(11):1334–1349, December 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Don:2010:MBS

[DMVW10] Gary W. Don, Karyn K. Muir, Gordon B. Volk, and James S. Walker. Music: broken symmetry, geometry, and complexity. Notices of the American Mathematical Society, 57(1):30– 49, January 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Donaldson:2011:WLN

[Don11]  Simon Donaldson.  On the

work of Louis Nirenberg. Notices of the American Mathematical Society, 58(3):469–472, March 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Donovan:2014:ATM

[Don14] Peter W. Donovan. Alan Turing, Marshall Hall, and the alignment of WW2 Japanese Naval intercepts. Notices of the American Mathematical Society, 61(3):258–264, March 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201403/
rnoti-p258.pdf.

Donoho:2018:BBH

[Don18] David Donoho. From blackboard to bedside: high-dimensional geometry is transforming the MRI industry. Notices of the American Mathematical Society, 65 (1):40–44, January 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Donaldson:2019:KUC

[Don19] Simon Donaldson. Karen Uhlenbeck and the calculus of variations. Notices of the American Mathematical Society, 66 (3):303–313, March 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Duran:2014:MTM

[DPV14] Antonio J. Durán, Mario Pérez, and Juan L. Varona. The misfortunes of a trio of mathematicians using computer algebra systems. Can we trust in them? Notices of the American Mathematical Society, 61(10):1249–1252, November 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

                                        www.ams.org/notices/201410/
                                        rnoti-p1249.pdf.

Drasin:2015:ABI

[Dra15a] David Drasin. Albert Baernstein II 1941–2014. Notices of the American Mathematical Society, 62(7):815–818, August 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Drasin:2015:MLF

[Dra15b] David Drasin. My life and functions [book review, Logic Press, 2014]. Notices of the American Mathematical Society, 62 (5):517–520, May 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Damron:2016:RGM

[DRAS16] Michael Damron, Firas Rassoul-Agha, and Timo Seppäläinen. Random growth models. Notices of the American Mathematical Society, 63(9):1004–1008, October 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2018:IAL

[DS18] Bjørn Ian Dundas and Christian Skau. Interview with Abel Laureate Yves Meyer [reprint of MR3726777]. Notices of the American Mathematical Society, 65(5):520–529, May 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2019:IAL

[DS19] Bjørn Ian Dundas and Christian Skau. Interview with Abel Laureate Robert P. Langlands. Notices of the American Mathematical Society, 66(4):494–503, April 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted from the Newsletter of European Mathematical Society, September 2018.

Dumbaugh:2017:WIL

[DSG17] Della Dumbaugh and Amy Shell-Gellasch. The “wide influence” of Leonard Eugene Dickson. Notices of the American Mathematical Society, 64(7):772–776, August 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Demeyer:2016:BBB

[DSW16] Jeroen Demeyer, William Stein, and Ursula Whitcher. Beyond the black box. Notices of the American Mathematical Society, 63(8):928–929, September 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DAngelo:2010:ICR

[DT10] John P. D’Angelo and Jeremy T. Tyson. An invitation to Cauchy– Riemann and sub-Riemannian geometries. Notices of the American Mathematical Society, 57 (2):208–219, February 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dudley:2010:WM

[Dud10] Underwood Dudley. What is mathematics for? Notices of the American Mathematical Society, 57(5):608–613, May 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dujella:2016:WLD

[Duj16] Andrej Dujella. What is … a Diophantine m-tuple? Notices of the American Mathematical Society, 63(7):772–774, August 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Duke:2014:ACA

[Duk14] W. Duke. Almost a century of answering the question: what is a mock theta function? Notices of the American Mathematical Society, 61(11):1314–1320, December 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2019:COB

[Dum19a] Della Dumbaugh. Creating opportunities and building confidence: Clare Boothe Luce’s unexpected support of women in math and science. Notices of the American Mathematical Society, 66(3):387–394, March 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2019:ECR

[Dum19b] Della Dumbaugh. Extensive cooperation with rugged individualism: George Mackey’s guide for practitioners of mathematics. Notices of the American Mathematical Society, 66(6):883–891, June/July 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2019:BPB

[Dum19c] Della Dumbaugh. From breakthrough prizes to breakout graduate fellowships: ‘mathematics is a community’. Notices of the American Mathematical Society, 66(8):1294–1297, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2019:MGA

[Dum19d] Della Dumbaugh. Maria Gaetana Agnesi: 1718–1799. Notices of the American Mathematical Society, 66(3):414–415, March 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2019:RLH

[Dum19e] Della Dumbaugh. Rudy Lee Horne: the hidden figure of Hidden Figures, 1968–2017. Notices of the American Mathematical Society, 66(2):202–210, February 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2019:ERPa

[Dun19a] Edward Dunne. Everything in its right place an expert guide to searching with Math-SciNet. Notices of the American Mathematical Society, 66 (8):1320–1324, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2019:ERPb

[Dun19b] Edward Dunne. Everything in its right place: Part II: An expert guide to searching with Math-SciNet. Notices of the American Mathematical Society, 66(9): 1501–1506, October 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2019:LML

[Dun19c] Edward Dunne. Looking at the mathematics literature. Notices of the American Mathematical Society, 66(2):227–230, February 2019. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Dupont:2012:WSC

[Dup12] Johan L. Dupont. What is a scissors congruence? Notices of the American Mathematical Society, 59(9):1242–1244, October 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Durrett:2013:CMP

[Dur13] Rick Durrett. Cancer modeling: a personal perspective. Notices of the American Mathematical Society, 60(3):304–309, March 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dyson:2013:WTJ

[Dys13] Freeman Dyson. A walk through Johnny von Neumann’s garden. Notices of the American Mathematical Society, 60(2):154–161, February 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://www.ams.org/notices/
201302/rnoti-p154.pdf. Talk

given at Brown University, Providence, Rhode Island, May 4, 2010.

E:2019:MLM

[E19] Weinan E. Machine learning:

mathematical theory and scientific applications. Notices of the American Mathematical Society, 66(11):1813–1820, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Edelman:2013:BPR

[Ede13] Paul H. Edelman. Burden of proof: a review of Math on Trial [book review of MR3025050]. Notices of the American Mathematical Society, 60(7):910–914, August 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Edwards:2012:GCR

[Edw12] Harold M. Edwards. Galois for 21st-century readers. Notices of the American Mathematical Society, 59(7):912–923, August 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eisworth:2010:MOV

[Eis10] Todd Eisworth. Mathematicians: an outer view of the inner world [book review of MR2542778]. Notices of the American Mathematical Society, 57(7):861–863, August 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Elran:2019:TTT

[Elr19] Yossi Elran. Twists, tilings, and tessellations [book review of MR3753635]. Notices of the American Mathematical Society, 66(8):1326–1329, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ezhov:2011:CTG

[EMS11] Vladimir Ezhov, Ben McLaughlin, and Gerd Schmalz. From Cartan to Tanaka: getting real in the complex world. Notices of the American Mathematical Society, 58(1):20–27, January 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ewing:2011:WMC

[Ewi11a] John Ewing. I want to be a mathematician: a conversation with Paul Halmos [film review, Math. Assn. of Amer., Washington, DC, 2009]. Notices of the American Mathematical Society, 58(6):810–811, June/July 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ewing:2011:MID

[Ewi11b] John Ewing. Mathematical intimidation: driven by the data. Notices of the American Mathematical Society, 58(5):667–673, May 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ewing:2012:UGU

[Ewi12]  John Ewing.  Uneducated

guesses: using evidence to uncover misguided education policies [book review, Princeton Univ. Press, Princeton, NJ, 2011]. Notices of the American Mathematical Society, 59(6): 815–818, June/July 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Faris:2012:BIE

[Far12a] William G. Faris. The beginning of infinity: Explanations that transform the world [book review, Viking Penguia, New York, 2011]. Notices of the American Mathematical Society, 59 (4):544–552, April 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Farris:2012:FS

[Far12b] Frank A. Farris. Forbidden symmetries. Notices of the American Mathematical Society, 59 (10):1386–1390, November 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Feldman:2014:PAC

[Fel14] Marcus Feldman. Probably approximately correct [book review, Basic Books, New York, 2013]. Notices of the American Mathematical Society, 61(10):1222– 1223, November 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fenyvesi:2012:MAB

[Fen12] Kristóf Fenyvesi. Mathematical aesthetics in a beautiful town: Bridges Coimbra 2011. Notices of the American Mathematical Society, 59(5):648–652, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Held at the University of Coimbra, July 27– 31, 2011.

Ferguson:2010:CMS

[FF10] Helaman Ferguson and Claire Ferguson. Celebrating mathematics in stone and bronze. Notices of the American Mathematical Society, 57(7):840–850, August 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fefferman:2012:PLA

[FF12] Charles Fefferman and Robert Fefferman. Princeton lectures in analysis [book reviews of MR1970295, MR1976398, MR2129625, MR2827930]. Notices of the American Mathematical Society, 59(5):641–647, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions from Paul Hagelstein, Nataša Pavlović and Lillian Pierce.

Forsen:2013:WSW

[FGLG13] Sture Forsén, Harry B. Gray, L. K. Olof Lindgren, and Shirley B. Gray. Was something wrong with Beethoven’s metronome? Notices of the American Mathematical Society, 60(9):1146–1153, October 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ferreira:2011:ESS

[FH11] Paulo J. S. G. Ferreira and Rowland Higgins. The establishment of sampling as a scientific principle — a striking case of multiple discovery. Notices of the American Mathematical Society, 58(10):1446–1450, November 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Flapan:2017:KSI

[Fla17] Laure Flapan. Karen E. Smith interview. Notices of the American Mathematical Society, 64 (7):718–720, August 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Flapan:2018:JHI

[Fla18] Laure Flapan. Joe Harris interview. Notices of the American Mathematical Society, 65(6): 661–662, June/July 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Floystad:2015:EAC

[Flø15] Gunnar Fløystad. The exterior algebra and central notions in mathematics. Notices of the American Mathematical Society, 62(4):364–371, April 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Folsom:2010:WLM

[Fol10] Amanda Folsom. What is … a mock modular form? Notices of the American Mathematical Society, 57(11):1441–1443, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Folland:2011:BWM

[Fol11] Gerald B. Folland. The best writing on mathematics 2010[book review of MR2777703]. Notices of the American Mathematical Society, 58(10):1451–1452, November 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Folland:2012:UZW

[Fol12a] Gerald B. Folland. The universe in zero words and In pursuit of the unknown [book review of MR2908506; MR2896048]. Notices of the American Mathematical Society, 59(11):1562–1564, December 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Folland:2012:WBM

[Fol12b] Gerald B. Folland. Why beliefs matter: reflections on the nature of science[book review of MR2674693]. Notices of the American Mathematical Society, 59(4):553–559, April 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Folland:2016:BFS

[Fol16] Gerald B. Folland. Birds and frogs: selected papers, 1990–2014 [book review of MR3287223]. Notices of the American Mathematical Society, 63(4):442–443, April 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Foreman:2018:WBR

[For18] Matthew Foreman. What is a Borel reduction? Notices of the American Mathematical Society, 65(10):1263–1268, November 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fox:2012:WLA

[Fox12] Daniel J. F. Fox. What is … an affine sphere? Notices of the American Mathematical Society, 59(3):420–423, March 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Franzoni:2012:KBV

[Fra12] Gregorio Franzoni. The Klein bottle: variations on a theme. Notices of the American Mathematical Society, 59(8):1076– 1082, September 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fresan:2010:PRM

[Fre10a] Javier Fresán. Pythagoras’ revenge: a mathematical mystery [book review, Princeton University Press, Princeton, NJ, 2009]. Notices of the American Mathematical Society, 57(5):637–638, May 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Frey:2010:ABC

[Fre10b] Gerhard Frey. The arithmetic behind cryptography. Notices of the American Mathematical Society, 57(3):366–374, March 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.

ams.org/notices/201003/.

Frenkel:2018:AEP

[Fre18] Edward Frenkel. AMS Einstein public lecture: imagination and knowledge. Notices of the American Mathematical Society, 65 (4):410–412, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedman:2010:WMB

[Fri10] Avner Friedman. What is mathematical biology and how useful is it? Notices of the American Mathematical Society, 57 (7):851–857, August 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedlander:2015:PNM

[Fri15] John Friedlander. Prime numbers: a much needed gap is finally found. Notices of the American Mathematical Society, 62(6):660–664, June/July 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedberg:2018:WIL

[Fri18] Solomon Friedberg. What is … the Langlands program? Notices of the American Mathematical Society, 65(6):663–665, June/July 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fujiwara:2010:HPB

[Fuj10] Koji Fujiwara. The housekeeper and the professor [book review, Picadoz, London, 2009]. Notices of the American Mathematical Society, 57(5):635–636, May 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Furman:2011:WLS

[Fur11] Alex Furman. What is … a stationary measure? Notices of the American Mathematical Society, 58(9):1276–1277, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gaboriau:2010:WLC

[Gab10] Damien Gaboriau. What is … cost? Notices of the American Mathematical Society, 57 (10):1295–1296, November 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gangbo:2018:OPC

[Gan18] Wilfrid Gangbo. Open problems concerning Michell trusses. Notices of the American Mathematical Society, 65(3):313–316, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garibaldi:2010:WLL

[Gar10] Skip Garibaldi. What is … a linear algebraic group? Notices of the American Mathematical Society, 57(9):1125–1126, October 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garcia:2017:WLC

[Gar17] Stephan Ramon Garcia. What is … a complex symmetric operator? Notices of the American Mathematical Society, 64(1):35– 37, January 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garfunkel:2019:MCG

[Gar19] Solomon Garfunkel. Modeling competitions and gender equity. Notices of the American Mathematical Society, 66(10):1685– 1688, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goodstein:2012:FLF

[GB12] Judith Goodstein and Donald Babbitt. A fresh look at Francesco Severi. Notices of the American Mathematical Society, 59(8):1064–1075, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201208/
rtx120801064p.pdf.

Goodstein:2013:BMC

[GB13] Judith R. Goodstein and Donald Babbitt. E. T. Bell and mathematics at Caltech between the wars. Notices of the American Mathematical Society, 60(6): 686–698, June/July 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garcia:2018:RDS

[GCCUD18] Stephan Ramon Garcia, Sun-Yung Alice Chang, David Cruz-Uribe, and John Doyle. Remembering Donald Sarason (1933– 2017). Notices of the American Mathematical Society, 65(2): 195–200, February 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gray:2015:MAP

[GDG^{+}15] Shirley B. Gray, Daniel Ye Ding, Gustavo Gordillo, Samuel Landsberger, and Cye Waldman. The method of Archimedes: propositions 13 and 14. Notices of the American Mathematical Society, 62(9):1036–1040, October 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gerovitch:2016:IHD

[Ger16] Slava Gerovitch. Infinitesimal: how a dangerous mathematical theory shaped the modern world [book review, Sci. Am./Farrar, Straus and Giroux, New York, 2014]. Notices of the American Mathematical Society, 63 (5):571–574, May 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gerhardt:2019:IRE

[Ger19]  Teena Gerhardt.  Invariants

of rings via equivariant homotopy. Notices of the American Mathematical Society, 66 (8):1353–1354, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Geelen:2014:SRC

[GGW14] Jim Geelen, Bert Gerards, and Geoff Whittle. Solving Rota’s conjecture. Notices of the American Mathematical Society, 61 (7):736–743, August 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garcia:2015:GHM

[GHL15] Stephan Ramon Garcia, Trevor Hyde, and Bob Lutz. Gauss’s hidden menagerie: from cyclotomy to supercharacters. Notices of the American Mathematical Society, 62(8):878–888, September 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ghomi:2018:DUP

[Gho18] Mohammad Ghomi. Dürer’s unfolding problem for convex polyhedra. Notices of the American Mathematical Society, 65(1):25– 27, January 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Graham:2018:MGB

[GHPW18] Erica Graham, Raegan Higgins, Candice Price, and Shelby Wilson. The Mathematically Gifted and Black website. Notices of the American Mathematical Society, 65(2):124–126, February 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Giv:2019:WLH

[Giv19] Hossein Hosseini Giv. What is … a hereditarily indecomposable Banach space? Notices of the American Mathematical Society, 66(9):1474–1475, October 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gurski:2019:TIH

[GJO19] Nick Gurski, Niles Johnson, and Angélica M. Osorno. Topological invariants from higher categories. Notices of the American Mathematical Society, 66 (8):1225–1237, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ge:2019:MRV

[GJRR19] Liming Ge, Arthur Jaffe, Marc Rieffel, and Mikael Rørdam. In memoriam: Richard V. Kadison (1925–2018). Notices of the American Mathematical Society, 66(9):1453–1463, October 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goemans:2011:MWD

[GK11] Michel X. Goemans and Jonathan A.

Kelner. The mathematical work of Daniel Spielman. Notices of the American Mathematical Society, 58(9):1288–1291, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldberg:2018:MAR

[GK18] David Goldberg and Phil Kutzko.

The Math Alliance and its roots in the African American community. Notices of the American Mathematical Society, 65(2): 127–131, February 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Glass:2013:CET

[Gla13] Darren B. Glass. Coble and Eisenhart: two Gettysburgians who led mathematics. Notices of the American Mathematical Society, 60(5):558–566, May 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Glasscock:2015:WLG

[Gla15] Daniel Glasscock. What is … a graphon? Notices of the American Mathematical Society, 62(1):46–48, January 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Glutsyuk:2014:NIT

[Glu14] Alexey Glutsyuk. Naming infinity: a true story of religious mysticism and mathematical creativity [book review of MR2526973]. Notices of the American Mathematical Society, 61(1):62–65, January 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grieser:2013:HST

[GM13] Daniel Grieser and Svenja Maronna. Hearing the shape of a triangle. Notices of the American Mathematical Society, 60 (11):1440–1447, December 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldfeld:2019:ASW

[GM19] Dorian Goldfeld and Simone Munao. Alexander Smith wins the first David Goss Prize in number theory. Notices of the American Mathematical Society, 66 (11):1875–1878, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goff:2015:APM

[Gof15]  Christopher Goff.  Animat-

ing popular mathematics: “The Simpsons and their mathematical secrets” [book review of MR3155039]. Notices of the American Mathematical Society, 62(1):40–44, January 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goins:2018:TQJ

[Goi18] Edray Herber Goins. Three questions: the journey of one black mathematician. Notices of the American Mathematical Society, 65(2):144–147, February 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goins:2019:UEC

[Goi19] Edray Herber Goins. The ubiquity of elliptic curves. Notices of the American Mathematical Society, 66(2):169–174, February 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldberg:2019:BWB

[Gol19]  Lisa R. Goldberg.  The

book of why [book review of MR3791446]. Notices of the American Mathematical Society, 66(7):1093–1098, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:2011:WNL

[Grä11] George Grätzer. What is new in LaTEX? IV. WYSIWYG LaTEX. Notices of the American Mathematical Society, 58(6):828–830, June/July 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://www.ams.org/notices/
201106/rtx110600828p.pdf.

Grabiner:2013:SLB

[Gra13a]  Judith V. Grabiner.  Se-

duced by logic [book review of MR3013691]. Notices of the American Mathematical Society, 60(6):744–745, June/July 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:2013:WNL

[Grä13b]  George Grätzer.  What is

new in LaTEX? V. LaTEX on an iPad. foundation. Notices of the American Mathematical Society, 60(3):332–334, March 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201303/
rnoti-p332.pdf.

Gray:2016:WWW

[Gra16] Jeremy Gray. Who would have won the Fields Medal 150 years ago? Notices of the American Mathematical Society, 63 (3):269–274, March 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grabiner:2017:RCB

[Gra17] Judith V. Grabiner. The real and the complex [book review of MR3380969]. Notices of the American Mathematical Society, 64(11):1322–1326, December 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grcar:2010:TBG

[Grc10] Joseph F. Grcar. Topical bias in generalist mathematics journals. Notices of the American Mathematical Society, 57 (11):1421–1424, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grcar:2011:MGE

[Grc11] Joseph F. Grcar. Mathematicians of Gaussian elimination. Notices of the American Mathematical Society, 58(6):782–792, June/July 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grcar:2013:ECM

[Grc13] Joseph F. Grcar. Errors and corrections in mathematics literature. Notices of the American Mathematical Society, 60 (4):418–425, April 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Green:2012:WLA

[Gre12a] Ben Green. What is … an approximate group? Notices of the American Mathematical Society, 59(5):655–656, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Green:2012:FRT

[Gre12b] Judy Green. Film review: Top secret Rosies: The female computers of World War II [website:

http://www.topsecretrosies.com].

Notices of the American Mathematical Society, 59(2):308–311, February 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gross:2018:WLR

[Gro18] Jacob Gross. What is … Riemannian holonomy? Notices of the American Mathematical Society, 65(7):795–796, August 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goodman-Strauss:2010:CDU

[GS10] Chaim Goodman-Strauss. Can’t decide? Undecide! Notices of the American Mathematical Society, 57(3):343–356, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Greenwald:2015:MTO

[GT15] Sarah J. Greenwald and Jill E. Thomley. The mαth βook [book review of MR3183832]. Notices of the American Mathematical Society, 62(4):384–385, April 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldstein:2010:WDF

[GTTZ10] Sheldon Goldstein, Daniel V. Tausk, Roderich Tumulka, and Nino Zanghì. What does the free will theorem actually prove? Notices of the American Mathematical Society, 57(11):1451–1453, December 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Guillen:2019:PGJ

[Gui19] Nestor Guillen. A primer on generated Jacobian equations: geometry, optics, economics. Notices of the American Mathematical Society, 66(9):1401–1411, October 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gundy:2010:TSF

[Gun10] Richard F. Gundy. Tilings, scaling functions, and a Markov process. Notices of the American Mathematical Society, 57(9): 1094–1104, October 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gross:2015:MRK

[GW15] Kenneth I. Gross and Edward N.

Wilson. In memory of Ray Kunze. Notices of the American Mathematical Society, 62 (10):1190, November 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldbring:2019:INA

[GW19] Isaac Goldbring and Sean Walsh.

An invitation to nonstandard analysis and its recent applications. Notices of the American Mathematical Society, 66(6): 842–851, June/July 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Haggstrom:2011:NSH

[Häg11] Olle Häggström. Nonsense on stilts: how to tell science from bunk [book review, University of Chicago Press, Chicago, 2010]. Notices of the American Mathematical Society, 58(4):582–584, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Haggstrom:2012:CSB

[Häg12] Olle Häggström. The calculus of selfishness [book review of MR2590059]. Notices of the American Mathematical Society, 59(1):47–49, January 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Haggstrom:2013:PPT

[Häg13] Olle Häggström. Paradoxes in probability theory [book review of MR2963841]. Notices of the American Mathematical Society, 60(3):329–331, March 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hales:2014:NBD

[Hal14] Thomas C. Hales. The NSA back door to NIST. Notices of the American Mathematical Society, 61(2):190–192, February 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hamblin:2019:SBM

[Ham19] Elizabeth Platt Hamblin. Sharing the beauty of mathematics. Notices of the American Mathematical Society, 66 (8):1312–1318, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hannah:2012:MNF

[Han12] John Hannah. The man of numbers: Fibonacci’s arithmetic revolution [book review, Walker and Co., New York, NY, 2011]. Notices of the American Mathematical Society, 59(5):661–664, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hasselblatt:2019:AKH

[Has19] Boris Hasselblatt. Anatole Katok — a half-century of dynamics. Notices of the American Mathematical Society, 66 (5):708–719, May 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hayes:2013:FTB

[Hay13] Brian Hayes. Figures of thought [book review, Green Lion Press, Santa Fe, NM, 2005]. Notices of the American Mathematical Society, 60(9):1173–1176, October 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hayes:2014:DDS

[Hay14] Brian Hayes. Doing data science [book review, O’Reilly Media, 2014]. Notices of the American Mathematical Society, 61(9):

1068–1071, October 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hayes:2017:MVE

[Hay17] Brian Hayes. The mathematics of various entertaining subjects [book review of MR3444773]. Notices of the American Mathematical Society, 64(2):163–166, February 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Heil:2013:WLF

[Hei13] Christopher Heil. What is … a frame? Notices of the American Mathematical Society, 60(6): 748–750, June/July 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Helgason:2013:GML

[Hel13a]  Sigurdur Helgason.  Gösta

Mittag-Leffler. A man of conviction [book review of MR2743189]. Notices of the American Mathematical Society, 60(8):1054– 1055, September 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Heller:2013:DQN

[Hel13b] Michael Heller. Deep questions on the nature of mathematics [book review of Meaning in mathematics, MR2895623]. Notices of the American Mathematical Society, 60(5):592–594, May 2013. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Henson:2010:MAB

[HH10] Shandelle M. Henson and James L. Hayward. The mathematics of animal behavior: an interdisciplinary dialogue. Notices of the American Mathematical Society, 57(10):1248–1258, November 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hoshino:2013:GTS

[HiK13] Richard Hoshino and Ken ichi Kawarabayashi. Graph theory and sports scheduling. Notices of the American Mathematical Society, 60(6):726–731, June/July 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hitchin:2011:SIS

[Hit11] Nigel Hitchin. The shape of inner space [book review of MR2722198]. Notices of the American Mathematical Society, 58(2):311–312, February 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harris:2019:LLH

[HLS19] Pamela E. Harris, Alicia Prieto Langarica, and Luis Sordo Vieira. 2019 Lathisms: Latinxs and Hispanics in the mathematical sciences. Notices of the American Mathematical Society, 66(8):1285–1289, September 2019. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Hill:2017:HGL

[HN17] C. Denson Hill and Paweł Nurowski. How the green light was given for gravitational wave search. Notices of the American Mathematical Society, 64 (7):686–692, August 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Paper title also given as: How the green light was given for gravitational wave research.

Hughes-Oliver:2019:APA

[HO19] Jacqueline M. Hughes-Oliver. Assessment of prediction algorithms for ranking objects. Notices of the American Mathematical Society, 66(2):182–190, February 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hodge:2011:BRN

[Hod11]  Jonathan K. Hodge.  Book

review: Numbers rule: the vexing mathematics of democracy, from Plato to the present [MR2604887]. Notices of the American Mathematical Society, 58(1):59–61, January 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hollings:2013:SAI

[Hol13] Christopher Hollings. The struggle against idealism: Soviet ideology and mathematics. Notices of the American Mathematical Society, 60(11):1448–1458, December 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Holm:2016:WIS

[Hol16] Tara S. Holm. What is … symplectic geometry. Notices of the American Mathematical Society, 63(11):1252–1254, December 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Honigs:2016:MG

[Hon16]  Katrina Honigs.  Meeting

Grothendieck, 2012. Notices of the American Mathematical Society, 63(3):266–267, March 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hosack:2019:IPM

[Hos19] John Hosack. An inclusive philosophy of mathematics. Notices of the American Mathematical Society, 66(9):1433–1437, October 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2018:CFS

[Hou18a] Johnny L. Houston. Clarence F. Stephens (1917–2018). Notices of the American Mathematical Society, 65(7):848–849, August 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2018:TAA

[Hou18b] Johnny L. Houston. Ten African American pioneers and mathematicians who inspired me. Notices of the American Mathematical Society, 65(2):139–143, February 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2019:CYR

[Hou19a] Johnny L. Houston. A centennial year (2019) reflection on the life and contributions of mathematician David H. Blackwell (1919– 2010). Notices of the American Mathematical Society, 66(2): 221–226, February 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2019:LPC

[Hou19b] Johnny L. Houston. The life and pioneering contributions of an African American centenarian: mathematician Katherine G. Johnson. Notices of the American Mathematical Society, 66(3):324–329, March 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harrington:2017:WLC

[HR17] Phillip S. Harrington and Andrew Raich. What is … a CR submanifold? Notices of the American Mathematical Society, 64(7):722–724, August 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Heilman:2010:LEH

[HS10] Steven M. Heilman and Robert S.

Strichartz. Localized eigenfunctions: here you see them, there you don’t. Notices of the American Mathematical Society, 57 (5):624–629, May 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hogben:2019:JMR

[HS19] Leslie Hogben and T. Christine Stevens. Joining a mathematical research community. Notices of the American Mathematical Society, 66(7):1101–1107, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Helgason:2015:RKP

[HSPM15] Sigurdur Helgason, Dana Scott, Alice Peters, and David Mumford. Remembering Klaus Peters. Notices of the American Mathematical Society, 62(3): 264–268, March 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Huber:2019:TGI

[Hub19] Mark Huber. Ten great ideas about chance [book review of MR3702017]. Notices of the American Mathematical Society, 66(6):917–921, June/July 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hurtubise:2014:TVB

[Hur14] Jacques Hurtubise. Théorème vivant [book review, Grasset et Fasquelle, Paris, 2012]. Notices of the American Mathematical Society, 61(2):172–173, February 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hales:2011:MWF

[HWWA11] Thomas C. Hales, Benjamin Weiss, Wendelin Werner, and Luigi Ambrosio. The mathematical work of the 2010 Fields medalists. Notices of the American Mathematical Society, 58 (3):453–458, March 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hynd:2019:SPD

[Hyn19] Ryan Hynd. Sticky particle dynamics on the real line. Notices of the American Mathematical Society, 66(2):162–168, February 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Illusie:2010:RGH

[IBB^{+}10] Luc Illusie, Alexander Beilinson, Spencer Bloch, Vladimir Drinfeld, et al. Reminiscences of Grothendieck and his school. Notices of the American Mathematical Society, 57(9):1106–1115, October 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Illusie:2019:MR

[Ill19] Luc Illusie. Michel Raynaud (1938–2018). Notices of the American Mathematical Society, 66(1):37–41, January 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Iosevich:2019:WLF

[Ios19] Alex Iosevich. What is … Falconer’s conjecture? Notices of the American Mathematical Society, 66(4):552–555, April 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jabbusch:2019:WMR

[Jab19] Kelly Jabbusch. Writing for Mathematical Reviews(R). Notices of the American Mathematical Society, 66(2):197–198, February 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2010:KRN

[Jac10] Allyn Jackson. Kalman receives National Medal of Science. Notices of the American Mathematical Society, 57(1):56–57, January 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2011:PVI

[Jac11]  Allyn Jackson.  Presidential

views: interview with George Andrews. Notices of the American Mathematical Society, 58(2):

318–320, February 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2013:PVIb

[Jac13a]  Allyn Jackson.  Presidential

views: interview with David Vogan. Notices of the American Mathematical Society, 60(3): 336–338, March 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2013:PVIa

[Jac13b]  Allyn Jackson.  Presidential

views: interview with Eric Friedlander. Notices of the American Mathematical Society, 60(2): 218–220, February 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2014:HCS

[Jac14] Allyn Jackson. Happy 91st, Cathleen Synge Morawetz. Notices of the American Mathematical Society, 61(5):510–511, May 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2016:APR

[Jac16] Allyn Jackson. AMS President Robert Bryant: his early life and his views on mathematics. Notices of the American Mathematical Society, 63(6):614–620, June/July 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2017:MAM

[Jac17] Allyn Jackson. Michèle Audin, mathematician and writer. Notices of the American Mathematical Society, 64(7):761–762, August 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ji:2016:CMI

[Ji16] Lizhen Ji. Chinese mathematics and ICCM. Notices of the American Mathematical Society, 63 (7):780–782, August 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Johnson:2014:JBT

[Joh14] Dale M. Johnson. L. E. J.

Brouwer — topologist, intuitionist, philosopher: how mathematics is rooted in life [book review of MR2905809]. Notices of the American Mathematical Society, 61(6):607–610, June/July 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jones:2012:NVS

[Jon12] Rafe Jones. Numbers: a very short introduction [book review of MR2848884]. Notices of the American Mathematical Society, 59(1):50–51, January 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jones:2016:MPE

[Jon16] Christopher K. R. T. Jones.

Mathematics of planet earth: mathematicians reflect on how to discover, organize, and protect our planet [book review of MR3486163]. Notices of the American Mathematical Society, 63(11):1276–1277, December 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jones:2019:SMP

[Jon19] Rafe Jones. A singular mathematical promenade [book review of MR3702027]. Notices of the American Mathematical Society, 66(11):1867–1870, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jarnicki:2011:DRV

[JP11] Marek Jarnicki and Peter Pflug. Directional regularity vs. joint regularity. Notices of the American Mathematical Society, 58 (7):896–904, August 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jacobson:2014:ISI

[JS14] M. J. Jacobson, Jr. and R. Scheidler. Infrastructure: structure inside the class group of a real quadratic field. Notices of the American Mathematical Society, 61(1):36–46, January 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kalai:2016:QCP

[Kal16] Gil Kalai. The quantum computer puzzle. Notices of the American Mathematical Society, 63(5):508–516, May 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Karigiannis:2011:WLM

[Kar11] Spiro Karigiannis. What is … a G_{2}-manifold? Notices of the American Mathematical Society, 58(4):580–581, April 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Katz:2017:WLT

[Kat17] Eric Katz. What is … tropical geometry? Notices of the American Mathematical Society, 64(4):380–382, April 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2010:PRO

[Keh10] Elaine Kehoe. Popa receives Ostrowski Prize. Notices of the American Mathematical Society, 57(3):404–405, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2011:MRN

[Keh11] Elaine Kehoe. Mumford receives National Medal of Science. Notices of the American Mathematical Society, 58(2):317, February 2011. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:ASB

[Keh12a] Elaine Kehoe. 2012 AMS–SIAM Birkhoff Prize. Notices of the American Mathematical Society, 59(4):576–577, April 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:CPA

[Keh12b] Elaine Kehoe. 2012 Cole Prize in algebra. Notices of the American Mathematical Society, 59(4): 567, April 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:CP

[Keh12c] Elaine Kehoe. 2012 Conant Prize. Notices of the American Mathematical Society, 59(4): 568, April 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:MP

[Keh12d] Elaine Kehoe. 2012 Morgan Prize. Notices of the American Mathematical Society, 59 (4):569–571, April 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:SP

[Keh12e] Elaine Kehoe. 2012 Steele Prizes. Notices of the American Mathematical Society, 59(4):563–566, April 2012. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:WP

[Keh12f] Elaine Kehoe. 2012 Whiteman Prize. Notices of the American Mathematical Society, 59 (4):572–575, April 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:ADP

[Keh12g] Elaine Kehoe. Award for distinguished public service. Notices of the American Mathematical Society, 59(4):575–576, April 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:BTA

[Keh12h] Elaine Kehoe. Bourgain and Tao awarded 2012 Crafoord Prize. Notices of the American Mathematical Society, 59(5):676–677, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2012:TVR

[Keh12i] Elaine Kehoe. Tapia and Varadhan receive National Medal of Science. Notices of the American Mathematical Society, 59(1):52– 54, January 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2013:ACA

[Keh13a] Elaine Kehoe. Aschbacher and Caffarelli awarded 2012 Wolf Prize. Notices of the American Mathematical Society, 60 (4):474–475, April 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2013:DMR

[Keh13b] Elaine Kehoe. Daubechies and Mumford receive BBVA Foundation Award. Notices of the American Mathematical Society, 60(5):603–604, May 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201305/
rnoti-p603.pdf.

Kehoe:2016:HAS

[Keh16] Elaine Kehoe. Hitchin awarded Shaw Prize. Notices of the American Mathematical Society, 63(8):940–941, September 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2018:BDA

[Keh18] Elaine Kehoe. Beilinson and Drinfeld awarded 2018 Wolf Prize in mathematics. Notices of the American Mathematical Society, 65(6):697–698, June/ July 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kjos-Hanssen:2019:OHB

[KH19] Bjørn Kjos-Hanssen. Only human [book review of The Turing guide, MR3618873]. Notices of the American Mathematical Society, 66(4):556–561, April 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khammash:2016:SGK

[Kha16] Ahmed A. Khammash. Sandy Green as I knew him. Notices of the American Mathematical Society, 63(6):657–658, June/July 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Keen:2018:CJE

[KK18] Linda Keen and Irwin Kra. Clifford J. Earle: a life in mathematics and music (1935–2017). Notices of the American Mathematical Society, 65(1):52–53, January 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kuznetsov:2014:LVA

[KKK^{+}14] Nikolay Kuznetsov, Tadeusz Kulczycki, Mateusz Kwaśnicki, Alexander Nazarov, Sergey Poborchi, Iosif Polterovich, and Bartłomiej Siudeja. The legacy of Vladimir Andreevich Steklov. Notices of the American Mathematical Society, 61(1):9–22, January 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khavinson:2014:TEP

[KL14] Dmitry Khavinson and Erik Lundberg. A tale of ellipsoids in potential theory. Notices of the American Mathematical Society, 61(2):148–156, February 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kontorovich:2019:WLT

[KLLR19] Alex Kontorovich, D. Darren Long, Alexander Lubotzky, and Alan W. Reid. What is … a thin group? Notices of the American Mathematical Society, 66(6): 905–910, June/July 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kazdan:2014:PRG

[KLN^{+}14] Jerry L. Kazdan, Peter D. Lax, Albert B. J. Novikoff, C. Denson Hill, Antony Jameson, Eva V. Swenson, Ruth Bers Shapiro, Cathy Garabedian, and Emily Garabedian. Paul Roesel Garabedian (1927–2010). Notices of the American Mathematical Society, 61(3):244–255, March 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Koblitz:2010:BNW

[KM10] Neal Koblitz and Alfred Menezes. The brave new world of bodacious assumptions in cryptography. Notices of the American Mathematical Society, 57 (3):357–365, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/notices/201003/.

Khavinson:2008:FTA

[KN08] Dmitry Khavinson and Genevra Neumann. From the fundamental theorem of algebra to astrophysics: A “Harmonious” path. Notices of the American Mathematical Society, 55(6): 666–675, June/July 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic).  URL http://www.
ams.org/journals/notices/
200806/noti-jun08-cov1.jpg.

See follow-on article [Pet10].

Knapp:2014:LMH

[Kna14] Anthony W. Knapp. Love and math: the heart of hidden reality [book review of MR3155773]. Notices of the American Mathematical Society, 61(9):1056– 1060, October 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knobloch:2015:LEE

[Kno15] Eberhard Knobloch. On Leibniz, expanded edition [book review of MR3330133]. Notices of the American Mathematical Society, 62(7):820–821, August 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kobayashi:2014:TEM

[Kob14] Shoshichi Kobayashi. Two essays from Mathematicians who lost their faces. Notices of the American Mathematical Society, 61 (11):1343–1344, December 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Translated by Hisashi Kobayashi.

Kobayashi:2016:MSR

[Kob16]  Mei Kobayashi.  My search

for Ramanujan [book review of MR3495254]. Notices of the American Mathematical Society, 63(8):890–892, September 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Koberda:2018:WLA

[Kob18] Thomas Koberda. What is … an acylindrical group action? Notices of the American Mathematical Society, 65(1):31– 34, January 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kollar:2015:TCF

[Kol15] János Kollár. Is there a curse of the Fields Medal? Notices of the American Mathematical Society, 62(1):21–25, January 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kossak:2011:WLR

[Kos11] Roman Kossak. What is … a resplendent stucture? Notices of the American Mathematical Society, 58(6):812–814, June/July 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kovacs:2018:MTS

[Kov18] Sándor Kovács. Moduli theory and singularities. Notices of the American Mathematical Society, 65(4):406–408, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kritchman:2010:SEP

[KR10] Shira Kritchman and Ran Raz.

The surprise examination paradox and the second incompleteness theorem. Notices of the American Mathematical Society, 57(11):1454–1458, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kelly:2012:WEM

[KR12] Susan E. Kelly and Sarah A.

Rozner. Winifred Edgerton Merrill: “She opened the door”. Notices of the American Mathematical Society, 59(4):504–521, April 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Katz:2012:LLC

[KS12a] Mikhail G. Katz and David M.

Sherry. Leibniz’s laws of continuity and homogeneity. Notices of the American Mathematical Society, 59(11):1550–1558, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kra:2012:CM

[KS12b] Irwin Kra and Santiago R.

Simanca. On circulant matrices.

Notices of the American Mathematical Society, 59(3):368–377, March 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Krizek:2012:TCL

[KŠŠ12] Michal Křížek, Jakub Šolc, and Alena Šolcová. Is there a crystal lattice possessing five-fold symmetry? Notices of the American Mathematical Society, 59(1):22– 30, January 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kelly:2017:ELH

[KSZ17] Susan E. Kelly, Carly Shinners, and Katherine Zoroufy. Euphemia Lofton Haynes: bringing education closer to the “goal of perfection”. Notices of the American Mathematical Society, 64(9):995–1003, October 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kuznetsov:2015:PCT

[Kuz15] Nikolay Kuznetsov. 333 pages that changed theory of water waves. Notices of the American Mathematical Society, 62 (10):1208–1209, November 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kramer:2016:SMM

[KvP16] Jürg Kramer and Anna-Maria von Pippich. Snapshots of modern mathematics from Oberwolfach: special values of zeta functions and areas of triangles. Notices of the American Mathematical Society, 63(8):917–922, September 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khoshnevisan:2017:CMF

[KW17] Davar Khoshnevisan and Edward Waymire. A conversation with Mu-Fa Chen. Notices of the American Mathematical Society, 64(6):616–619, June/July 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Klartag:2018:SOP

[KW18] Bo’az Klartag and Elisabeth Werner. Some open problems in asymptotic geometric analysis. Notices of the American Mathematical Society, 65(6):657–659, June/July 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lanius:2017:RLI

[Lan17]  Melinda Lanius.  Reinhard

Laubenbacher interview. Notices of the American Mathematical Society, 64(5):456–460, May 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Landsberg:2018:GMM

[Lan18a]  J. M. Landsberg.  On the

geometry of matrix multiplication. Notices of the American Mathematical Society, 65(4):402, April 2018. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Lanius:2018:KHI

[Lan18b] Melinda Lanius. Kathryn Hess interview. Notices of the American Mathematical Society, 65 (3):317–319, March 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lanius:2018:SAI

[Lan18c] Melinda Lanius. Sharon Arroyo interview. Notices of the American Mathematical Society, 65(1):29–30, January 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Larson:2010:SMP

[Lar10a] Loren Larson. Solving mathematical problems: a personal perspective [book review of MR2265113]. Notices of the American Mathematical Society, 57(2):244–247, February 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Larusson:2010:WLO

[Lár10b] Finnur Lárusson. What is … an Oka manifold? Notices of the American Mathematical Society, 57(1):50–52, January 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Larvor:2018:WAP

[Lar18] Brendan Larvor. A welcome addition to the philosophy of mathematical practice [book review of Making and breaking mathematical sense: histories and philosophies of mathematical practice [MR3585691]]. Notices of the American Mathematical Society, 65(6):692–695, June/July 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lawrence:2019:PNB

[Law19] Emille Davie Lawrence. Power in numbers [book review, Race Point Publishing, New York, 2018]. Notices of the American Mathematical Society, 66(2): 251–253, February 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

LeBrun:2017:MOR

[LBG^{+}17] Claude LeBrun, Gérard Besson, Misha Gromov, Jim Simons, Jeff Cheeger, Jean-Pierre Bourguignon, Dennis Sullivan, Jacques Lafontaine, Jerry Kazdan, Marie-Louise Michelsohn, Pierre Pansu, David Ebin, and Karsten Grove. Marcel Berger remembered. Notices of the American Mathematical Society, 64(11):1285–1295, December 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Edited and translated by LeBrun.

Li:2018:LHS

[Li18] Y. Charles Li. Linear hydrodynamic stability. Notices of the American Mathematical Society, 65(10):1255–1259, November 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lin:2010:EWM

[Lin10] Bor-Luh Lin. Every waking moment Ky Fan (1914–2010). Notices of the American Mathematical Society, 57(11):1444–1447, December 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Linton:2012:CSM

[Lin12] Christopher Linton. Chasing shadows: mathematics, astronomy, and the early history of eclipse reckoning [book review of MR2919869]. Notices of the American Mathematical Society, 59(3):430–433, March 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lipton:2011:ATG

[Lip11] Richard J. Lipton. The annotated Turing: a guided tour through Alan Turing’s historic paper on computability and the Turing machine [book review, Wiley, Indianapolis, IN, 2008]. Notices of the American Mathematical Society, 58(8):1120– 1121, September 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lohr:2019:GMC

[Loh19] Sharon L. Lohr. Gertrude M.

Cox and statistical design. Notices of the American Mathematical Society, 66(3):317–323, March 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Levine:2010:WLS

[LP10] Lionel Levine and James Propp.

What is … a sandpile? Notices of the American Mathematical Society, 57(8):976–979, September 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

LeGall:2014:MY

[LP14] Jean-François Le Gall and Jim Pitman. Marc Yor (1949–2014). Notices of the American Mathematical Society, 61(5):508–509, May 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lenstra:2016:PTS

[LS16] Jan Karel Lenstra and David Shmoys. In pursuit of the traveling salesman: mathematics at the limits of computation [book review of MR2866515]. Notices of the American Mathematical Society, 63(6):635–638, June/July 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lindenstrauss:2019:RWU

[LSW19] Elon Lindenstrauss, Peter Sarnak, and Amie Wilkinson. Ratner’s work on unipotent flows and its impact. Notices of the American Mathematical Society, 66(3):373–377, March 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Labate:2013:W

[LWW13] Demetrio Labate, Guido Weiss, and Edward Wilson. Wavelets. Notices of the American Mathematical Society, 60(1):66–76, January 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lynch:2013:ISR

[Lyn13] Peter Lynch. Invisible in the storm: the role of mathematics in understanding weather [book review of MR3024839]. Notices of the American Mathematical Society, 60(8):1051–1053, September 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lagarias:2012:MPR

[LZ12] Jeffrey C. Lagarias and Chuanming Zong. Mysteries in packing regular tetrahedra. Notices of the American Mathematical Society, 59(11):1540–1549, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201211/
rtx121101540p.pdf.

Maggioni:2012:DM

[Mag12] Mauro Maggioni. Data mining. Notices of the American Mathematical Society, 59(4):532–533, April 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Magid:2013:WNB

[Mag13] Andy R. Magid. A wealth of numbers [book review of MR2906064]. Notices of the American Mathematical Society, 60(3):326–328, March 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Magid:2014:UHP

[Mag14] Andy Magid. Undiluted hocuspocus [book review, Princeton Univ. Press, Princeton, NJ, 2013]. Notices of the American Mathematical Society, 61 (3):281–284, March 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Maggioni:2015:GDB

[Mag15] Mauro Maggioni. Geometry of data and biology. Notices of the American Mathematical Society, 62(10):1185–1188, November 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Manin:2010:WTP

[Man10] Yuri Manin. What then? Plato’s ghost: the modernist transformation of mathematics [book review of MR2452344]. Notices of the American Mathematical Society, 57(2):239–243, February 2010. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Manaster:2015:WMG

[Man15] Alfred Manaster. What’s math got to do with it? [book review, Penguin, New York, 2015]. Notices of the American Mathematical Society, 62(9):1077–1079, October 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Martin:2010:EGB

[Mar10] Jeremy L. Martin. Euler’s gem [book review of MR2440945]. Notices of the American Mathematical Society, 57(11):1448– 1450, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Marateck:2012:YMB

[Mar12] Samuel L. Marateck. Yang– Mills and beyond. Notices of the American Mathematical Society, 59(6):802–809, June/July 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Martin:2016:SPT

[Mar16a] Gaven J. Martin. Siegel’s problem in three dimensions. Notices of the American Mathematical Society, 63(11):1244–1247, December 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Martin:2016:HBP

[Mar16b] Jeremy L. Martin. How to bake π [book review of MR3363668].

Notices of the American Mathematical Society, 63(9):1053– 1054, October 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Margalit:2019:MJB

[Mar19] Dan Margalit. The mathematics of Joan Birman. Notices of the American Mathematical Society, 66(3):341–353, March 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Masser:2019:AB

[Mas19] David Masser. Alan Baker 1939– 2018. Notices of the American Mathematical Society, 66(1):32– 35, January 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Matschke:2014:SSP

[Mat14] Benjamin Matschke. A survey on the square peg problem. Notices of the American Mathematical Society, 61(4):346–352, April 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

May:2019:MTG

[May19a] J. Peter May. Memorial tribute:

L. Gaunce Lewis Jr. (1949–2006) and Mark Steinberger (1950– 2018). Notices of the American Mathematical Society, 66(6): 864–869, June/July 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mayo:2019:PYF

[May19b] Talea L. Mayo. Predicting the 100-year flood to improve hurricane storm surge resilience. Notices of the American Mathematical Society, 66(2):176–181, February 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McCleary:2017:OHT

[McC17] John McCleary. One hundred twenty-one days [book review of 3497071]. Notices of the American Mathematical Society, 64 (7):758–760, August 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McCleary:2018:MAB

[McC18] John McCleary. Mathematics without apologies [book review of MR3289987]. Notices of the American Mathematical Society, 65(10):1280–1283, November 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McDuff:2017:ML

[McD17] Dusa McDuff. My life. Notices of the American Mathematical Society, 64(8):892–896, September 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Menne:2017:CV

[Men17] Ulrich Menne. The concept of varifold. Notices of the American Mathematical Society, 64 (10):1148–1152, November 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2019:CTC

[Mer19a] Sophia D. Merow. Has Carmichael’s totient conjecture been proven? No, no, it has not. Notices of the American Mathematical Society, 66(5):759–761, May 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2019:OSA

[Mer19b] Sophia D. Merow. One spark is all you need: Germain gets the Hamilton treatment. Notices of the American Mathematical Society, 66(8):1309–1311, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Martinez-Finkelshtein:2016:WIM

[MFV16] Andrei Martínez-Finkelshtein and Walter Van Assche. What is … a multiple orthogonal polynomial? Notices of the American Mathematical Society, 63(9): 1029–1031, October 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milnor:2011:DTF

[Mil11] John Milnor. Differential topology forty-six years later. Notices of the American Mathematical Society, 58(6):804–809, June/July 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milne:2012:WLS

[Mil12] James S. Milne. What is … a Shimura variety? Notices of the American Mathematical Society, 59(11):1560–1561, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miller:2015:FFM

[Mil15] Ezra Miller. Fruit flies and moduli: interactions between biology and mathematics. Notices of the American Mathematical Society, 62(10):1178–1184, November 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miller:2016:PGM

[Mil16] Stephen D. Miller. “I plan to be a great mathematician”: an NFL offensive lineman shows he’s one of us. Notices of the American Mathematical Society, 63(2): 148–151, February 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milicevic:2018:GPC

[Mil18] Elizabeth T. Milićević. Genius at play: the curious mind of John Horton Conway [book review of MR3329687]. Notices of the American Mathematical Society, 65(10):1277–1279, November 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miller:2019:MPB

[Mil19] Steven J. Miller. Maker of patterns [book review, Liveright, New York, 2018]. Notices of the American Mathematical Society, 66(5):750–753, May 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Marcellan:2017:WLS

[MMB17] Francisco Marcellán and Juan J.

Moreno-Balcázar. What is … a Sobolev orthogonal polynomial? Notices of the American Mathematical Society, 64(8):873–875, September 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moll:2010:SO

[Mol10] Victor H. Moll. Seized opportunities. Notices of the American Mathematical Society, 57 (4):476–484, April 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Montelle:2013:CPN

[Mon13] Clemency Montelle. The crest of the peacock: non-European roots of mathematics [book review of MR2722967]. Notices of the American Mathematical Society, 60(11):1459–1463, December 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Montalban:2019:MCC

[Mon19] Antonio Montalbán. Martin’s conjecture: a classification of the naturally occurring Turing degrees. Notices of the American Mathematical Society, 66 (8):1209–1215, September 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moosa:2010:MTC

[Moo10] Rahim Moosa. Model theory and complex geometry. Notices of the American Mathematical Society, 57(2):230–235, February 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moricz:2014:MIM

[Mór14] Ferenc Móricz. Moment inequalities for maxima of partial sums in probability with applications in the theory of orthogonal series. Notices of the American Mathematical Society, 61(6): 576–585, June/July 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Morse:2018:CCS

[Mor18] Jennifer Morse. Computing, combinatorics, and k-Schur functions. Notices of the American Mathematical Society, 65 (4):403–404, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McAdams:2010:CWA

[MOT10] Aleka McAdams, Stanley Osher, and Joseph Teran. Crashing waves, awesome explosions, turbulent smoke, and beyond: applied mathematics and scientific computing in the visual effects industry. Notices of the American Mathematical Society, 57 (5):614–623, May 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Muller-Stach:2014:WLP

[MS14] Stefan Müller-Stach. What is … a period? Notices of the American Mathematical Society, 61(8):898–899, September 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ma:2019:RAA

[MS19] Linquan Ma and Karl Schwede.

Recent applications of p-adic methods to commutative algebra. Notices of the American Mathematical Society, 66(6): 820–831, June/July 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mumford:2010:MIB

[Mum10] David Mumford. Mathematics in India [book review of MR2468443]. Notices of the American Mathematical Society, 57(3):385–390, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mumford:2011:IRE

[Mum11] David Mumford. Intuition and rigor and Enriques’s quest. Notices of the American Mathematical Society, 58(2):250–260, February 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mummert:2018:RMB

[Mum18] Carl Mummert. Reverse mathematics [book review of MR3729321]. Notices of the American Mathematical Society, 65(9):1098–1102, October 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Murray:2012:WTN

[Mur12] J. D. Murray. Why are there no 3-headed monsters? Mathematical modeling in biology. Notices of the American Mathematical Society, 59(6):785–795, June/July 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mordukhovich:2012:JW

[MZ12] Boris Mordukhovich and Qiji Zhu. Jack Warga (1922–2011). Notices of the American Mathematical Society, 59(8):1109, September 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Naimpally:2013:JGH

[Nai13] Som Naimpally. John G. Hocking. Notices of the American Mathematical Society, 60(3):

335, March 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nathanson:2014:AMM

[Nat14] Melvyn B. Nathanson. An American mathematician in Moscow, or how I destroyed the Soviet Union. Notices of the American Mathematical Society, 61(2): 186–189, February 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nkwanta:2018:ELG

[NB18] Asamoah Nkwanta and Janet E.

Barber. Episodes in the life of a genius: J. Ernest Wilkins Jr. Notices of the American Mathematical Society, 65(2):135–138, February 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Needell:2019:LDA

[Nee19] Deanna Needell. Large data analysis and Lyme disease. Notices of the American Mathematical Society, 66(1):8–15, January 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nelson:2011:CRK

[Nel11] Sam Nelson. The combinatorial revolution in knot theory. Notices of the American Mathematical Society, 58(11):1553–1561, December 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nelson:2016:WLQ

[Nel16] Sam Nelson. What is … a quandle? Notices of the American Mathematical Society, 63 (4):378–380, April 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Neumann:2011:MWE

[Neu11] Peter M. Neumann. The Mathematical Writings of Évariste Galois. European Mathematical Society, Zürich, Switzerland, 2011. ISBN 3-03719-104-X. x + 410 pp. LCCN QA3 .G18 2011.

Nicholson:2012:PWU

[Nic12] Jason Scott Nicholson. A perspective on Wigner’s “Unreasonable effectiveness of mathematics”. Notices of the American Mathematical Society, 59(1):38–42, January 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/201201/
                                        rtx120100038p.pdf.

Nielsen:2018:WLI

[Nie18] Frank Nielsen. What is … an information projection? Notices of the American Mathematical Society, 65(3):321–324, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Napoletani:2014:BDE

[NPS14] Domenico Napoletani, Marco Panza, and Daniele C. Struppa.

Is big data enough? A reflection on the changing role of mathematics in applications. Notices of the American Mathematical Society, 61(5):485–490, May 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nazarov:2010:WGE

[NS10] Fedor Nazarov and Mikhail Sodin. What is … a Gaussian entire function? Notices of the American Mathematical Society, 57(3):375–377, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Novak:2011:WLF

[NŚ11] Jonathan Novak and Piotr Śniady. What is … a free cumulant? Notices of the American Mathematical Society, 58(2): 300–301, February 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Osborne:2013:GSE

[OH13] Jason M. Osborne and Gregory P. Hicks. The geodesic spring on the Euclidean sphere with parallel-transport-based damping. Notices of the American Mathematical Society, 60 (5):544–556, May 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Okoudjou:2019:IGA

[Oko19] Kasso A. Okoudjou. An invitation to Gabor analysis. Notices of the American Mathematical Society, 66(6):808–819, June/July 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Olver:2011:JF

[Olv11] Peter J. Olver. Journals in flux. Notices of the American Mathematical Society, 58 (8):1124–1126, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ono:2010:LWG

[Ono10] Ken Ono. The last words of a genius. Notices of the American Mathematical Society, 57 (11):1410–1419, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ooguri:2015:IEW

[Oog15] Hirosi Ooguri. Interview with Edward Witten. Notices of the American Mathematical Society, 62(5):491–506, May 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Oort:2013:DMS

[Oor13] Frans Oort. David Mumford — selected papers, volume II [book review, Springer, Berlin, 2010]. Notices of the American Mathematical Society, 60(2):214–217, February 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ornstein:2013:NLC

[Orn13] Don Ornstein. Newton’s laws and coin tossing. Notices of the American Mathematical Society, 60(4):450–459, April 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

OShea:2011:BRP

[O’S11] Donal O’Shea. Book review:

Prefect rigor: a genius and the mathematical breakthrough of the century [MR2598223]. Notices of the American Mathematical Society, 58(1):56–59, January 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

OShea:2018:PNR

[O’S18] Donal O’Shea. Prime numbers and the Riemann hypothesis [book review of MR3616260]. Notices of the American Mathematical Society, 65(7):811–815, August 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Osserman:2010:MGA

[Oss10] Robert Osserman. Mathematics of the gateway arch. Notices of the American Mathematical Society, 57(2):220–229, February 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Owhadi:2019:SNA

[OSS19] Houman Owhadi, Clint Scovel, and Florian Schäfer. Statistical numerical approximation. Notices of the American Mathematical Society, 66(10):1608–1617, November 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pak:2019:CI

[Pak19] Igor Pak. Combinatorial inequalities. Notices of the American Mathematical Society, 66(7): 1109–1112, August 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pansu:2018:JFH

[Pan18] Pierre Pansu. Jacqueline Ferrand and her oeuvre. Notices of the American Mathematical Society, 65(2):201–205, February 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Papadopoulos:2017:MAS

[Pap17] Jim Papadopoulos. Mathematics applied to the study of bicycles. Notices of the American Mathematical Society, 64(8):888–890, September 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Papadopoulos:2018:HB

[Pap18] Athanase Papadopoulos. Herbert Busemann. Notices of the American Mathematical Society, 65(3):341–343, March 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Papadopoulos:2019:MLD

[Pap19] Athanase Papadopoulos. Maps with least distortion between surfaces: from geography to brain warping. Notices of the American Mathematical Society, 66(10):1628–1639, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Peifer:2017:DRM

[Pei17] David Peifer. Dorothea Rockburne and Max Dehn at Black Mountain College. Notices of the American Mathematical Society, 64(11):1313–1318, December 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pengelley:2011:MES

[Pen11] David Pengelley. Mathematics emerging: a sourcebook 1540–1900 [book review of MR2473212]. Notices of the American Mathematical Society, 58(6):815–819, June/July 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Perlick:2000:ROF

[Per00] Volker Perlick. Ray Optics, Fermat’s Principle, and Applications to General Relativity, volume 61 of Lecture notes in physics: New series m, Monographs; Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66898-5. x + 220 pp. LCCN QC389 .P37 2000.

Perez:2011:BHA

[Pér11] Rodrigo A. Pérez. A brief but historic article of Siegel. Notices of the American Mathematical Society, 58(4):558–566, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Perez:2017:NGA

[Pér17] Joaquín Pérez. A new golden age of minimal surfaces. Notices of the American Mathematical Society, 64(4):347–358, April 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Perea:2019:TTS

[Per19a] Jose A. Perea. Topological times series analysis. Notices of the American Mathematical Society, 66(5):686–694, May 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Perutz:2019:MMS

[Per19b] Timothy Perutz. The mirror’s magic sights: an update on mirror symmetry. Notices of the American Mathematical Society, 66(4):483–492, April 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Petters:2010:GAL

[Pet10] Arlie O. Petters. Gravity’s action on light. Notices of the American Mathematical Society, 57 (11):1392–1409, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic).  URL http://www.
ams.org/journals/notices/
201011/201011-full-issue.

pdf. See preceding article [KN08], and book-length treatment [Per00].

Petters:2018:B

[Pet18] Arlie O. Petters. Belonging.

Notices of the American Mathematical Society, 65(2):120–123, February 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Petrovic:2019:WLM

[Pet19] Sonja Petrović. What is … a Markov basis? Notices of the American Mathematical Society, 66(7):1088–1092, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Philippon:2012:FT

[Phi12a] Patrice Philippon. A Farey tail.

Notices of the American Mathematical Society, 59(6):746–757, June/July 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Phillips:2012:GMR

[Phi12b]  Anthony Phillips.  Galileo’s

muse: Renaissance mathematics and the arts [book review, Harvard Univ. Press, Cambridge, MA, 2011]. Notices of the American Mathematical Society, 59 (10):1435–1442, November 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Panaretos:2012:IMB

[PM12] John Panaretos and Chrisovalandis Malesios. Influential mathematicians: birth, education, and affiliation. Notices of the American Mathematical Society, 59(2): 274–286, February 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Peters:2012:ANS

[PN12] Jim Peters and Som Naimpally. Applications of near sets. Notices of the American Mathematical Society, 59(4):536–542, April 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Poovey:2017:WMD

[Poo17]  Mary Poovey.  Weapons of

math destruction [book review of MR3561130]. Notices of the American Mathematical Society, 64(8):933–935, September 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Porter:2010:CBU

[Por10] Mason A. Porter. Can baseball be used to teach statistics? [book review, Math. Assoc. America, Washington D.C., 2003]. Notices of the American Mathematical Society, 57(4):503–507, April 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Porter:2018:WLM

[Por18] Mason A. Porter. What is … a multilayer network? Notices of the American Mathematical Society, 65(11):1419–1423, December 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Petit:2013:RC

[PQ13] Christophe Petit and Jean-Jacques Quisquater. Rubik’s for cryptographers. Notices of the American Mathematical Society, 60(6):733–740, June/July 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pitman:2012:AGS

[PR12] Jim Pitman and Nathan Ross.

Archimedes, Gauss, and Stein. Notices of the American Mathematical Society, 59(10):1416– 1421, November 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Quarteroni:2018:TMH

[Qua18] Alfio Quarteroni. Taking mathematics to heart. Notices of the American Mathematical Society, 65(1):19–24, January 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Quertermous:2019:WMB

[Que19] Katie Spurrier Quertermous. Women in mathematics [book review of MR3775369]. Notices of the American Mathematical Society, 66(3):395–398, March 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Quinn:2012:RMW

[Qui12] Frank Quinn. A revolution in mathematics? What really happened a century ago and why it matters today. Notices of the American Mathematical Society, 59(1):31–37, January 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Radin:2013:RM

[Rad13] Charles Radin. A revolutionary material. Notices of the American Mathematical Society, 60(3):310–315, March 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Radin:2017:OMC

[Rad17] Charles Radin. The open mathematics of crystallization. Notices of the American Mathematical Society, 64(6):551–556, June/July 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Range:2012:WLP

[Ran12] R. Michael Range. What is … a pseudoconvex domain? Notices of the American Mathematical Society, 59(2):301–303, February 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Range:2014:DDP

[Ran14] R. Michael Range. Descartes’s double point method for tangents: an old idea suggests new approaches to calculus. Notices of the American Mathematical Society, 61(4):387–389, April 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rao:2014:PSM

[Rao14] Arni S. R. Srinivasa Rao. Population stability and momentum. Notices of the American Mathematical Society, 61(9): 1062–1065, October 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ravenel:2016:RMM

[Rav16] Douglas C. Ravenel. Remembering Mark Mahowald 1931–2013. Notices of the American Mathematical Society, 63(6):652–656, June/July 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Razborov:2013:WLF

[Raz13] Alexander A. Razborov. What is … a flag algebra? Notices of the American Mathematical Society, 60(10):1324–1327, November 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rehmeyer:2010:MDI

[Reh10] Julie Rehmeyer. The mathematical dramatist: interview with Gioia De Cari. Notices of the American Mathematical Society, 57(6):740–743, June/July 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reichstein:2012:WLE

[Rei12] Zinovy Reichstein. What is … essential dimension? Notices of the American Mathematical Society, 59(10):1432–1434, November 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ribenboim:2019:EGK

[Rib19]  Paulo Ribenboim.  Excerpt

from The Grothendieck I knew: telling, not hiding, not judging. Notices of the American Mathematical Society, 66(7): 1069–1077, August 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Richert:2014:MRC

[Ric14] Norman Richert. Mathematical Reviews celebrates 75 years. Notices of the American Mathematical Society, 61(11):1355–1356, December 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Riviere:2016:EUW

[Riv16] Tristan Rivière. Exploring the unknown: the work of Louis Nirenberg on partial differential equations. Notices of the American Mathematical Society, 63(2): 120–125, February 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Roberts:2010:AEP

[Rob10] Catherine A. Roberts. Awareness of ethical pitfalls: a requirement for professional protection. Notices of the American Mathematical Society, 57(4):485–489, April 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Robinson:2015:MRH

[Rob15] E. Arthur Robinson, Jr. Man Ray’s human equations. Notices of the American Mathematical Society, 62(10):1192–1198, November 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Robinson:2019:HFS

[Rob19] Michael Robinson. Hunting for foxes with sheaves. Notices of the American Mathematical Society, 66(5):661–676, May 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rockmore:2011:CGT

[Roc11] Dan Rockmore. Complexity: a guided tour [book review of MR2641048]. Notices of the American Mathematical Society, 58(4):585–587, April 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rockmore:2012:OMR

[Roc12] Dan Rockmore. One in a million — remembrances of Laurie Snell (1925–2011). Notices of the American Mathematical Society, 59(9):1256–1260, October 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Roitman:2010:LES

[Roi10] Judith Roitman. Logicomix: an epic search for truth [book review of MR2884886]. Notices of the American Mathematical Society, 57(11):1427–1431, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Roitman:2017:MGR

[Roi17] Judith Roitman. Math graphics: a review of two books [book review of MR3558830 and MR3561047]. Notices of the American Mathematical Society, 64(5):504–507, May 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rosendal:2011:DCT

[Ros11] Christian Rosendal. Descriptive classification theory and separable Banach spaces. Notices of the American Mathematical Society, 58(9):1251–1262, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ross:2014:TCA

[Ros14] Kenneth A. Ross. A trip from classical to abstract Fourier analysis. Notices of the American Mathematical Society, 61(9): 1032–1038, October 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ross:2019:CCA

[Ros19]  William T. Ross.  A com-

prehensive course in analysis [book review of MR3408971; MR3443339; MR3364090; MR3410783; MR3364494]. Notices of the American Mathematical Society, 66(1):74–76, January 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rothman:2017:MEG

[Rot17] Daniel H. Rothman. Mathematical expression of a global environmental catastrophe. Notices of the American Mathematical Society, 64(2):138–140, February 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rousseau:2019:IDM

[Rou19] Christiane Rousseau. International Day of Mathematics. Notices of the American Mathematical Society, 66(8):1306–1307, September 2019. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rowlett:2014:LST

[Row14] Julie Rowlett. The level sets of typical games. Notices of the American Mathematical Society, 61(8):840–847, September 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rowland:2015:WLA

[Row15] Eric Rowland. What is … an automatic sequence? Notices of the American Mathematical Society, 62(3):274–276, March 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rowe:2019:EIM

[Row19] David E. Rowe. Einstein’s Italian mathematicians: Ricci, Levi-Civita, and the birth of general relativity [book review of MR3821554]. Notices of the American Mathematical Society, 66(9):1477–1483, October 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Richert:2013:ACA

[RR13] Benjamin P. Richert and Norman Richert. An avuncular chat about reviewing for Mathematical Reviews. Notices of the American Mathematical Society, 60(6):775–776, June/July 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2010:IMG

[RS10] Martin Raussen and Christian Skau. Interview with Mikhail Gromov. Notices of the American Mathematical Society, 57 (3):391–403, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2011:IAL

[RS11] Martin Raussen and Christian Skau. Interview with Abel Laureate John Tate. Notices of the American Mathematical Society, 58(3):444–452, March 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2012:IJM

[RS12] Martin Raussen and Christian Skau. Interview with John Milnor. Notices of the American Mathematical Society, 59(3): 400–408, March 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2013:IES

[RS13] Martin Raussen and Christian Skau. Interview with Endre Szemerédi [reprint of MR2986979]. Notices of the American Mathematical Society, 60(2):221–231, February 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2014:IPD

[RS14] Martin Raussen and Christian Skau. Interview with Pierre Deligne [reprint of MR3114790]. Notices of the American Mathematical Society, 61(2):177–185, February 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2016:IAL

[RS16a] Martin Raussen and Christian Skau. Interview with Abel laureate John F. Nash Jr. [reprint of MR3409221]. Notices of the American Mathematical Society, 63(5):486–491, May 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2016:ILN

[RS16b] Martin Raussen and Christian Skau. Interview with Louis Nirenberg [reprint of MR3409067]. Notices of the American Mathematical Society, 63(2):135–140, February 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2017:IAL

[RS17] Martin Raussen and Christian Skau. Interview with Abel Laureate Sir Andrew J. Wiles. Notices of the American Mathematical Society, 64(3):198–208, March 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprint of [MR3526311].

Reiner:2014:WLC

[RSW14] Victor Reiner, Dennis Stanton, and Dennis White. What is … cyclic sieving? Notices of the American Mathematical Society, 61(2):169–171, February 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Readdy:2018:EWM

[RT18a] Margaret A. Readdy and Christine Taylor. Early women mathematicians in Princeton. Notices of the American Mathematical Society, 65(3):304–305, March 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Readdy:2018:WHM

[RT18b] Margaret A. Readdy and Christine Taylor. Women’s history month. Notices of the American Mathematical Society, 65 (3):248–249, March 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rao:2012:ISP

[RTKB12] Arni S. R. Srinivasa Rao, Kurien Thomas, Sudhakar Kurapati, and Ramesh Bhat. Improvement in survival of people living with HIV/AIDs and requirement for 1st- and 2nd-line ART in India: a mathematical model.

Notices of the American Mathematical Society, 59(4):560–562, April 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Serre:2010:THC

[SA10] Jean-Pierre Serre and Michael Atiyah. A tribute to Henri Cartan. Notices of the American Mathematical Society, 57(8): 946–949, September 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saaty:2013:MIP

[Saa13] Thomas L. Saaty. On the measurement of intangibles. A principal eigenvector approach to relative measurement derived from paired comparisons. Notices of the American Mathematical Society, 60(2):192–208, February 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saclolo:2011:HMT

[Sac11] Michael P. Saclolo. How a medieval troubadour became a mathematical figure. Notices of the American Mathematical Society, 58(5):682–687, May 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sangwin:2011:MJE

[San11] Chris Sangwin. Modelling the journey from elementary word problems to mathematical research. Notices of the American Mathematical Society, 58 (10):1436–1445, November 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sarnak:2010:RPC

[Sar10] Peter Sarnak. Remembering Paul Cohen (1934–2007). Notices of the American Mathematical Society, 57(7):824–838, August 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saul:2014:RNG

[Sau14] Mark Saul. Ramanujan and the nature of genius [film review: The genius of Srinivasa Ramanujan]. Notices of the American Mathematical Society, 61 (11):1348–1350, December 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saul:2016:WDW

[Sau16] Mark Saul. Why do we do mathematics? Thoughts on the film A brilliant young mind [movie review, Samuel Goldwyn Films, 2014]. Notices of the American Mathematical Society, 63(9): 1050–1051, October 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saxe:2018:CMH

[Sax18] Karen Saxe. Can math help the Supreme Court free Congress from Gerrymanders? Notices of the American Mathematical Society, 65(1):37–39, January 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schattschneider:2010:MSM

[Sch10]  Doris Schattschneider.  The

mathematical side of M. C. Escher. Notices of the American Mathematical Society, 57(6): 706–718, June/July 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schneiderman:2011:COH

[Sch11] Rob Schneiderman. Can one hear the sound of a theorem? Notices of the American Mathematical Society, 58(7):929–937, August 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schleicher:2013:IJH

[Sch13a] Dierk Schleicher. Interview with John Horton Conway. Notices of the American Mathematical Society, 60(5):567–575, May 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schmidt:2013:MLB

[Sch13b] William Schmidt. A mathematician’s lament [book review, Bellevue Library Press, New York, 2009]. Notices of the American Mathematical Society, 60(4):461–462, April 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scherich:2018:DYP

[Sch18a] Nancy Scherich. Dance your PhD: representations of the braid groups. Notices of the American Mathematical Society, 65(4):420–421, April 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schochet:2018:CSN

[Sch18b] Claude L. Schochet. A case study in noncommutative topology. Notices of the American Mathematical Society, 65(6): 653–656, June/July 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Skoda:2014:MPL

[SDS14] Henri Skoda, Jean-Pierre Demailly, and Yum-Tong Siu. In memory of Pierre Lelong. Notices of the American Mathematical Society, 61(6):586–595, June/July 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Henri Skoda, coordinating editor.

Senechal:2013:TTJ

[Sen13] Marjorie Senechal. Transcending tradition: Jewish mathematicians in German-speaking academic culture [book review of MR3087339]. Notices of the American Mathematical Society, 60(2):209–213, February 2013.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Serafino:2014:ODW

[Ser14] Loris Serafino. Optimizing without derivatives: what does the No Free Lunch Theorem actually say? Notices of the American Mathematical Society, 61 (7):750–755, August 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sethuraman:2010:DAW

[Set10] B. A. Sethuraman. Division algebras and wireless communication. Notices of the American Mathematical Society, 57 (11):1432–1439, December 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sharifi:2019:ITC

[Sha19] Romyar Sharifi. Iwasawa theory: a climb up the tower. Notices of the American Mathematical Society, 66(1):16–26, January 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shlyakhtenko:2018:CHT

[Shl18]  Dimitri Shlyakhtenko.  A

(co)homology theory for subfactors and planar algebras. Notices of the American Mathematical Society, 65(4):408–409, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shparlinski:2010:NWP

[Shp10] Igor E. Shparlinski. Numbers at work and play. Notices of the American Mathematical Society, 57(3):334–342, March 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/notices/201003/.

Shu:2012:EAS

[Shu12] Chi-Wang Shu. Efficient algorithms for solving partial differential equations with discontinuous solutions. Notices of the American Mathematical Society, 59(5):615–621, May 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Silver:2010:DDH

[Sil10] Daniel S. Silver. Duel at dawn:

heroes, martyrs, and the rise of modern mathematics [book review of MR2654369]. Notices of the American Mathematical Society, 57(10):1297–1299, November 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201302/
rnoti-p154.pdf.

Silverman:2013:WLA

[Sil13] Joseph H. Silverman. What is … the p-adic Mandelbrot set? Notices of the American Mathematical Society, 60 (8):1048–1050, September 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Silver:2014:KNC

[Sil14] Daniel S. Silver. Knots in the nursery: (Cats) Cradle Song of James Clerk Maxwell. Notices of the American Mathematical Society, 61(10):1186–1194, November 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Silva:2016:WLE

[Sil16] Cesar E. Silva. What is … an ergodic transformation? Notices of the American Mathematical Society, 63(1):26–27, January 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schaeffer:2012:TLP

[SLL12] Philip Schaeffer, Simon Levin, and Howard A. Levine. Tribute to Lawrence E. Payne. Notices of the American Mathematical Society, 59(5):653–654, May 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sloane:2018:LEI

[Slo18] Neil J. A. Sloane. The online encyclopedia of integer sequences. Notices of the American Mathematical Society, 65(9): 1062–1074, October 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Smith:2014:HJD

[Smi14] Alvy Ray Smith. His just deserts: a review of four books [book reviews of MR2963548, MR2919681, MR3185259, and Alan Turing’s electronic brain, Oxford University Press, 2012]. Notices of the American Mathematical Society, 61(8):891–895, September 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://www.ams.org/notices/
201408/rnoti-p891.pdf. Re-

views of: Alan Turing: The Enigma: The Centenary Edition, Andrew Hodges, Princeton University Press, May 2012 US$24.95, 632 pages, ISBN-13 978-0-691-15564-7; Alan M. Turing: Centenary Edition, Sara Turing, Cambridge University Press, April 2012, US$31.99, 193 pages, ISBN-13 978-1-107-02058-0; Alan Turing’s Electronic Brain: The Struggle to Build the ACE, the World’s Fastest Computer, B. Jack Copeland and others, Oxford University Press, May 2012, US$33.97, 592 pages, ISBN-13 978-0-19-960915-4; Turing: Pioneer of the Information Age, Jack Copeland, Oxford University Press, January 2013, US$21.95, 224 pages, ISBN-13 978-0-19-963979-3.

Snaith:2013:HAK

[Sna13] Victor P. Snaith. A history of the Arf–Kervaire invariant problem. Notices of the American Mathematical Society, 60 (8):1040–1047, September 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sobczyk:2012:CMG

[Sob12] Garret Sobczyk. Conformal mappings in geometric algebra. Notices of the American Mathematical Society, 59(2):264–273, February 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Soberon:2017:GST

[Sob17]  Pablo Soberón.  Gerryman-

dering, sandwiches, and topology. Notices of the American Mathematical Society, 64(9): 1010–1013, October 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Solomon:2018:CFS

[Sol18] Ronald Solomon. The classification of finite simple groups: a progress report. Notices of the American Mathematical Society, 65(6):646–651, June/July 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Simoes-Pereira:2018:WLM

[SP18] J. M. S. Simões-Pereira. What is … a matroidal family of graphs? Notices of the American Mathematical Society, 65(8):947–948, September 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Spencer:2010:SPN

[Spe10a] Joel Spencer. The solitude of prime numbers [book review, Pamela Dorman Books, New York, 2010]. Notices of the American Mathematical Society, 57(8):980–981, September 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Spencer:2010:GCG

[Spe10b] Joel Spencer. The giant component: the golden anniversary. Notices of the American Mathematical Society, 57(6):720–724, June/July 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Skinner:2017:MWA

[SRM^{+}17] Christopher Skinner, Karl Rubin, Barry Mazur, Mirela Çiperiani, Chandrashekhar Khare, and Jack Thorne. The mathematical works of Andrew Wiles. Notices of the American Mathematical Society, 64(3):217–227, March 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schleimer:2015:PC

[SS15a] Saul Schleimer and Henry Segerman. Puzzling the 120-cell. Notices of the American Mathematical Society, 62(11):1309– 1316, December 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Serrano:2015:MMN

[SS15b] Isabel M. Serrano and Bogdan D. Suceavă. A medieval mystery: Nicole Oresme’s concept of curvitas. Notices of the American Mathematical Society, 62(9):

1030–1034, October 2015. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Siegmund-Schultze:2015:VWT

[SS15c] Reinhard Siegmund-Schultze. Van der Waerden in the Third Reich [book review of MR3243751]. Notices of the American Mathematical Society, 62(8):924–929, September 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sidman:2017:RFT

[SS17] Jessica Sidman and Audrey St. John. The rigidity of frameworks: theory and applications. Notices of the American Mathematical Society, 64(9):973–978, October 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Steinsaltz:2011:VNR

[Ste11] David Steinsaltz. The value of nothing: a review of The Quants [book review, Crown Business, New York, 2010]. Notices of the American Mathematical Society, 58(5):699–704, May 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stewart:2014:NYT

[Ste14] Ian Stewart. The New York Times book of mathematics [book review of MR3137670]. Notices of the American Mathematical Society, 61(5):495–497, May 2014. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Stillwell:2014:HPS

[Sti14] John Stillwell. Henri Poincaré.

A scientific biography [book review of MR2986502]. Notices of the American Mathematical Society, 61(4):378–383, April 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stovall:2019:WST

[Sto19] Betsy Stovall. Waves, spheres, and tubes: a selection of Fourier restriction problems, methods, and applications. Notices of the American Mathematical Society, 66(7):1013–1022, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stroock:2015:BTB

[Str15] Daniel W. Stroock. Birth of a theorem [book review, Farrar Straus Giroux, New York, NY, 2015]. Notices of the American Mathematical Society, 62 (10):1214–1215, November 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Striker:2017:DAC

[Str17] Jessica Striker. Dynamical algebraic combinatorics: promotion, rowmotion, and resonance. Notices of the American Mathematical Society, 64(6):543–549, June/July 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Street:2018:WEA

[Str18] Brian Street. What else about … hypoellipticity? Notices of the American Mathematical Society, 65(4):424–425, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sturmfels:2016:TTE

[Stu16] Bernd Sturmfels. Tensors and their eigenvectors. Notices of the American Mathematical Society, 63(6):604–606, June/July 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sethi:2016:WLN

[SW16] Rajiv Sethi and Jörgen Weibull. What is … Nash equilibrium? Notices of the American Mathematical Society, 63(5):526–528, May 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Silver:2017:WLV

[SW17] Daniel S. Silver and Susan G. Williams. What is … a virtual knot? Notices of the American Mathematical Society, 64(5):461–464, May 2017. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stroock:2011:RPM

[SYK^{+}11] Daniel W. Stroock, Marc Yor, Jean-Pierre Kahane, Richard Gundy, Leonard Gross, and Michèle Vergne. Remembering Paul Malliavin. Notices of the American Mathematical Society, 58(4):568–573, April 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saaty:2013:PIP

[SZ13] Thomas L. Saaty and H. J. Zoffer. Principles for implementing a potential solution to the Middle East conflict. Notices of the American Mathematical Society, 60(10):1300–1322, November 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tefera:2010:WLW

[Tef10] Akalu Tefera. What is … a Wilf–Zeilberger pair? Notices of the American Mathematical Society, 57(4):508–509, April 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Thompson:2014:DDT

[Tho14] Abigail Thompson. Does diversity trump ability? An example of the misuse of mathematics in the social sciences. Notices of the American Mathematical Society, 61(9):1024–1030, October 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Toro:2019:GMT

[Tor19] Tatiana Toro. Geometric measure theory — recent applications. Notices of the American Mathematical Society, 66 (4):474–481, April 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tosatti:2018:MLC

[Tos18] Valentino Tosatti. Metric limits of Calabi–Yau manifolds. Notices of the American Mathematical Society, 65(4):413–414, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trefethen:2016:IY

[Tre16] Lloyd N. Trefethen. Inverse Yogiisms. Notices of the American Mathematical Society, 63 (11):1281–1285, December 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tu:2011:WLE

[Tu11] Loring W. Tu. What is … equivariant cohomology? Notices of the American Mathematical Society, 58(3):423–426, March 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tu:2015:GWH

[Tu15] Loring W. Tu. On the genesis of the Woods Hole fixed point theorem. Notices of the American Mathematical Society, 62 (10):1200–1206, November 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tumpach:2016:GID

[Tum16] Alice Barbara Tumpach. Gauge invariance of degenerate Riemannian metrics. Notices of the American Mathematical Society, 63(4):342–350, April 2016. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tenner:2019:ACI

[TW19] Bridget Eileen Tenner and Gregory S. Warrington. Accumulation charts for instant-runoff elections. Notices of the American Mathematical Society, 66 (11):1793–1799, December 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tran:2017:ATM

[TY17] Ngoc M. Tran and Jed Yang. Antibiotics time machines are hard to build. Notices of the American Mathematical Society, 64 (10):1136–1140, November 2017. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

VanOsdol:2012:MMB

[Van12] Donovan H. Van Osdol. Magical mathematics [book review of MR2858033]. Notices of the American Mathematical Society, 59(7):960–961, August 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vandervelde:2015:PSC

[Van15] Sam Vandervelde. Professor Stewart’s Casebook of mathematical mysteries [book review, Basic Books, New York, 2014]. Notices of the American Mathematical Society, 62(11):1361–1363, December 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

vandenBerg:2015:RND

[vdBL15] Jan Bouwe van den Berg and Jean-Philippe Lessard. Rigorous numerics in dynamics. Notices of the American Mathematical Society, 62(9):1057–1061, October 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

VerdeSaraiva:2013:GSS

[Ver13] José Cloves Verde Saraiva. Geraldo Severo de Souza Ávila — a math professor. Notices of the American Mathematical Society, 60(5):595, May 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vezzosi:2011:WLD

[Vez11]  Gabriele Vezzosi.  What is

… a derived stack? Notices of the American Mathematical Society, 58(7):955–958, August 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vidick:2019:OAC

[Vid19] Thomas Vidick. From operator algebras to complexity theory and back. Notices of the American Mathematical Society, 66 (10):1618–1627, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vinzant:2014:WLS

[Vin14] Cynthia Vinzant. What is … a spectrahedron? Notices of the American Mathematical Society, 61(5):492–494, May 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vejdemo-Johansson:2019:MHM

[VJCC19] Mikael Vejdemo-Johansson, Justin Curry, and Julie Corrigan. Mental health in the mathematics community. Notices of the American Mathematical Society, 66(7):1079–1084, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Walter:2010:EWM

[Wal10] Martin E. Walter. Earthquakes and weatherquakes: mathematics and climate change. Notices of the American Mathematical Society, 57(10):1278–1284, November 2010. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Walker:2015:CSA

[Wal15] James S. Walker. Creating symmetry: the artful mathematics of wallpaper patterns [book review of MR3362188]. Notices of the American Mathematical Society, 62(11):1350–1354, December 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Warren:2013:NEC

[War13] Richard H. Warren. Numeric experiments on the commercial quantum computer. Notices of the American Mathematical Society, 60(11):1434–1438, December 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Washington:2018:BES

[Was18] Talitha M. Washington. Behind every successful woman, there are a few good men. Notices of the American Mathematical Society, 65(2):132–134, February 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wayne:2011:VTD

[Way11] C. Eugene Wayne. Vortices and two-dimensional fluid motion. Notices of the American Mathematical Society, 58(1):10– 19, January 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weinberger:2011:WLP

[Wei11] Shmuel Weinberger. What is … persistent homology? Notices of the American Mathematical Society, 58(1):36–39, January 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weil:2018:MFA

[Wei18] Sylvie Weil. My father André Weil. Notices of the American Mathematical Society, 65(1):54– 57, January 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weibel:2019:VV

[Wei19] Charles Weibel. Vladimir Voevodsky. Notices of the American Mathematical Society, 66 (4):526–533, April 2019. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wickelgren:2016:WLA

[Wic16] Kirsten Wickelgren. What is … an Anabelian scheme? Notices of the American Mathematical Society, 63(3):285–286, March 2016. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wickelgren:2018:ACL

[Wic18] Kirsten Wickelgren. An arithmetic count of the lines on a smooth cubic surface. Notices of the American Mathematical Society, 65(4):404–405, April 2018. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wigderson:2014:QCS

[Wig14] Avi Wigderson. Quantum computing since Democritus [book review of MR3058839]. Notices of the American Mathematical Society, 61(10):1218–1220, November 2014. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Williams:2018:IHB

[Wil18a] Talithia Williams. An introduction [In honor of black history]. Notices of the American Mathematical Society, 65(2):119, February 2018. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Williams:2018:PC

[Wil18b] Talithia Williams. Personal conclusions. Notices of the American Mathematical Society, 65 (2):148, February 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wegert:2011:PPC

[WS11] Elias Wegert and Gunter Semmler. Phase plots of complex functions: a journey in illustration. Notices of the American Mathematical Society, 58(6):768–780, June/July 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wu:2017:AWW

[Wu17] Wentsun Wu. Autobiography of Wentsun Wu (1919–2017). Notices of the American Mathematical Society, 64(11):1319–1320, December 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Yu:2019:RGD

[YHH19] Guoliang Yu, Roger Howe, and Shana Hutchins. Ronald G. Douglas: a master in the art of transcending problems. Notices of the American Mathematical Society, 66(10):1672–1684, November 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Yau:2011:STG

[YN11] Shing-Tung Yau and Steve Nadis. String theory and the geometry of the universe’s hidden dimensions. Notices of the American Mathematical Society, 58(8):1067–1076, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201108/
rtx110801067p.pdf.

Yong:2014:CHC

[Yon14] Alexander Yong. Critique of Hirsch’s citation index: a combinatorial Fermi problem. Notices of the American Mathematical Society, 61(9):1040–1050, October 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zakharevich:2019:ATT

[Zak19] Inna Zakharevich. Attitudes of K-theory: topological, algebraic, combinatorial. Notices of the American Mathematical Society, 66(7):1034–1044, August 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zelditch:2019:MQC

[Zel19] Steve Zelditch. Mathematics of quantum chaos in 2019. Notices of the American Mathematical Society, 66(9):1412–1422, October 2019. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ziegler:2011:CPP

[Zie11] Günter M. Ziegler. 3N colored points in a plane. Notices of the American Mathematical Society, 58(4):550–557, April 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ziegler:2018:EOF

[Zie18] Günter M. Ziegler. For example: on occasion of the fiftieth anniversary of Grünbaum’s Convex polytopes. Notices of the American Mathematical Society, 65(5):531–536, May 2018. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zitarelli:2011:SLC

[Zit11] David E. Zitarelli. The 1904 St. Louis Congress and westward expansion of American mathematics. Notices of the American Mathematical Society, 58 (8):1100–1111, September 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zitarelli:2015:DRM

[Zit15] David E. Zitarelli. David Rittenhouse: modern mathematician. Notices of the American Mathematical Society, 62(1):11– 14, January 2015. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zong:2013:WLL

[Zon13] Chuanming Zong. What is … the Leech lattice? Notices of the American Mathematical Society, 60(9):1168–1169, October 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zorn:2012:ODN

[Zor12] Paul Zorn. Our days are numbered: how mathematics orders our lives [book review, McClelland & Stewart Ltd., Toronto, ON, 2009]; The big questions: mathematics [book review, Quercus Publishing, London, 2011]. Notices of the American Mathematical Society, 59(9):1253– 1255, October 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).